A Complete Bibliography of Publications in Applied Mathematics and Computation: 2025–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 November 2025 Version 1.22

Title word cross-reference

(g + 1) [EJJ26]. (k, g) [EJJ26]. (P_{2} ∪ kP_{1}) [HWS25]. 0/1 [ZZL^{+}25]. 1 [HWS25, JSZ25, WWK25]. 1n [AA25]. 2 [ISC25, JCHW25, WHCB25]. 2 − (v, k, 1) [LZCD25]. 2 × 2 [ZMM26]. 3 [MR25, ZHF25, ZZJZ25]. 3 − (v, k, 2) [GLD25]. 4 [LZ25a]. 5 [WWK25]. _{s}igma [ZT25]. A_{4,18}^{b} [AB25]. A_{α} [LL25a]. α [GCV25, ZT25]. B [CCLL25]. C [Wan25]. D^{L} [ZWF25]. D^{Q} [ZWF25]. D_{2n} × Z_{m} [WZY25]. F_{k,4} [LX25]. h [ZBM25, ZYY26]. H^{1} [JT25]. H_{2}/H_{∞} [DZML25]. H_{∞} [LL25b, PPC^{+}25, WL26, WZJW25]. k [GK26, HWS25, WLZ25b]. K_{4} [YZH25]. K_{a,b} [LL25a]. L [BD25]. L2 [JT25]. L^{1} [LLMT25]. L^{p} [CDM25]. l_{0} [WMZ25b]. L_{1} [LZJ^{+}25]. ⌊n/2⌋ [XW25]. M [AA25, BFPS26, MAS26]. Q [ZZ25a]. N [SGH^{+}25, WWLZ25, ZWF25]. p [LHLL25, ZYL25]. {P_{2} ∪ P_{3} [LLW26]. P_{5} [JLJZ26]. Q [ZZ25b]. r [JLJZ26, YZM25]. S [BFH^{+}25, JLG25]. σ [KSFD25]. θ [BDD25, NNW25]. X + AX^{−1}B = Q [MW26]. |b| ≤ 1 [AB25].

-arc-transitive [JLG25]. -Cayley [ZWF25]. -conditional [ZBM25]. -connected [HWS25]. -cycles [EJJ26, WWK25]. -degenerate [WWK25]. -domination [ET26]. -Edge [WHCB25]. -eigenvalues [Wan25]. -equivalent [CCLL25]. -factorization [JSZ25]. -fractal [GCV25]. -free [HWS25, JLJZ26]. -function-based [ZYY26]. -graphs [EJJ26]. -hued [JLJZ26]. -independence [LZ25a]. -integral [ZWF25]. -irregularity [KSFD25]. -Laplacian [ZYL25]. -learning [ZZ25b]. -matchings [GK26]. -matrices [BFPS26].

-methods [BDD25, NNW25]. -minor [LL25a, YZH25]. -norm [JT25]. -normal [LHLL25]. -packing [BFH^{+}25]. -path-edge-connectivity [WLZ25b]. -Planar [WWK25]. -player [WWLZ25, SGH^{+}25]. -regularized [WMZ25b]. -restricted [YZM25]. -robust [ZT25]. -saturating [LX25]. -spectral [LL25a, ZZ25a]. -splines [BD25]. -structure [YZM25]. -tough [HWS25]. -type [CDM25]. -version [JT25]. -walk-regular [ZHF25]. -WG [MAS26]. -Wright [AA25].

1 [Ano25a, Ano25b, Ano25c, Ano25d, Ano25e, Ano25g, Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano26a, Ano26b, ZT25]. 1-planar [TQWW26]. 15 [Ano25m, Ano25n, Ano25o, Ano25p, Ano25q, Ano25s, Ano25r, Ano25t, Ano25u, Ano25v, Ano25w, Ano25x, Ano26d].

2025 [Ano25a, Ano25m, Ano25b, Ano25n, Ano25c, Ano25o, Ano25d, Ano25p, Ano25e, Ano25q, Ano25g, Ano25s, Ano25r, Ano25h, Ano25t, Ano25i, Ano25u, Ano25j, Ano25v, Ano25k, Ano25w, Ano25l, Ano25x]. 2026 [Ano26a, Ano26d, Ano26b]. 2D [GSB26].

362 [PPS25, RRR20]. 395 [AS25].

483 [SS25]. 489 [PPN^{+}25a].

500 [Grz26]. 503 [WMZ25a].

6-coloring [WWL25].

Abel [AZD25]. acceleration [PPS25, RRR19, RRR20]. accounting [LL25d]. accuracy [RPD25]. accurate [ECD^{+}25]. achieved [YK26]. Acoustics [ALC^{+}25]. action [MMNP25]. active [CGH^{+}25]. actuated [WS26]. actuator [BAT25, GLY26, HWC25, SNP25, WZJW25, XZSC26, ZLCS26, ZZ25e]. actuators [CZGY25]. acyclic [TQWW26]. AD [DCC^{+}25]. adaptations [HC26]. Adaptive [LSH25, MLW^{+}25, WS26, YYMY26, ZZ25f, BAT25, CL25, CSB25, GLY26, LHZL25, LYZ^{+}26, LWS^{+}25, LZP25, LYD25, ML26, Mut26, SSK^{+}26, SZSW25, SWLC25, TXJW26, WZJW25, XZSC26, ZLA25, ZLCS26, ZHWW26, ZLCL25]. additive [Kam26, ZD25]. ADI [CLLF25, JCHW25]. adjoint [BD25, ML26]. adjustment [LW25b, TLSVGR25, VL26, XTGZ25]. advection [LWY25, LL25c, MCKZ25, PP25, PSS^{+}25, dOKA25]. advection-diffusion [PP25]. affects [MMTU25]. affine [DZML25, Grz26, IJM25, ZHWW26]. against [CH26, HY25, PPC^{+}25, SNP25, XZSC26, ZZ25c, ZLJY25, ZC25, ZXS^{+}25]. agent [CWS^{+}25, CXZ26, LSP^{+}25, LHZL25, LHZL26a, LLZL26, LZY^{+}25, LZZ^{+}26, MS26, QMWY25, SDH25, SZLG26, WLZ26, XTGZ25, YGY^{+}25, YQQ^{+}26, ZXS^{+}25, ZZS^{+}26, ZLL26, ZZ25e, ZZ26b]. agents [Zha25]. AI [ABC^{+}26]. algebra [AB25]. algebraic [JK25, JZGV25, WLLL26b, WY26a]. algorithm [COP^{+}25, CSB25, GX25, JZGV25, KCLZ25, LSF^{+}25, LZY^{+}25, LWML26, MCHB25, Pal25, YWJ^{+}25]. algorithms [KXL^{+}25, LL25c, LTH^{+}25]. Allee [CMG25]. Allen [OGPFVV25, WWQX25]. allocation [LZY^{+}25, RS25]. allowing [WPTK26]. almost [CMG25]. alternative [IÖÜP25]. altruism [WSY25]. amalgamated [GK26]. American [Hai26]. among [Kis26, YLR^{+}25, Zha25]. Analogy [LLZ26]. Analysis [FBK25, GCV25, LD25, MCKZ25, TF25, TLW^{+}25, ZLHS25, ALC^{+}25, ALV25, BSG^{+}25, BMM25, BM25, CWWT25, COP^{+}25, CRS25, FLSS26, GSB26, KXL^{+}25, KSR25, KGJ25, LLLH25, LZ26, ML26, NVGA^{+}26, PTB25, RPS^{+}25, RR26, SS24, SS25, WF25, WLW26, WLLL26b, WY26a, WLLL26a, XLLC25, YY25a, YYT25, ZFFM25, ZQS26, ZZT25, ZZJZ25]. Analytical [NS25, WLLL26a]. Analytically [Ruj25]. angle [MNE26]. anisotropic [WQLW25]. antagonistic [LHL^{+}26]. Anti [CZGY25, LMQZ25, SP25a]. anti-phase [SP25a]. anti-Ramsey [LMQZ25]. Anti-windup [CZGY25]. antiregular [KSFD25]. antistrong [YM25]. aortic [AT25]. aperiodic [WSX^{+}25, ZXS^{+}25]. aperiodically [WSZM25]. Appl [AS25, PPN^{+}25a, SS25]. Application [RQWX25, ADR^{+}26, BM25, CDYL25, CjZjHbX25, Grz26, KCLZ25, LSF^{+}25, LSP^{+}25, LYLS25, NKSB25, RK25, TQWW26, WZJW25, WS26, ZJX25]. Applications [MMF26, BEG^{+}25, Gao25, Li26, Liu25, MWS25, WX25, WLZ26]. Applied [WMZ25a, PPS25, RRR19, RRR20, TNL25]. approach [Abb26, ALNP25, AAS25, CL25, DGD^{+}26, DAG25, DZML25, EST^{+}25, EP25, FA26, HS25, HZF^{+}25, KM25, LZZ25, LWL^{+}25, LCLY25, LDXL26, LW25a, LD26, MBH^{+}25, MeFW25, PPC^{+}25, QX25b, WZP^{+}25, XC25, XZSC26, YYT25, YWJ^{+}25]. approximate [FLSS26, HYK^{+}25]. approximately [YGY^{+}25]. approximation [AZD25, CCD26, CCC^{+}26, FLLR25, IJM25, LYQZ26, NKSB25, RPD25, VMM26]. approximations [BKE25, BM25]. April [Ano25a, Ano25m]. arbitrarily [GAYKS25]. Arbitrary [NVGA^{+}26, CWWT25]. Arbitrary-order [NVGA^{+}26]. arc [JLG25]. arcs [BK25, SZLG26]. arguments [PPH26, ST25]. arising [LXM25, PMVA26]. arterial [PPS25, RRR19, RRR20]. artificial [AS21, AS25]. Aspects [JDES25, SBK25]. aspiration [HC26, ZWCT26]. aspiration-based [ZWCT26]. aspiration-driven [HC26]. assessment [YZM25]. asset [QZ25]. assets [XZN^{+}25]. assimilation [AT25, JH25]. associated [Gao25]. associative [SNP25]. assumptions [KGJ25]. asymmetric [BK25, GLHS25, LHLL25, SDH25]. asymmetrical [MBH^{+}25]. asymmetrically [PPN^{+}25a, PPN^{+}25b]. Asymptotic [ZQS26, AKM25, LSH25, WDMM25]. Asymptotics [CS25, Mao25]. Asynchronous [ZZ25c, LLWC26, MCH25b, YFZ^{+}25]. atmosphere [MCKZ25]. atom [MK25, YL26]. atom-bond [YL26]. Attack [MCH25b]. Attack-compensated [MCH25b]. attacks [BAT25, CGS^{+}26, JCW26, LHZL26a, SW25, SWW25, WL26, WZJW25, XWW^{+}26, YZ25, ZZ25c, ZLJY25, ZC25, ZLCS26, ZXW^{+}25, ZXS^{+}25, ZLCL25]. attenuated [MH25]. attitude [WS26, ZLCL25]. attractive [CYZ26]. attractivity [ZLW^{+}26]. Augmented [Mut26, BGV26]. August [Ano25b, Ano25n]. automata [WMW^{+}25]. automatic [ML26, NVGA^{+}26]. autonomous [KH26, LWL^{+}25]. auxiliary [LCW25, ZZD26]. avoidance [FA26, ZLA25, ZLL26]. awareness [LH25].

B [PP25]. B-splines [PP25]. background [RS25]. Baker [ACC25]. balanced [BGV26, WLZ25b]. Banach [SKM26]. bandwidth [LHZL26b, WL26, YQQ^{+}26, ZD25]. banner} [LLW26]. banner}-free [LLW26]. Barrier [ZLA25, YWJ^{+}25]. Barrier-function [ZLA25]. barycentric [LL26a]. based [AdFS26, CGS^{+}26, CG25a, CL25, CWWT25, CDYL25, CC26, CYZ26, CSB25, CXZ26, DWL^{+}25, DCC^{+}25, EST^{+}25, FLSS26, GCW^{+}26, GLY26, HYK^{+}25, Han26, HLT25, HWC25, HY25, HCCH25, IÖÜP25, JYLS25, LCW25, LHLL25, LZJ^{+}25, LZD^{+}25, LHZL25, LLZL26, LF26, LWS^{+}25, LL25b, LCRZ25, LLWC26, LZZ^{+}26, LWML26, LM25, Mez25, MNE26, MeFW25, PP25, Pal25, PPC^{+}25, QX25a, QMWY25, RB25, SZM25, SWW25, TF25, TX25, TXJW26, WZWC25a, WSX^{+}25, WWQX25, WXD^{+}25, WLZ25a, WQLW25, WLLL26a, WPTK26, WZWC25b, XC25, YBSC25, YY25b, YZM25, YZ25, YYMY26, YHSZ26, ZWQC25, ZLA25, ZWL25, ZYZ^{+}25, ZLJY25, ZC25, ZZL^{+}25, ZHLL25, ZYY26, ZZD26, ZMM26, ZZTL25, ZXS^{+}25, ZZS^{+}26, ZWCT26]. bases [DEMFRL26]. basis [HS25, LD26, TNL25, WQLW25]. Basket [ZG26a]. bathymetric [RQWX25]. BDF [JZ25]. BDF2 [AZD25]. be [YK26]. bead [MBF25]. bead-rod [MBF25]. beam [NKSB25]. beams [EP25]. behavior [FGJ25, Kab25, KC26, SA26, YNZ25, YH25]. behaviors [SGH^{+}25, WWY^{+}25]. behaviour [LLS^{+}26]. Belonging [QAS26]. Benjamin [WMWY25]. Bernoulli [EP25, NKSB25]. Bessel [ADZ25, Han26, JK25, KXL^{+}25]. best [LLWC26]. best-response [LLWC26]. between [AZZ^{+}26, QZ25]. bi [JLG25]. bi-Cayley [JLG25]. biased [LZD^{+}25]. bidirectional [SNP25]. bifurcation [SS24, SS25]. Bifurcations [ZZJZ25]. bilinear [IJM25]. binary [WLW26]. biodiversity [BFdOdS25]. biological [LM25]. biorthogonal [RNL25]. bipancyclicity [DFKY26]. Bipartite [CWS^{+}25, ZZ26b, CG25b, SDH25, SW25, SHY25, WLZ25b, YT25, YHSZ26]. biquaternion [LLLZ25]. bisimulations [MeFW25]. bit [ZZ26b]. bit-flips [ZZ26b]. Black [XGL25]. blind [IAK25]. block [GLD25, LXM25, LZCD25, ZMM26]. block-transitive [GLD25, LZCD25]. blood [AT25, BGV26, RPS^{+}25]. Board [Ano25y, Ano25z, Ano25-27, Ano25-28, Ano25-29, Ano25-30, Ano25-31, Ano25-32, Ano25-33, Ano25-34, Ano25-35, Ano25-36, Ano25-37, Ano25-38, Ano25-39, Ano25-40, Ano25-41, Ano25-42, Ano25-43, Ano25-44, Ano25-45, Ano25-46, Ano25-47, Ano25-48, Ano26e, Ano26f, Ano26g, Ano26h]. body [PPS25, RRR19, RRR20]. Boltzmann [YQYS25]. Bona [WMWY25]. bond [YL26]. Boolean [ALV25, CLWZ26, CC25, MeFW25, YFZ^{+}25]. Boros [DS25]. bound [LL25c]. bound-preserving [LL25c]. boundary [ALC^{+}25, AT25, BSM^{+}26, BEtM25, GCV25, HRS25, HCTS25, HS25, NS25, OGPFVV25, ZYZ^{+}25, ZQS26]. Bounded [LZZ25, CGS^{+}26, PPC^{+}25]. bounding [ZYY26]. Bounds [ZZ25a, SP25b]. Branch [HHSW25, BVZ25]. Bratu [SSH25]. brief [XS26]. Brinkman [RR26, LL25c]. Brownian [MBF25, WXT25, ZHLL25]. bubble [DFKY26]. bubble-sort [DFKY26]. Buchanan [BS25]. build [ABC^{+}26]. bulk [TFCL25]. bulk-surface [TFCL25]. bullying [Cro25]. Bumpless [WZP^{+}25, ZZ25c]. bundle [HZF^{+}25]. Burgers [WMWY25]. burning [LSQ25].

C [ML26]. cable [LYQZ26]. Cahn [OGPFVV25, WWQX25]. Calculating [Wan25]. calculator [ECD^{+}25]. calculus [LYLS25]. Calibrating [ZG26a]. Campbell [ACC25]. Can [YK26]. canonical [WLLL26b]. capability [LLZ26]. capital [TLSVGR25]. capped [Ruj25]. capped/floored [Ruj25]. Caputo [Abb26, PPH26, ZT25]. cardinality [BFH^{+}25]. Carleman [ALNP25]. case [BS25, MH25, Ruj25, ZHWW26]. Catalyst [QFNW25]. Cauchy [JK25]. Causal [HZ26a]. Cayley [JLG25, WZY25, ZWF25]. cell [BM25, LM25]. cellular [WMW^{+}25]. center [LM25, WHCB25]. center-based [LM25]. chain [HZ25, ZFFM25, ZZT25]. chains [MBF25, ZW25]. Change [HCCH25]. channel [ZZ26b]. channels [YQQ^{+}26]. chaos [ZFFM25]. chaotic [FZY^{+}25, LMMG25]. Characterization [WHCB25, MLZ26, WLLL26a]. characterizing [YNZ25]. Chebyshev [GKM25]. chemical [AD25, ADR^{+}26]. Chialvo [KC26]. chimeras [PPN^{+}25a, PPN^{+}25b]. Chinese [DCC^{+}25]. choice [PDR25]. Choosability [PC25]. chorded [CG25b]. chromatic [LLW26]. circuit [FZY^{+}25]. circuits [LSP^{+}25]. circular [LSF^{+}25]. class [AK26, BAT25, BD25, FGR25, KSR25, KGJ25, MLW^{+}25, MNE26, PDR25, PMVA26, SFL26, WDMM25, Wu25, ZL25, ZYZ^{+}25]. classes [FHX26, JY25b, ZW25]. Classical [YM25, MBH^{+}25]. classification [LZD^{+}25]. classifying [MBH^{+}25]. closed [DZY^{+}25]. closed-loop [DZY^{+}25]. Closeness [FHX26]. clustered [CGS^{+}26]. co [LH25, YH25, ZCJL25]. co-evolution [LH25]. co-evolutionary [YH25]. co-integrated [ZCJL25]. cobweb [LYLS25]. codes [Kov25]. coefficient [JZ25]. coefficients [ALNP25, CWWT25, Hag26, ISC25, LHLL25, Mao25, MW26, QXZ25, TO25, YQYS25]. coevolutionary [GPH25]. Cohen [ZN25]. Cohn [MMF26]. Collaborated [CH26]. collective [ZCJL25, ZWCT26]. collision [ZLA25, ZLL26]. collision-avoidance [ZLL26]. collocation [BSM^{+}26, JLZ25, LL26a, PP25, PTB25, TFCL25, WQLW25]. colored [ZD25]. coloring [JLJZ26, Kam26, TQWW26, WWL25]. colorings [BFH^{+}25]. combination [Liu25]. combinatorial [ALV25, YY25b]. combined [Chr26]. combining [LSF^{+}25, Tak25, Tak26]. Command [YBSC25]. Command-filter-based [YBSC25]. commerce [ZFFM25]. commitment [KH25]. common [GPH25, KGJ25]. communication [LHZL26b, LZY^{+}25, WSX^{+}25, YHSZ26, ZXW^{+}25, ZZS^{+}26, ZZ25e]. communications [RK25]. Community [CCD26, LLMR25]. commutative [JZGV25]. compact [CLLF25, HCTS25, JT25, MR25, WMWY25].

Comparative [BMM25]. compensated [MCH25b]. compensation [GPH25, HLT25]. competition [AZZ^{+}26, SW25]. competitions [PNN25]. competitive [FBK25, WWY^{+}25]. complementarity [GZL^{+}26, Mez25, PSM25]. complements [AAV25, ZMM26]. complete [AD25, JSZ25, LYM25, WLZ25b, WJ25]. completion [HYK^{+}25, LLLZ25]. complex [CC26, LHZL26b, SG26, WSZM25, WZWC25b, XWW^{+}26, ZYL25]. complex-valued [XWW^{+}26, ZYL25]. complexity [MBH^{+}25]. compliance [QWWW26]. Component [ZLW25]. Components [LWLZ26, BQS25]. composed [CRHTV25]. Composite [SSK^{+}26, CL25, WLLL26b]. composition [NKSB25]. compressible [LL25c]. Compt [AS25, PPN^{+}25a]. Comput [SS25]. Computation [JY25b, PPS25, RRR20, WMZ25a, BQS25]. Computational [SBK25, BEG^{+}25]. compute [ACC25]. computing [VL26]. concave [YWJ^{+}25]. condition [HWS25, NNW25]. Conditional [ZZ26a, ZBM25]. conditions [ALC^{+}25, BD25, BDD25, GCV25, KL25, Liu25, MM25, NS25, OGPFVV25, WZXY26, YM25, ZQS26]. conduction [WQLW25, XLLC25]. conduction-convection [XLLC25]. configuration [MBF25]. conflict [WWL25]. conflict-free [WWL25]. conformal [HRS25]. conformists [AZZ^{+}26]. congestion [AKM25]. conical [MH25]. conjecture [BS25, DS25]. conjugate [CFS25, KCLZ25]. connected [BBV25, lCXZ^{+}25, HWS25, LCRZ25, LLZ26, YYZG25]. connectedness [WHCB25]. connectivity [FT25, WLZ25b, YL26, ZLW25, ZZ26a, ZHY26]. Consensus [MS26, YQQ^{+}26, CWS^{+}25, CGS^{+}26, GPL25, HY25, LZZ25, LLZL26, SZLG26, WLZ26, XTGZ25, YHSZ26, ZC25, ZXW^{+}25, ZZS^{+}26, ZZ25e]. conservation [CG25a, KBM25, MCKZ25, SH25, TNL25]. conserving [LCS^{+}26b]. considering [LZ25b, MZX25, WSX^{+}25]. constant [PPH26]. constrained [GPL25, Hai26, LCW25, LHZL26b, LZP25, YYMY26]. constraint [APJ25, AKM25, HWC25, LHLL25, LZJ^{+}25, LDXL26, WZLS26, Wu25]. constraints [FA26, LWS^{+}25, ZHWW26]. Construction [PBS25, AdFS26, RNL25]. consumer [CA25]. consumer-resource [CA25]. contact [CjZjHbX25, TLW^{+}25]. Contagion [KK25a]. containing [SLX^{+}26]. containment [ZXS^{+}25, ZZ26b]. Context [LCS26a]. Context-sensitive [LCS26a]. continuous [ACC25, Seb25, YYT25]. contraction [ADR^{+}26]. Contractivity [BDD25]. contrast [LLQ25]. contribution [ZCJL25]. control [AT25, BAT25, CL25, CLWZ26, CGH^{+}25, DZML25, FY25, FA26, GX25, GCW^{+}26, GPL25, GLY26, HWC25, HY25, HCQ26, HLZ25, IJM25, JXZ^{+}25, JYLS25, KNKR25, KMK26, LZZ25, LCW25, LHLL25, LZJ^{+}25, LSP^{+}25, LHZL25, LHZL26a, LDXL26, LSH25, LWS^{+}25, LLLH25, LC25, LZP25, LCRZ25, LLWC26, MCH25b, MLTL25, MeFW25, Mut26, PFZ25, PPC^{+}25, QX25a, QMWY25, RHSZ26, SDH25, SZM25, SSK^{+}26, SZSW25, SWLC25, SWXL25, SNP25, TXJW26, WDMM25, WZWC25a, WLZ26, WZLS26, WWS^{+}25, WSZM25, WZJW25, WZP^{+}25, WZWC25b, WYKS25, WS26, XC25, XWW^{+}26, XZSC26, YBSC25, YGY^{+}25, YZ25, YFZ^{+}25, YWR^{+}25, YYMY26, YQQ^{+}26, ZWQC25, ZZ25c, ZLA25, ZFW25, ZLJY25, ZD25, ZC25, ZLCS26, ZHWW26, ZLL26, ZLH26, ZZ25e, ZZ26b, ZLHS25, ZLCL25, ZHZ^{+}26]. control-affine [IJM25]. Controllability [LW25d, BEtM25, KH26, YHGZ25]. controller [JXZ^{+}25]. convection [XLLC25, YQYS25]. convection-diffusion [YQYS25]. Convergence [Hag26, KSR25, NNW25, PSS^{+}25, PTB25, WXT25, YY25a, YYT25, BSG^{+}25, BM25, JT25, MCKZ25, PSM25, XTGZ25, ZYZ^{+}25]. convex [YY25a]. convexities [ADPS25]. cooling [LDXL26]. Cooperation [GSL25, QTW25, WSY25, YXG26, BPL^{+}25, BHL25, FWLZ25, GLHS25, HC26, Kis26, LLMR25, LWH^{+}25, LHJ^{+}25, LYZ^{+}26, LF26, LLFS25, LL25d, LW25b, LW25c, LL26b, QFNW25, SW25, SIL25, TS25, Tak25, Tak26, YY25b, YYW^{+}25, YLR^{+}25, ZZZM25, Zha25, ZWL25]. cooperation-competition [SW25]. cooperative [Kab25, MLTL25, WY26b, YY26, ZZ25b]. coordinates [Gao25]. coordination [LLWC26]. copies [LLX25]. core [AMMT26]. correlated [QLCW25]. Correlation [Gao25, QZ25]. Corrigendum [AS25, BDTP25, PPS25, RRR20, SS25, WMZ25a]. cospectral [LSWZ25]. cost [WL26, YWJ^{+}25]. costly [AJK25]. costs [LCS26a]. counterexample [ABD25]. Counting [ZLY26]. Coupled [TGS25, WWY^{+}25, BEtM25, KCLZ25, MCKZ25, NNW25, PPN^{+}25a, PPN^{+}25b, SKM26, SP25a, TFCL25, YWR^{+}25, ZQS26]. coupling [KM25]. covariance [ZFW25]. cover [DFKY26, LCL25, ZHL25]. covers [JS25, QM25]. Cowell [WZXY26]. CPSs [SWW25]. CR [ADR^{+}26]. crash [SL26]. criteria [LZP25]. cross [FBK25]. cross-immunity [FBK25]. crystal [QHZ25]. cubature [Som25]. Cubic [JLG25, ET26, THPA25, FP25]. current [SP25a]. curvature [SBK25]. curves [KMK26]. cyber [CGS^{+}26, DZY^{+}25, SW25, ZLJY25]. cyber-attacks [CGS^{+}26, SW25]. cyber-physical [DZY^{+}25]. Cycle [SP25a, DFKY26, LCL25]. cycles [CG25b, ET26, EJJ26, HWS25, WWL25, WWK25]. cyclic [ZW25]. cylinder [LSF^{+}25, LWLZ26, MH25].

D [ISC25, JCHW25, MR25, ZZJZ25]. damped [CRHTV25]. damping [CS25]. Darcy [LL25c, RR26]. Data [IJM25, QMWY25, SDH25, ALNP25, AT25, BPL^{+}25, CA25, CLWZ26, COP^{+}25, HZF^{+}25, JH25, LHZL26a, MLW^{+}25, RQWX25, TF25, THPA25, WHCB25, WDMM25, WSX^{+}25, WLZ25a, YZ25, YFZ^{+}25]. Data-driven [IJM25, SDH25]. Data-driven-based [QMWY25]. Davey [HAD25]. DCMIP [WZJW25]. DDSIIM [CWWT25]. dead [LHZL25, WZWC25a]. dead-zone [LHZL25]. deadline [QZ25]. deblurring [IAK25]. December [Ano25c, Ano25o]. Decentralized [SW25]. deception [JCW26, WZJW25, ZZ25c, ZLCL25]. decision [YNZ25, ZJX25]. decision-making [YNZ25, ZJX25]. decisions [AJK25]. Decoding [MBH^{+}25]. decomposition [LSF^{+}25, LWML26, LLLZ25]. Decoupled [LL25c, QHZ25]. decoupling [JDES25]. Deep [AAS25, GLY26, EP25, VL26, ZZL^{+}25]. Deep-time [AAS25]. defense [Kab25]. defined [SZSW25, ZCJL25]. degeneracy [YZH25]. degenerate [KK25b, WZJW25, WWK25]. degree [YL26]. degrees [GAYKS25]. delay [BKE25, CZGY25, Dib26, DZY^{+}25, EWH25, GSB26, LLLH25, LW25d, MMF26, PFZ25, PTB25, SWW25, SGH^{+}25, WSX^{+}25, WZXY26, WY26a, YHGZ25, YZ25, ZZ25d]. delay-integro-differential-algebraic [WY26a]. delayed [CCLL25, LSP^{+}25, LL25b, LLS^{+}26, SA26, SZY25, SNP25, WXT25, WDMM25, WSX^{+}25, XWW^{+}26, ZYY26, ZHZ^{+}26]. delays [HLZ25, NSX^{+}25, ZN25]. demand [ZLHS25]. denoising [LLMT25]. Density [ZG26b, BCCF25]. dependent [ABD25, AT25, BQS25, CCLL25, ISC25, LZJ^{+}25, QTW25, YQYS25]. depression [LZD^{+}25]. derivation [RR26]. derivative [CK25, LT25, PPH26, WQLW25]. derivatives [AS21, AS25]. descent [YWJ^{+}25]. described [XZN^{+}25]. description [MMNP25]. design [CZGY25, ZKZ26]. designs [GLD25, LZCD25]. detection [CCD26, HCCH25]. determinant [Ued25, ZJX25]. Determination [JH25]. deterministic [DAG25, MMTU25]. developments [XS26]. device [WZJW25, YK26]. diagnosis [LLZ26]. diagonalization [LCLY25]. diameter [LSWZ25, WJBL25]. diffeomorphism [LBLZ25]. difference [CPR26, Dib26, JT25, OGPFVV25, SSH25, WMWY25, WH26, XGL25]. different [MS26]. different-order [MS26]. differentiable [HLZ25, ZL25]. differential [Abb26, AZD25, BBG^{+}26, CjZjHbX25, GSB26, GPL25, HS25, JY25a, LZ25b, LD26, MMF26, PPH26, QXZ25, SS24, SS25, TX25, WZXY26, WLLL26b, WY26a, WPTK26, XZN^{+}25, YYT25, ZLSQ26]. differential-algebraic [WLLL26b]. differentiation [ML26, NVGA^{+}26]. diffusion [BM25, CS25, Grz26, ISC25, LWY25, LL25c, LD26, MZX25, MCKZ25, PP25, PSS^{+}25, QX25b, RHSZ26, Ruj25, WWY^{+}25, WWG^{+}25, YQYS25, ZT25, ZWL25, dOKA25]. diffusive [WF25]. digraphs [AHMSV25, JS25, YM25]. dilemma [BHL25, FWLZ25, LLMR25, LWH^{+}25, LW25b, LW25c, QLCW25, TS25, Tak26, YYW^{+}25, ZWCT26, KH25]. dilemmas [DGD^{+}26, LYZ^{+}26, QTW25, WL25, YNZ25]. dilute [MBF25]. dimensional [AAS25, EP25, FGR25, LJ25, MMF26, PBS25, RQWX25, SST26, SSH25, Wan25, WQLW25, WMW^{+}25, WWS^{+}25]. dimensionality [MeFW25]. dimensions [PJK25]. directed [CWS^{+}25, LZY^{+}25]. direction [YWJ^{+}25]. directional [NS25]. directions [QX25a, SZLG26]. directly [Wan25]. Dirichlet [ALC^{+}25]. Discerning [ABC^{+}26]. disclosure [QWWW26]. discontinuities [WH26]. discontinuous [CWWT25, Hag26, JY25a, WF25]. Discrete [DZY^{+}25, MK25, AZD25, Dib26, HLZ25, IJM25, JCHW25, JT25, LYLS25, LZ26, ML26, NSX^{+}25, Ruj25, SZM25, ZYL25, ZYY26, ZKZ26]. Discrete-time [DZY^{+}25, HLZ25, IJM25, NSX^{+}25, SZM25, ZYY26, ZKZ26]. discretized [XGL25]. disease [FBK25, LGW^{+}26]. Disjoint [QM25, DFKY26, JS25, LLX25, LCL25]. disordering [JXZ^{+}25]. disparities [BHL25]. displacement [GKR^{+}25]. dissipative [LDXL26, LLLH25, SZM25, WWS^{+}25]. dissociation [HZ26b]. Distance [AHMSV25, ZW25, BFH^{+}25, WQLW25]. distances [JY25b]. distinct [GAYKS25]. Distinguishing [AAV25, Sta25]. Distributed [CXZ26, LYD25, LZZ25, LZY^{+}25, LC25, LZZ^{+}26, QMWY25, WLZ26, WPTK26, ZXW^{+}25, ZZ25e, ZZ25f]. Distribution [HHSW25, MP25, QX25b, RMS25, YXG26]. Disturbance [LWS^{+}25, GCW^{+}26, SWXL25, YYMY26, ZZ25d]. disturbances [CWS^{+}25, CGH^{+}25, SSK^{+}26, WSZM25, YGY^{+}25, ZLL26]. diverse [SGH^{+}25]. diversity [SIL25]. do [NS25]. do-nothing [NS25]. domain [MCH25a, WLLL26a]. domains [CWWT25, Chr26, TKM25]. domatic [GVW25]. domination [ET26, GXY25, GVW25]. domino [LWLZ26]. double [AK26, CWS^{+}25, FA26, GXY25, RKS25, ZLJY25]. double-layer [ZLJY25]. drawdown [SL26]. Drazin [AMMT26]. drift [Hag26]. driven [lCXZ^{+}25, CC26, GPL25, HC26, IJM25, LHZL26b, LW25b, LW25c, QMWY25, Ruj25, SDH25, WXT25, WL25, ZYL25, ZZ25b]. drives [ZZZM25]. Dual [LCW25, MZX25, MWS25, ZZ25b]. dual-reputation [ZZ25b]. due [DS25, GSB26]. Durán [RBT25]. Durán-type [RBT25]. Dynamic [GPL25, LHZL26a, ZZZM25, ZLCL25, CWS^{+}25, CLLF25, CDYL25, HWC25, HC26, LSF^{+}25, LYZ^{+}26, LDXL26, LLZL26, PPC^{+}25, QZ25, SA26, SW25, SWW25, VL26, WZP^{+}25, WYKS25, YYMY26, ZLH26]. dynamical [CGH^{+}25, SKSC26]. dynamical-systems [SKSC26]. Dynamics [KC26, LGW^{+}26, YY26, ALV25, ABdF^{+}25, CMG25, CA25, Che25, Cro25, DGD^{+}26, HC26, IJM25, LLWC26, LDT^{+}25, MBH^{+}25, MMTU25, QAS26, QWWW26, QTW25, SP25a, SG26, SGH^{+}25, Tak26, WWY^{+}25, Wet25, YNZ25, ZWL25].

E-commerce [ZFFM25]. E2P [ZKZ26]. eavesdropping [HY25]. eccentricity [HZ25]. eco [SKSC26]. eco-epidemic [SKSC26]. ecological [GPH25, LLS^{+}26]. ecosystem [ZWL25]. Edge [WHCB25, WLZ25b, WLW26, XTGZ25, ZZ26a, ZHY26]. edge-connectivity [ZHY26]. edge-heterogeneous [WLW26]. edges [BBV25, JS25, LX25]. Editorial [Ano25y, Ano25z, Ano25-27, Ano25-28, Ano25-29, Ano25-30, Ano25-31, Ano25-32, Ano25-33, Ano25-34, Ano25-35, Ano25-36, Ano25-37, Ano25-38, Ano25-39, Ano25-40, Ano25-41, Ano25-42, Ano25-43, Ano25-44, Ano25-45, Ano25-46, Ano25-47, Ano25-48, Ano26e, Ano26f, Ano26g, Ano26h]. Effect [LH25, CMG25, GPH25, MZX25, MLTL25]. Effective [COP^{+}25, ABC^{+}26]. Effects [YY25b, FBK25, KH26, LYD25, SGH^{+}25]. efficacy [XZSC26]. Efficient [HRS25, LT25, TKM25, AAS25, ACC25, CRHTV25, YYT25]. efficiently [HS25, LD26]. Eigenvalue [SP25b, CCC^{+}26, LYM25, MP25, XSB25, ZHF25, ZQS26]. eigenvalues [GAYKS25, Wan25]. eighth [CTTN25]. eighth-order [CTTN25]. elastic [KM25]. electro [PPS25, RRR19, RRR20]. electromagnetic [KC26]. element [AZD25, BRZS25, EP25, HJF25, HAD25, LT25, LD25, LJ25, Tan25, TLW^{+}25, TKM25, ZT25]. element-free [LD25]. elements [HS25, NVGA^{+}26]. elliptic [BSM^{+}26, CWWT25, GSB26, PMVA26]. emergencies [LLS^{+}26]. emotion [LL26b]. enabled [ML26]. end [BD25]. endogenous [GSL25]. enemy [ZZJZ25]. energies [ZW25]. Energy [LCS^{+}26b, AKM25, MCKZ25, SFL26, ZHLL25, ZLCS26]. Energy-conserving [LCS^{+}26b]. enforcement [LCS26a]. enhanced [HY25, LT25, ZYZ^{+}25]. Enhancement [LDT^{+}25]. enhances [FWLZ25, LHJ^{+}25]. entropy [BGV26, CLLC25]. Enumerating [GK26]. environment [WY26b, ZHLL25]. Environmental [LWH^{+}25, LHJ^{+}25, LF26, WL25, YLR^{+}25]. Environmental-feedback-driven [WL25]. environments [ACCO25, MMTU25]. EP [AMMT26]. Epidemic [FBK25, Chr26, LLS^{+}26, SKSC26, SLX^{+}26]. epidemics [CH26, LH25, MMNP25, YH25]. epistemic [LZ25b]. epitaxial [PMVA26]. equal [LL26a, XW25]. equality [LL25d]. equation [AZD25, Dib26, FP25, GKM25, ISC25, JCHW25, LL26a, LYQZ26, LZ25b, LD26, LTH^{+}25, MW26, NKSB25, PP25, QX25b, SSH25, WF25, WWQX25, WMWY25, WGeL25, WPTK26, YQYS25]. equation-based [WPTK26]. equations [Abb26, ALNP25, ALC^{+}25, BBG^{+}26, BKE25, BMM25, CLLF25, CYZ26, CRHTV25, DN25, FLLR25, GSB26, HAD25, HS25, JY25a, JT25, JLZ25, JZ25, KSR25, KCLZ25, LCS^{+}26b, LZ26, MMF26, MR25, NS25, OGPFVV25, PSS^{+}25, PTB25, QXZ25, SKM26, SST26, SS24, SS25, SFL26, TGS25, Tan25, TX25, TZL26, TKM25, WQLW25, WZXY26, WLLL26b, WY26a, XLLC25, XZN^{+}25, YYT25, ZYL25, ZT25, ZL25]. equienergetic [LSWZ25]. equilibria [BAG^{+}25]. Equilibrium [WLW26, CA25]. equitable [AdFS26]. equivalent [CCLL25, Liu25]. Erdos [MMNP25]. Erratum [PPN^{+}25a]. error [CWWT25, CLLC25, HAD25, YYMY26]. error-based [YYMY26]. errors [ZJX25]. Escher [OCN^{+}26]. Escher-like [OCN^{+}26]. ESO [LCRZ25]. ESO-based [LCRZ25]. essence [Cam26]. estimates [TNL25]. Estimation [XZN^{+}25, ZWQC25, BBG^{+}26, EST^{+}25, HAD25, LL25b, LZ25b, LYD25, SZY25, SZM25, WXD^{+}25, ZZD26, dOKA25]. Estimation-based [ZWQC25]. estimator [TXJW26]. estimator-based [TXJW26]. ETF [WZJW25]. Euler [AKM25, EP25, NKSB25, TKM25, WLLL26b]. evader [ZLSQ26]. evaluation [Han26, ZBM25]. Event [APJ25, HLT25, HY25, HCQ26, LHLL25, PFZ25, QX25a, SWLC25, YGY^{+}25, ZZ25e, CWS^{+}25, DZY^{+}25, GPL25, HWC25, JCW26, LZJ^{+}25, LHZL25, LHZL26a, LLZL26, LC25, LZZ^{+}26, Mut26, PPC^{+}25, QMWY25, SW25, SWW25, SZSW25, SWXL25, WZP^{+}25, WYKS25, XWW^{+}26, YBSC25, YZ25, ZWQC25, ZLCS26, ZXS^{+}25, ZZS^{+}26, ZLCL25, ZHZ^{+}26]. Event-based [HY25, LHLL25]. event-driven [GPL25]. Event-triggered [APJ25, HLT25, HCQ26, PFZ25, QX25a, SWLC25, YGY^{+}25, ZZ25e, CWS^{+}25, DZY^{+}25, HWC25, JCW26, LZJ^{+}25, LHZL25, LHZL26a, LC25, LZZ^{+}26, Mut26, PPC^{+}25, QMWY25, SW25, SWW25, SZSW25, SWXL25, WYKS25, XWW^{+}26, YZ25, ZWQC25, ZLCS26, ZXS^{+}25, ZZS^{+}26, ZLCL25, ZHZ^{+}26]. event-triggered-based [LLZL26]. event-triggering [WZP^{+}25, YBSC25]. events [SKSC26]. Evolution [GLHS25, LYZ^{+}26, LLS^{+}26, SIL25, Tak25, Zha25, BHL25, DWL^{+}25, KH26, LH25, MCH25a, QFNW25, TF25, WSY25, YY25b, YY26, ZZ25b, ZWCT26]. Evolutionary [AZZ^{+}26, BPL^{+}25, Tak26, HC26, LLMR25, LWL^{+}25, LL25d, LLWC26, MZX25, QWWW26, SGH^{+}25, XS26, YH25, ZZTL25, ZLHS25]. exchange [LHZL26b]. Exemplary [FWLZ25]. exhibiting [GSB26]. Existence [PMVA26, ZL25]. expansion [BSG^{+}25, Han26]. expectation [LL26b]. Experience [LW25b]. Experience-driven [LW25b]. experimental [FZY^{+}25]. Explicit [EWH25, THPA25, ML26, dOKA25]. exploitative [PNN25]. Exploring [ABdF^{+}25, MMNP25, YT25, ZWL25]. Exponential [CCLL25, LWY25, MR25, SFL26, XWW^{+}26, ZN25]. expressions [WWLZ25]. extend [CLLC25]. Extended [CC25, LLLH25, WLLL26b]. extension [AB25]. external [WY26b, ZLL26]. extinction [CMG25]. extrema [LZ25a]. Extremal [CG25b, LL25a]. extreme [SKSC26]. Extremizing [KSFD25].

face [WY26b]. facility [RMS25]. factorization [JSZ25, DCC^{+}25]. factors [YH25]. fading [YQQ^{+}26]. failure [HHSW25, SNP25]. failures [ZLCS26]. fairness [DWL^{+}25, Kis26, Zha25]. fairness-seeking [Zha25]. false [LHZL26a, YZ25]. Fan [HWS25]. Fan-type [HWS25]. Fast [KXL^{+}25, ZZS^{+}26, CYZ26, ECD^{+}25, Pal25, Tan25, XGL25]. Fault [KNKR25, XZSC26, YZM25, CGH^{+}25, EST^{+}25, HLT25, HY25, LCW25, LLZ26, PPC^{+}25, SSK^{+}26, SZSW25, SWLC25, WS26, ZLCS26]. fault-compensation-based [HLT25]. fault-tolerance [HY25]. Fault-tolerant [KNKR25, XZSC26, CGH^{+}25, LCW25, PPC^{+}25, SZSW25, ZLCS26]. faults [BAT25, CGH^{+}25, GLY26, HWC25, WZJW25, XZSC26, YHSZ26, ZHWW26, ZZ25e]. faulty [JS25]. Favard [Yu25]. Favard-type [Yu25]. fear [Kab25, MLTL25]. fear-induced [Kab25]. feature [LZD^{+}25]. Features [FP25]. February [Ano25d, Ano25p, Ano26c]. feedback [BAT25, BS25, CL25, GX25, GCW^{+}26, GLY26, HLZ25, HZ26a, LWH^{+}25, LDXL26, MLW^{+}25, MCH25b, PFZ25, SNP25, WL25, WZLS26, YLR^{+}25]. feeding [CA25]. FFT [CWWT25]. FFT-based [CWWT25]. fidelity [LLMT25]. fiducial [CLLC25]. field [HJF25, MMNP25, PPS25, QHZ25, RRR19, RRR20, ZZT25]. fields [MBF25]. Filippov [ZZT25, ZZJZ25]. film [PBS25]. filter [CL25, CLLC25, YBSC25, ZKZ26]. filter-based [CL25]. filtering [BGQR25, WL26]. Finite [BRZS25, IJM25, WWS^{+}25, ZKZ26, CF25, CPR26, DZML25, EST^{+}25, EWH25, HJF25, HLT25, HWC25, LT25, LZJ^{+}25, LJ25, Mut26, NVGA^{+}26, OGPFVV25, PPC^{+}25, SSH25, Tan25, TLW^{+}25, WMW^{+}25, WMWY25, WLZ26, WH26, ZT25, ZD25, ZJX25, ZZS^{+}26, ZLL26]. finite-difference [CPR26, OGPFVV25]. finite-horizon [DZML25]. Finite-region [WWS^{+}25]. Finite-time [ZKZ26, EST^{+}25, EWH25, HLT25, HWC25, LZJ^{+}25, PPC^{+}25, WLZ26, ZZS^{+}26, ZLL26]. First [CTTN25, GKM25, ZQS26]. Fixed [CGH^{+}25, GCW^{+}26, LZP25, LCRZ25, MLW^{+}25, QX25a, SZSW25, WZWC25b, ZWQC25, ZZTL25]. Fixed-time [CGH^{+}25, GCW^{+}26, LZP25, LCRZ25, MLW^{+}25, QX25a, SZSW25, WZWC25b, ZWQC25]. flight [KNKR25]. flip [LWLZ26]. flipped [BEG^{+}25]. flips [ZZ26b]. flocks [QAS26]. floored [Ruj25]. flow [AT25, BGV26, LD25, LSF^{+}25, LL25c, MBF25, PPS25, PBS25, RRR19, RRR20, SBK25]. flows [lCXZ^{+}25]. fluid [KM25, KBM25, PPS25, RRR19, RRR20, RR26, ZG26b]. fluid-poroelastic [KBM25]. fluid-porous [RR26]. fluid-structure [ZG26b]. flux [KC26]. follower [SZLG26, YY26, YHSZ26]. following [CWS^{+}25, LLZL26, Wet25]. food [ZZT25]. foraging [QAS26]. forced [SKSC26]. forced-transmission [SKSC26]. forces [LM25]. form [KCLZ25, KL25, MLZ26, MR25, TGS25]. formation [KNKR25, YGY^{+}25, ZLH26]. forms [AAV25, LBLZ25, THPA25]. formula [ACC25]. formulae [WX25]. formulas [MH25]. four [LSWZ25]. Fourier [CDM25, Gao25]. fourth [JT25, KGJ25, PP25, PMVA26]. fourth-order [JT25]. fractal [GCV25, WQLW25]. fractional [Abb26, BGI25, CLLF25, CjZjHbX25, CDM25, CYZ26, DN25, EWH25, FZY^{+}25, Gao25, Han26, HZ26a, JCHW25, JLZ25, LLMT25, LWY25, LYQZ26, LMMG25, LYLS25, NKSB25, PDR25, PPH26, QX25b, SHY25, Tan25, TX25, WMZ25a, WXT25, WMZ25b, WWQX25, WXD^{+}25, WLLL26a, XZN^{+}25, XGL25, YHGZ25, YWR^{+}25, ZT25, ZN25, ZHLL25, ZZD26, ZLW^{+}26]. fractional-order [WXD^{+}25, WLLL26a, YHGZ25, ZZD26, ZLW^{+}26]. fractured [LL25c]. fragile [SNP25, ZLJY25]. framework [DZML25, HY25, LZD^{+}25, Mez25, MAMFBM^{+}25, RMS25, TLW^{+}25]. frameworks [MBH^{+}25]. Fredholm [QXZ25]. free [HWS25, JLJZ26, Kis26, LT25, LD25, LLZ26, LLW26, QX25a, SZSW25, Sta25, WWL25, YZH25]. free-riders [Kis26]. frequency [NVGA^{+}26, ZLCS26]. frictional [CjZjHbX25]. frictions [RS25]. Frobenius [SP25b]. frozen [ST25]. FTC [CXZ26, HLT25]. FTPP [LSH25]. full [LWS^{+}25, SZM25]. Fully [LC25, ZXW^{+}25, JCHW25, JT25, JZ25, QMWY25, WS26]. fully-geometric [JZ25]. function [ADZ25, AA25, CG25b, LD26, TNL25, WQLW25, ZLA25, ZZL^{+}25, ZYY26]. functional [BMM25, WLLL26b]. functions [ABD25, CC25, GCV25, HS25, MCHB25, NVGA^{+}26, ZL25]. fundamental [ZG26b]. Fusion [DCC^{+}25]. Fuzzy [LYLS25, EST^{+}25, HLT25, KL25, LHZL26a, LDXL26, LWS^{+}25, LLLH25, PPH26, RK25, YYMY26, ZWQC25, ZLJY25, ZHLL25, ZLCS26, ZZ25d, ZKZ26, ZHZ^{+}26]. FVCOM [RQWX25].

gain [JXZ^{+}25, LSP^{+}25, LCRZ25, YYMY26]. gain-scheduling [JXZ^{+}25]. gain-waving [LSP^{+}25]. Galerkin [JY25a, LD25, WF25, YYT25]. Game [ZFFM25, BAG^{+}25, BHL25, DGD^{+}26, DZML25, GSL25, GPH25, HCQ26, LLMR25, LWH^{+}25, LWL^{+}25, LW25b, LW25c, LLS^{+}26, MZX25, SGH^{+}25, Tak25, Wet25, XS26, YYW^{+}25, ZZZM25, Zha25, ZZTL25, ZLHS25]. game-based [ZZTL25]. game-level-headed [LLS^{+}26]. games [BPL^{+}25, CC26, GPL25, IÖÜP25, LCS26a, LHJ^{+}25, LLFS25, LL25d, LLWC26, QLCW25, SIL25, Tak25, Tak26, Ued25, WSY25, WWLZ25, WLW26, ZXP^{+}25, ZJX25, ZLSQ26]. gaps [SG26]. gasket [GCV25]. gauge [SHY25]. Gauss [KMK26]. Gaussian [Hai26, ZD25]. GBF [CCD26]. GBF-PUM [CCD26]. gender [LZD^{+}25]. general [AD25, ADPS25, RKS25, TGS25, TO25, Wu25]. Generalized [WZXY26, AK26, HZF^{+}25, ISC25, KXL^{+}25, KCLZ25, LZ25a, LJ25, MLZ26, PJK25, WMWY25, Wu25, YY25a, YFZ^{+}25, ZHLL25, ZHL25]. Generalizing [Li26]. generally [PPC^{+}25]. generic [CA25]. genuine [ZHF25]. geometric [JZ25]. geometrical [BGV26]. geometries [DEMFRL26]. geostrophic [BGQR25]. Ginzburg [ZYL25]. given [AD25, LZ25a, YL26]. GL-ADI [JCHW25]. glioma [BM25]. Global [WDMM25, BDD25, FLLR25, LHLL25, XLLC25, YK26, ZLW^{+}26]. glucose [AS21, AS25]. glucose-insulin [AS21]. Godunov [RPD25]. Godunov-type [RPD25]. goods [CC26, GSL25, LCS26a, LHJ^{+}25, LLFS25, SIL25, WSY25, ZZZM25, Zha25, ZXP^{+}25]. Gordon [CLLF25]. governing [PBS25]. Gower [SA26]. gradient [CFS25, KCLZ25, LLQ25]. granular [LYLS25]. graph [ADPS25, CG25b, FHX26, GXY25, GED26, HCCH25, JLJZ26, TQWW26, WLZ25b]. graphical [GPL25]. Graphs [DMT25, GAYKS25, LSQ25, XW25, AAV25, AdFS26, BBV25, BFH^{+}25, BDTP23, BDTP25, CWS^{+}25, CCD26, CDD^{+}25, DFKY26, ET26, EJJ26, GK26, GMOS26, GVW25, HWS25, HZ25, HZ26b, JY25b, JLG25, KSFD25, LL25a, LYM25, LZ25a, LLZ26, LWLZ26, LLW26, MBH^{+}25, MP25, PC25, PJK25, QM25, RKS25, Sta25, WZY25, WWL25, WJ25, WMW^{+}25, WJBL25, WWK25, YT25, YZM25, YZH25, YL26, YYZG25, ZHF25, ZLY26, ZWF25]. Gravitational [MNE26]. green [WWY^{+}25]. grid [JZ25, RQWX25, Tan25, YY26, ZT25]. grids [JT25, WLZ26, XGL25, ZLHS25]. Grossberg [ZN25]. Group [QAS26, Wet25, LLS^{+}26, ZJX25]. groups [YLR^{+}25]. growth [BDL25, PMVA26]. Grünwald [LLMT25]. guaranteed [WL26]. guarantees [CXZ26]. guidance [YYW^{+}25].

H [DZY^{+}25]. H-infinity [DZY^{+}25]. Hadamard [ZT25]. Hamilton [MP26]. Hamiltonian [BRZS25, LCS^{+}26b, WHCB25]. Hammerstein [SST26, WXD^{+}25]. Hamming [YZM25]. hand [CF25]. handling [RNL25]. Hankel [Han26]. harmonic [KK25b]. Harnessing [LMMG25]. Hartmann [LD25]. harvesting [MMTU25]. Hausdorff [ACC25, WQLW25]. having [GAYKS25, MK25]. headed [LLS^{+}26]. heat [BEtM25, WQLW25]. Heisenberg [CDM25]. Hemivariational [KH26, CjZjHbX25]. Hermite [ADZ25, WWQX25]. hesitancy [LdMAS25]. heterogeneous [CGS^{+}26, LYZ^{+}26, LZY^{+}25, QLCW25, WLW26, ZWL25, ZZ25e]. HHO [LZD^{+}25]. hidden [LW25a, ZZ25c]. Hierarchical [ZZD26, FT25, YXG26, ZZ26a]. hierarchy [YY26]. High [CPR26, AAS25, CLLC25, HLZ25, LT25, LH25, MMF26, MR25, QXZ25, SSH25, WMWY25, WS26, YY25b, ZWQC25, ZLA25, ZYZ^{+}25, ZZS^{+}26, ZLL26]. high-dimensional [AAS25, MMF26, SSH25]. high-order [CLLC25, LT25, LH25, MR25, QXZ25, WMWY25, WS26, ZWQC25, ZLA25, ZYZ^{+}25, ZZS^{+}26, ZLL26]. high-reputation [YY25b]. higher [KK25b, LDT^{+}25, SLX^{+}26, SG26, SWLC25, YWR^{+}25, ZYZ^{+}25]. higher-order [KK25b, LDT^{+}25, SLX^{+}26, SG26, SWLC25, YWR^{+}25]. highly [BFdOdS25, KSR25]. Hilbert [Liu25, QXZ25]. Hilbert-type [Liu25]. Hinderance [Tak26]. holistic [IÖÜP25]. homogeneous [LJ25, PPH26]. homotopies [HHSW25]. horizon [DZML25, HZ26a, LYD25]. horizontal [Mez25]. hot [LDXL26]. hued [JLJZ26]. human [lCXZ^{+}25, LW25a, YNZ25, YYW^{+}25]. human-driven [lCXZ^{+}25]. human-machine [LW25a]. hunt [Tak25, Tak26]. hunting [MLTL25, ZWL25]. Hybrid [MLTL25, EP25, HLZ25, LSF^{+}25, LLZ26, SW25, SWW25, WL26, XWW^{+}26]. Hybridizable [JY25a, WF25]. hybridization [BRZS25]. hydrodynamic [PPS25, RRR19, RRR20]. hydrodynamics [RPD25]. hydrogen [MK25]. hyperbolic [CMG25, CS25, JY25a, OCN^{+}26, PP25, SBK25, SH25]. hyperchaotic [RK25]. hypercomplex [NVGA^{+}26]. hypercube [CDD^{+}25]. hypercube-like [CDD^{+}25]. hypercubes [FGW25, KV25, LMQZ25]. hypergraph [JSZ25]. hypergraphs [MAMFBM^{+}25, PDV25, ZZ25a, ZHY26]. hyperinterpolation [Som25]. hypernetworks [LHL^{+}26]. hypersingular [GKM25].

ideals [AHMSV25]. identification [QX25b, WZWC25b]. identifying [ISC25, LMMG25, TQWW26]. identities [KK25b]. ill [Li26]. ill-posed [Li26]. image [CRS25, CSB25, IAK25, LLQ25]. imbalance [MNE26]. imitation [DGD^{+}26]. imitators [AZZ^{+}26]. immersed [HCTS25]. immobile [Tan25]. immunity [FBK25]. impact [QLCW25]. impatient [LLS^{+}26]. Impenetrability [ALC^{+}25]. imperfect [LdMAS25]. imperfections [ZZ26b]. implementation [FZY^{+}25, ML26, Ued25, ZJX25]. implicit [MR25, XGL25]. improve [XTGZ25]. Improved [FLSS26, MCHB25, WH26, SH25]. improves [LZD^{+}25, LL26b]. Improving [RPD25, ZZ25d, AT25]. Impulse [CGS^{+}26]. Impulse-based [CGS^{+}26]. impulses [CCLL25]. impulsive [CjZjHbX25, JCW26, KH26, MM25, PFZ25, SWXL25, WZJW25, YHGZ25, ZN25]. In-phase [SP25a]. Incentive [LLWC26, YXG26, ZZ25b]. Incentive-based [LLWC26]. incidence [ZZ25a]. inclusions [CjZjHbX25]. Incorporating [YH25, QHZ25]. incremental [RHSZ26]. indefinite [LXM25]. independence [LZ25a]. independent [CG25b, WJ25]. index [AD25, GED26, WJBL25, YL26]. indices [BDTP23, BDTP25, HZ25, JY25b, WZY25, ZHL25]. indirect [QFNW25]. individual [LH25, Tak26]. Induced [ET26, Kab25, YYZG25]. inequalities [HZF^{+}25, KH26]. inequality [KL25, WSX^{+}25, Yu25]. inertia [HZ25]. inexact [HZF^{+}25]. infection [Kab25]. inference [CA25]. infinite [CF25, HZ26a, MW26, Sta25]. infinitesimal [CF25]. infinity [DZY^{+}25]. Influence [LdMAS25, BHL25, DAG25, FWLZ25, HC26, WWG^{+}25]. information [ABdF^{+}25, LHJ^{+}25, LHZL26b, LGW^{+}26, MZX25, SZM25, SLX^{+}26, WWY^{+}25, YY26, YH25]. initial [QX25b, VL26, WZXY26, ZT25]. injection [BAT25, LHZL26a, LBLZ25, YZ25]. inner [Li26]. Input [SWXL25, CL25, LCW25, LDXL26, LLZL26, Mut26, QX25a, SDH25, WZLS26, WSZM25, ZZ25f]. input-constrained [LCW25]. input-output [Mut26]. Input-to-state [SWXL25]. inputs [SZM25]. Insider [QZ25]. insulin [AS21, AS25]. insurable [RS25]. integrability [AB25]. Integral [PDV25, dOKA25, BKE25, BMM25, FLLR25, FA26, GKM25, HCQ26, KSR25, Liu25, LLLH25, PTB25, SKM26, SST26, TGS25, WSX^{+}25, Yu25, ZL25, ZWF25]. Integrated [RMS25, ZCJL25, ZZT25]. Integrating [LL26b]. integration [LM25]. integro [AZD25, JY25a, QXZ25, WZXY26, WY26a, WPTK26, YYT25]. integro-differential [AZD25, JY25a, QXZ25, WZXY26, WPTK26, YYT25]. Interaction [XTGZ25, FP25, KBM25, LH25, ZG26b]. interactions [LDT^{+}25, MCKZ25, QHZ25, YWR^{+}25]. interactive [LW25b, SIL25]. interconnected [JYLS25]. interdependence [ZWCT26]. interdependent [LLFS25]. interface [CWWT25, LJ25]. interference [PNN25, YFZ^{+}25]. Intergroup [CC26]. Intergroup-based [CC26]. interior [GSB26]. intermittent [CGH^{+}25, LSP^{+}25, WSZM25, YWR^{+}25]. internal [BEtM25]. interneurons [SP25a]. interpolating [KMK26, THPA25]. interpolation [BD25, RQWX25]. intersecting [PC25]. Interval [FY25, EST^{+}25, LHZL26a, LLLH25, SP25b, ZZ25d]. Interval-oriented [FY25]. intervention [WWY^{+}25]. Introduction [Kam26]. invariant [AB25]. Inverse [ISC25, RHSZ26, ST25, ACCO25, BFPS26, DN25, FLSS26, MAS26, XSB25, ZMM26]. inverses [MWS25]. Inversion [MH25].

Investigating [HC26]. investment [RS25, SGH^{+}25]. Invitation [KSSY26]. involving [ADR^{+}26, PPH26, ZT25]. irregular [LSP^{+}25, MCH25a]. irregularity [FGW25, GMOS26, KSFD25]. isometries [AMMT26]. Isospectral [BDST25]. issue [QLCW25]. iteration [ADPS25, GZL^{+}26]. iterations [ADR^{+}26]. iterative [CTTN25, XZSC26].

Jacobi [JLZ25, RB25, XSB25]. January [Ano25e, Ano25q, Ano26a, Ano26d]. Jarratt [BSG^{+}25]. Jarratt-type [BSG^{+}25]. job [AJK25]. join [ZLY26]. joint [KH25, SWW25]. July [Ano25r, Ano25f]. jump [JYLS25, PPC^{+}25, RHSZ26, Ruj25, SW25, SSK^{+}26, WZJW25, XZSC26, ZZ25c]. jump-diffusion [RHSZ26, Ruj25]. jumping [LL25b, WWS^{+}25]. June [Ano25g, Ano25s].

Kalman [CLLC25]. Kansa [LCS^{+}26b]. kernel [AZD25, CDYL25, Liu25, QXZ25, RB25, XC25]. kernels [KSR25, PTB25]. Keyfitz [PBS25]. kind [FLLR25, GKM25]. kinetics [AS21, AS25]. Kirchhoff [GED26, JY25b, WZY25]. Klein [CLLF25]. Koopman [IJM25]. Korteweg [SFL26]. Kranzer [PBS25]. Kutta [ML26, WY26a]. KVA [TLSVGR25].

l0 [WMZ25a]. l0-regularized [WMZ25a]. L2 [ZT25]. L2-1 [ZT25]. labelling [BVZ25]. labor [ZCJL25]. ladder [Seb25]. lag [MLTL25]. Lagrange [YWJ^{+}25]. Laguerre [ADZ25]. Landau [ZYL25]. Lanzhou [WJBL25]. Laplacian [LYM25, MP25, ZYL25, ZW25]. large [LD25, LXM25, PFZ25]. large-scale [PFZ25]. lattice [YQYS25]. law [LTH^{+}25, MMNP25]. laws [CG25a, SH25, TNL25]. layer [BGQR25, YY26, YXG26, ZLJY25]. layered [ACCO25]. layers [GSB26]. leader [CWS^{+}25, CC26, LLZL26, SZLG26, YY26, YHSZ26, ZLH26]. leader-driven [CC26]. leader-follower [SZLG26, YY26, YHSZ26]. leader-following [CWS^{+}25, LLZL26]. Learning [VMM26, WWG^{+}25, CXZ26, EP25, GCW^{+}26, HCQ26, JYLS25, LHJ^{+}25, LHZL25, LMMG25, LF26, LW25b, MNE26, PDR25, PPN^{+}25a, PPN^{+}25b, VL26, WZWC25a, WZLS26, WS26, XS26, XZSC26, YGY^{+}25, YY25b, YZ25, ZZ25b, ZLH26]. learning-based [CXZ26, JYLS25, LHZL25, WZWC25a]. least [JZGV25, LXM25, LYM25, MNE26, TFCL25, ZHF25]. least-squares [TFCL25]. Legendre [KMK26]. Leslie [SA26]. Letnikov [LLMT25]. Level [IAK25, LLS^{+}26, MR25, SWW25]. levels [BHL25]. Leveraging [TS25]. Lévy [ZYL25]. Lexicographic [AS21, AS25]. Lie [AB25]. life [LZ25b]. like [CDD^{+}25, DEMFRL26, FLSS26, MM25, OCN^{+}26]. Lillo [BS25]. limited [ALNP25, LHZL26b, WL26, YQQ^{+}26]. limits [MMNP25]. line [WXD^{+}25]. Linear [BGQR25, CF25, LL26a, Abb26, BM26, CWS^{+}25, CMG25, CS25, CTTN25, JYLS25, JZ25, KK25a, KL25, KM25, LXM25, Li26, LZY^{+}25, LW25d, Mez25, MR25, PPH26, QHZ25, QXZ25, SS24, SS25, SSK^{+}26, Tan25, TO25, TNL25, Wu25, YHSZ26, ZZ25f, dOKA25]. linear-elastic [KM25]. linear-hyperbolic [CMG25]. link [LW25b, YHSZ26, ZXW^{+}25]. Liouville [MK25, ST25, ZQS26]. Lipschitz [BDD25]. liquidity [QZ25]. list [JLJZ26]. load [ZLCS26]. Local [CLWZ26, KGJ25, LTH^{+}25, ABdF^{+}25, FGJ25, GMOS26, Pal25, PSM25, ZG26a]. localized [TFCL25]. logarithmic [Hag26, YWJ^{+}25]. Logistic [DCC^{+}25]. logistics [ZFFM25]. Long [JT25, MCH25a, QHZ25]. long-range [QHZ25]. long-term [MCH25a]. loop [DZY^{+}25]. lose [XZSC26]. losses [ZZ26b]. Lotka [Che25]. Low [HYK^{+}25, LLLZ25, HJF25]. Low-rank [HYK^{+}25, LLLZ25]. LP [LZJ^{+}25]. LP-based [LZJ^{+}25]. LQ [HZ26a]. Lur’e [HWC25]. Lurie [WLZ25a]. Lyapunov [ABD25, MM25, WYKS25]. Lyapunov-like [MM25].

Ma [DS25]. machine [LMMG25, LW25a, MNE26]. macroscopic [BCCF25]. magnetic [PPS25, RRR19, RRR20]. magnetohydrodynamic [LD25]. Mahony [WMWY25]. main [GAYKS25]. maintain [TS25]. maintains [YYW^{+}25]. majority [Kam26]. making [YNZ25, ZJX25]. management [BPL^{+}25, ZZT25]. Maneuvering [ZFW25]. manifolds [CFS25]. manipulating [KMK26]. manipulator [YYMY26]. many [GAYKS25, ST25]. map [KC26]. mappings [HRS25]. March [Ano25h, Ano25t, Ano26b]. margin [VL26]. markets [ZCJL25]. Markov [BM25, CDYL25, IÖÜP25, JYLS25, LW25a, PPC^{+}25, SW25, SSK^{+}26, WWS^{+}25, XC25, XZSC26]. Markovian [LL25b, WLZ26, ZZ25c]. Mass [KBM25, MMNP25, PSS^{+}25, SFL26, ZC25, ZLH26]. mass-momentum-energy-preserving [SFL26]. masses [Chr26]. matchable [ZLY26]. matching [LMQZ25, LYZ^{+}26, ZHL25]. matchings [GK26, YYZG25]. materials [FGR25, WQLW25]. Math [AS25, PPN^{+}25a, SS25]. Mathematical [PPS25, RRR19, RRR20, TLSVGR25, BK25, Cro25, PNN25]. Mathematics [PPS25, RRR20, WMZ25a]. matrices [AMMT26, BDST25, BEG^{+}25, BFPS26, MWS25, MLZ26, SP25b]. Matrix [DCC^{+}25, Han26, WWLZ25, Dib26, HYK^{+}25, IÖÜP25, KCLZ25, KL25, Li26, MMF26, MW26, MS26, WSX^{+}25, ZMM26]. Matrix-based [Han26]. matrix-separation-based [WSX^{+}25]. matrix-weighted [MS26]. Maximal [WJBL25, YYZG25, BCCF25]. Maximally [CRHTV25]. maximizing [AD25]. maximum [HZ26b, SL26, XW25, YYZG25]. May [Ano25i, Ano25u]. McKean [LZ26, WXT25]. MDMDT [LZJ^{+}25]. mean [CK25, MMNP25, NNW25, XWW^{+}26]. mean-reverting [CK25]. mean-square [XWW^{+}26]. measurement [PFZ25]. measurements [WXD^{+}25]. measures [ABdF^{+}25, HCCH25]. mechanical [BGV26, LM25]. mechanism [CWS^{+}25, Che25, DWL^{+}25, LWL^{+}25, Mut26, WY26b, WWS^{+}25, YBSC25, YY25b, ZLCS26]. mechanisms [LF26, SP25a, SG26, ZZ25b]. media [LL25c]. medical [YK26]. medicine [DCC^{+}25]. meet [ZZ26b]. memetic [Pal25]. memory [HWC25, LHZL26a, PPH26, SNP25, YY25b, YWR^{+}25, ZWL25, ZN25]. memory-based [HWC25, ZWL25]. memristive [PPN^{+}25a, PPN^{+}25b, SZY25, SNP25, XWW^{+}26]. Mental [LL25d]. Merton [Ruj25]. mesh [JT25, RBT25]. meshless [BSM^{+}26, MCH25a]. metastatic [BDL25]. Method [TO25, BKE25, BMM25, BK25, CFS25, CCLL25, DN25, EP25, FLLR25, FGJ25, GZL^{+}26, Hag26, HCTS25, HAD25, HS25, IAK25, IÖÜP25, JY25a, JT25, JLZ25, KSR25, KMK26, LT25, LD25, LXM25, LL26a, MCH25a, ML26, PP25, PSM25, QXZ25, RB25, SSH25, TFCL25, TLW^{+}25, TKM25, WF25, WXD^{+}25, WQLW25, WZLS26, WLLL26b, WLLL26a, WH26, YY25a, ZT25, ZLJY25, ZZL^{+}25, ZZD26, ZLHS25, ZG26b]. methods [BSG^{+}25, BDD25, BM26, CCD26, CPR26, CRHTV25, GKM25, HJF25, JK25, JZ25, KGJ25, LCS^{+}26b, MMF26, ML26, NNW25, RQWX25, RPD25, SST26, Tan25, TZL26, TNL25, TLSVGR25, WZXY26, WY26a]. Metric [PJK25]. micro [WLZ26, YNZ25]. micro-dynamics [YNZ25]. micro-grids [WLZ26]. microgrids [LC25]. migration [BM25, SIL25]. mill [LDXL26]. mimetic [OGPFVV25]. MIMO [HLT25, LBLZ25]. Min [WMZ25a, WMZ25b]. minimization [WMZ25a, WMZ25b]. Minimizing [LYM25]. Minimum [CLLC25, BM26, WLLL26a, YL26, YWJ^{+}25]. minimum-time [BM26]. mining [TF25]. minor [LL25a, YZH25]. Mitigating [Kab25, DGD^{+}26]. mixed [BS25, HRS25, HJF25, ZHLL25]. mobile [RMS25, Tan25]. mobile/immobile [Tan25]. Mode [CDYL25, KNKR25, LSF^{+}25, LZJ^{+}25, LLLH25, SSK^{+}26, WZWC25a, ZHZ^{+}26]. Mode-based [CDYL25]. mode-dependent [LZJ^{+}25]. model [BCCF25, BGQR25, BK25, BGV26, CMG25, Che25, CK25, Cro25, DGD^{+}26, FY25, Hai26, KK25a, KNKR25, LHL^{+}26, LLS^{+}26, MZX25, MLTL25, PBS25, PNN25, RK25, RPS^{+}25, Ruj25, SKSC26, SA26, TGS25, WPTK26, XC25, XGL25, YH25, YQYS25, ZG26a, ZZD26, ZZTL25, ZZT25]. Modeling [FZY^{+}25, QWWW26, LW25a, PPS25, RRR19, RRR20]. Modelling [DAG25]. models [BFdOdS25, BDL25, Chr26, LYLS25, LM25, MCKZ25, RR26, TLSVGR25]. modification [RBT25, XSB25]. modified [KCLZ25, QHZ25]. modifying [KSFD25]. modulus [Mez25]. modulus-based [Mez25]. Moll [DS25]. mollifier [QX25b]. momentum [SFL26]. Monitoring [BFdOdS25, LWL^{+}25]. monkeypox [DGD^{+}26]. monotonicity [NNW25]. Moore [BFPS26, MWS25, ZMM26]. most [LSWZ25]. Mostar [BDTP25, BDTP23]. motion [ZHLL25]. motions [FGR25, WXT25]. moving [Chr26, LYD25]. MPC [SWW25]. Multi [ALV25, ZLSQ26, CWS^{+}25, CXZ26, KNKR25, LSP^{+}25, LHZL25, LHZL26a, LLZL26, LZY^{+}25, LZZ^{+}26, LYD25, MZX25, MS26, QMWY25, QLCW25, SDH25, SZLG26, TX25, WLZ26, XTGZ25, YGY^{+}25, YQQ^{+}26, YXG26, ZXS^{+}25, ZZS^{+}26, ZLL26, ZZ25e, ZZ26b]. multi-agent [CWS^{+}25, CXZ26, LSP^{+}25, LHZL25, LHZL26a, LLZL26, LZY^{+}25, LZZ^{+}26, MS26, QMWY25, SDH25, SZLG26, WLZ26, XTGZ25, YGY^{+}25, YQQ^{+}26, ZXS^{+}25, ZZS^{+}26, ZLL26, ZZ25e, ZZ26b]. multi-information [MZX25]. multi-issue [QLCW25]. multi-layer [YXG26]. Multi-pursuer [ZLSQ26]. multi-sensor [LYD25]. Multi-state [ALV25]. multi-term [TX25]. multi-UAV [KNKR25]. multiagent [GPL25, HY25, LZZ25, YHSZ26, ZXW^{+}25]. multiband [LZD^{+}25]. multigame [AZZ^{+}26]. Multilayer [SG26]. multiplayer [CH26, HCQ26]. Multiple [ZN25, CH26, CRS25, LWY25, RQWX25, TLW^{+}25, WSZM25, WH26, XZN^{+}25, ZC25]. multiple-fractional [LWY25]. multiplex [JCW26, LH25, LGW^{+}26, SLX^{+}26, WWY^{+}25, YH25]. multiresolution [CRS25]. Multiscale [MBF25]. multispace [Kov25]. multispaces [Kov25]. Multistability [ZLW^{+}26, SP25a]. multistate [BQS25]. multistep [YYT25]. multivariate [MCHB25]. Müntz [WX25]. music [MBH^{+}25]. Mutual [CDD^{+}25, RKS25, KV25]. Mutual-visibility [RKS25, KV25]. Mycielskians [RKS25].

Nash [BAG^{+}25]. Natural [AB25, BD25, ZZJZ25]. Navier [JH25, JLZ25, NS25]. negative [BDST25, LGW^{+}26]. Neighbor [FT25, ZZ26a]. neighbors [ZZTL25]. nested [KL25]. Network [DCC^{+}25, CH26, GLY26, HC26, KC26, LZZ25, LLMR25, LH25, LD26, NKSB25, PPN^{+}25a, PPN^{+}25b, QTW25, WLW26, YYMY26]. network-based [GLY26, YYMY26]. networked [CZGY25, JXZ^{+}25, LLWC26, Mut26, XC25, YZ25, ZD25]. networks [AAS25, ALV25, CGS^{+}26, CDYL25, CC26, CLWZ26, FT25, HCCH25, JCW26, LCL25, LHZL26b, LZY^{+}25, LGW^{+}26, LYD25, MS26, MeFW25, MMNP25, SS24, SS25, SDH25, SZY25, SW25, SLX^{+}26, SG26, SHY25, SNP25, WWY^{+}25, WHCB25, WWG^{+}25, WSX^{+}25, WSZM25, WZWC25b, XWW^{+}26, YBSC25, YFZ^{+}25, YWR^{+}25, YY26, YH25, YXG26, ZLW25, ZN25, ZBM25, ZZL^{+}25, ZZ26a, ZLW^{+}26, ZXS^{+}25]. Neural [NKSB25, SS24, SS25, AAS25, GLY26, LZZ25, LD26, MLW^{+}25, SZY25, SW25, SNP25, WSX^{+}25, WWQX25, XWW^{+}26, YBSC25, YWR^{+}25, YYMY26, ZN25, ZZL^{+}25, ZLW^{+}26]. Neuro [ZHWW26]. Neuro-adaptive [ZHWW26]. neuron [PPN^{+}25a, PPN^{+}25b]. neurophysiological [ABdF^{+}25]. neutral [HLZ25, JZ25, MMF26, WXT25, WLLL26b, YYW^{+}25]. neutral-type [HLZ25]. Newton [BGI25, CRHTV25, LT25, YY25a]. Newton-type [CRHTV25]. Newtonian [PPS25, RRR20, RRR19]. no [WWK25]. noise [YFZ^{+}25, ZYL25, ZFW25, ZD25]. Non [LSWZ25, SNP25, BDST25, BDD25, CTTN25, DEMFRL26, HLZ25, JSZ25, KH26, LJ25, PPS25, PMVA26, PPH26, RRR19, RRR20, SSK^{+}26, Tan25, WZLS26, ZLJY25, ZHWW26, dOKA25]. non-affine [ZHWW26]. non-autonomous [KH26]. Non-cospectral [LSWZ25]. non-differentiable [HLZ25]. non-existence [PMVA26]. Non-fragile [SNP25, ZLJY25]. non-global [BDD25]. non-homogeneous [LJ25, PPH26]. non-linear [CTTN25, SSK^{+}26, Tan25, dOKA25]. non-negative [BDST25]. non-Newtonian [PPS25, RRR20, RRR19]. non-standard [DEMFRL26]. non-strict [WZLS26]. non-uniform [JSZ25, Tan25]. nonexponentially [WPTK26]. nonhomogeneous [JZ25, Ruj25]. Nonlinear [CFS25, LBLZ25, TXJW26, Abb26, APJ25, ALC^{+}25, BAT25, BKE25, BGQR25, BMM25, CL25, CYZ26, CS25, CRHTV25, CXZ26, DZML25, EST^{+}25, FLLR25, GX25, GCW^{+}26, GLY26, GZL^{+}26, HCTS25, HLT25, HCQ26, IJM25, JY25a, KSR25, LZZ25, LWS^{+}25, LZP25, LTH^{+}25, MLW^{+}25, MBH^{+}25, NSX^{+}25, PFZ25, PSM25, QX25a, QMWY25, RS25, RHSZ26, Ruj25, SZM25, SWW25, SWXL25, TGS25, TX25, WXT25, WXD^{+}25, WMWY25, WGeL25, WL26, WZXY26, WZLS26, WLLL26b, WY26a, WLLL26a, WYKS25, XC25, YBSC25, YYMY26, ZWQC25, ZLA25, ZFW25, ZC25, ZXW^{+}25, ZZS^{+}26, ZLH26]. nonlinearities [CYZ26]. nonlinearity [HLZ25, LTH^{+}25]. nonlinearly [LSH25, SWLC25]. nonlocal [CjZjHbX25]. nonnegative [SP25b]. nonperiodic [OGPFVV25]. nonsmooth [AS21, AS25]. nonstandard [CPR26, Li26, WPTK26]. nonuniform [JT25, XGL25, ZT25]. norm [IÖÜP25, JT25, LCS26a, Liu25]. norm-based [IÖÜP25]. normal [AAV25, AS21, AS25, LHLL25, LBLZ25, MLZ26]. Note [LMQZ25, FGW25, FGJ25, JSZ25, LZCD25]. nothing [NS25]. novel [KMK26, MCH25a, TFCL25, WPTK26, YWR^{+}25]. November [Ano25j, Ano25v]. NSDDEs [NNW25]. Null [BEtM25]. number [AD25, ADPS25, ET26, GK26, HZ26b, KSSY26, LSQ25, LMQZ25, LX25, LLX25, LZ25a, LLW26, ZZTL25]. numbers [GVW25, KK25b, LD25]. Numerical [ALC^{+}25, BM25, BDL25, BM26, HAD25, JK25, LYQZ26, LM25, Som25, TX25, WWQX25, ABD25, Chr26, GSB26, NS25, TLSVGR25, WPTK26, ZYZ^{+}25]. numerics [CCC^{+}26]. Nyström [BMM25, KSR25]. Nyström-type [KSR25].

objective [YH25]. objects [YY25b]. observable [LBLZ25]. observation [LZ26, Ruj25]. Observer [LZZ^{+}26, PPC^{+}25, SWW25, WZWC25b, XC25, YHSZ26, ZLJY25, EST^{+}25, GCW^{+}26, LHZL26b, LWS^{+}25, SZM25, YYMY26, ZXS^{+}25, ZZ25f]. Observer-based [LZZ^{+}26, PPC^{+}25, SWW25, WZWC25b, XC25, YHSZ26, ZLJY25, EST^{+}25, LWS^{+}25, SZM25, ZXS^{+}25]. observer-driven [LHZL26b]. obstacle [FA26]. ocean [BGQR25, MCKZ25]. October [Ano25k, Ano25w, WMZ25a]. Octonionic [RZ25]. odd [XW25]. ODE [BEtM25]. ODE-heat [BEtM25]. ODEs [FGJ25]. off [LCW25]. on-line [WXD^{+}25]. on/off [LCW25]. One [JS25, EP25, FGR25, WMW^{+}25, XSB25]. one-dimensional [EP25, WMW^{+}25]. One-to-one [JS25]. OP [LZD^{+}25]. OP-HHO [LZD^{+}25]. open [BPL^{+}25]. operation [TQWW26]. operations [GXY25]. operator [IJM25, Liu25, WGeL25]. operators [ADR^{+}26, MK25, NKSB25, OGPFVV25, ST25]. opinion [QTW25]. opinion-dependent [QTW25]. Optimal [AJK25, DZML25, RS25, SLX^{+}26, WLZ26, ZCJL25, ZD25, CTTN25, DMT25, GPL25, JDES25, MCHB25, MNE26, RHSZ26, WS26]. optimization [CFS25, GX25, LCLY25, XTGZ25, YY25a]. optimized [GCW^{+}26, YGY^{+}25, ZHWW26]. Optimizing [GKR^{+}25]. option [Hai26]. options [AJK25, Grz26, Ruj25, SL26, XZN^{+}25, ZG26a]. orbit [LWML26]. order [ALNP25, AA25, CG25a, CLLC25, CPR26, CTTN25, Dib26, FY25, JT25, KH26, KK25b, KGJ25, LT25, LZZ25, LH25, LZ25a, LDT^{+}25, MS26, MR25, NVGA^{+}26, PP25, PMVA26, PPH26, QXZ25, SZY25, SZM25, SLX^{+}26, SG26, SWLC25, Tan25, WXD^{+}25, WMWY25, WGeL25, WZXY26, WLLL26a, WS26, YHGZ25, YWR^{+}25, ZWQC25, ZLA25, ZYZ^{+}25, ZZD26, ZLW^{+}26, ZZS^{+}26, ZLL26]. ordering [Pal25]. ordinary [BBG^{+}26]. oriented [FY25]. orthogonal [LCLY25]. oscillators [FZY^{+}25]. oscillatory [JK25, KXL^{+}25, KSR25]. other [CC26]. other-regarding [CC26]. Out-of-equilibrium [CA25]. outbreaks [DGD^{+}26]. outerplanar [LLX25]. output [BAT25, CGS^{+}26, CL25, GX25, LDXL26, LBLZ25, MLW^{+}25, MCH25b, Mut26, PFZ25, SNP25, WZLS26, YZ25]. output-feedback [BAT25, CL25, MLW^{+}25, SNP25]. overshooting [APJ25].

packet [JXZ^{+}25, ZZ26b]. packet-losses [ZZ26b]. packing [BFH^{+}25]. paired [FP25]. pancake [FT25]. pancreatic [AS21, AS25]. pancyclicity [LCL25]. pantograph [JZ25]. paper [BFdOdS25, LDT^{+}25]. parabolic [ALNP25, DN25, EST^{+}25, MR25]. parallel [GZL^{+}26, Pal25]. Parameter [WXD^{+}25, ABD25, BBG^{+}26, JH25, PDR25, VMM26]. parameter-dependent [ABD25]. parameterized [KL25, LSH25, SWLC25]. parameters [LMMG25, XZN^{+}25]. Parking [Seb25]. partial [AZD25, AMMT26, HS25, LHZL26b, LD26, SS24, SS25]. partially [JYLS25, LL25b, NSX^{+}25]. particle [PSS^{+}25, RPD25]. partition [AdFS26]. Partitioning [BBV25]. passive [SW25]. passivity [SHY25]. Past [LW25c, TS25]. Past-performance-driven [LW25c]. path [JS25, QM25, WLZ25b]. paths [LLX25]. Pattern [Che25, YZM25]. patterns [DEMFRL26]. Pauli [CDM25]. payoff [LL25d]. payoffs [Ruj25]. PDE [EST^{+}25, PMVA26]. PDEs [AAS25]. PDT [ZLJY25]. pedestrian [BCCF25]. peers [LdMAS25]. pencil [MMF26]. pendant [AD25]. Penrose [BFPS26, MWS25, ZMM26]. perception [AZZ^{+}26, CH26, LHJ^{+}25]. Perfect [ZHL25]. performance [CXZ26, HWC25, HCQ26, LZD^{+}25, LSH25, LW25c, SZSW25, SZLG26, TXJW26, ZD25, ZZ25d]. period [XTGZ25]. periodic [OGPFVV25, PPS25, RRR19, RRR20, WMW^{+}25]. permeability [HJF25]. Perron [SP25b]. persistence [LDT^{+}25]. perspective [LL25d, MZX25, SKSC26, ZFFM25]. perturbative [ADR^{+}26]. perturbed [GSB26, MM25, RBT25, SSK^{+}26, TXJW26]. perverse [TS25]. pest [ZZT25, ZZJZ25]. pest-natural [ZZJZ25]. phase [KM25, QHZ25, SP25a]. phenomenon [AA25]. physical [DZY^{+}25]. Picard [ALNP25, ADR^{+}26]. Picard-CR [ADR^{+}26]. piece [LZJ^{+}25]. piece-wise [LZJ^{+}25]. piecewise [BM25, LLQ25, LW25d, PPH26]. piezoelectric [Wan25]. Pinning [SHY25, WZWC25b]. Planar [WWK25, PC25, TQWW26, WWL25]. Plane [MH25]. planning [RMS25]. plans [ZCJL25]. platform [ZFFM25]. platoon [LCRZ25]. player [SGH^{+}25, WWLZ25]. PMSM [RK25]. point [AK26, ADZ25, BSM^{+}26, GCV25, HCCH25, JDES25, Wu25]. points [CLLC25, HHSW25, KMK26]. Poisson [Ruj25]. polar [Gao25]. policy [WWY^{+}25]. polygon [KMK26]. polygons [Som25]. polymeric [MBF25]. polynomial [ECD^{+}25, KMK26, PP25, RB25, RNL25]. polynomials [DS25, RB25, RNL25, WX25]. polyomino [ZW25]. pool [GPH25]. population [CMG25, CA25]. populations [LM25, PNN25, QLCW25, YYW^{+}25]. poroelastic [KBM25]. poroelasticity [HJF25]. porous [RR26]. port [BRZS25]. port-Hamiltonian [BRZS25]. portfolio [AJK25, RS25]. posed [Li26]. position [ADPS25, RKS25]. positive [FWLZ25, LZJ^{+}25, NSX^{+}25, SP25b]. positivity [Hag26, QHZ25]. positivity-preserving [QHZ25]. potential [MK25]. potted [HZ26b]. poverty [DAG25]. power [LTH^{+}25, SG26, YXG26, ZLCS26]. Practical [HWC25]. preconditioner [FLSS26]. preconditioners [AK26, Wu25]. predator [MMTU25, MLTL25, SA26, TGS25]. predator-prey [SA26, TGS25]. predators [ZWL25]. Predefined [WZLS26, WYKS25, YBSC25, ZLCL25]. Predefined-time [WZLS26, WYKS25, ZLCL25]. predictive [KNKR25]. Predictor [YZ25, IJM25]. Predictor-based [YZ25]. preference [LF26]. preferences [CC26, QLCW25]. preferential [LWH^{+}25]. Prescribed [LZY^{+}25, RK25, ZLL26, ZLH26, CL25, HCQ26, LLZL26, MM25, TXJW26, WJBL25]. Prescribed-time [LZY^{+}25, RK25, ZLH26, CL25, LLZL26, MM25]. presence [BAT25]. preserving [AKM25, CLLF25, Hag26, LLQ25, LL25c, LTH^{+}25, QHZ25, SFL26, WGeL25, ZLSQ26]. prevalence [DAG25]. prey [Kab25, MMTU25, MLTL25, SA26, TGS25, ZZJZ25]. prey-predator [MMTU25, MLTL25]. pricing [CK25, Grz26, Hai26, Ruj25, SL26]. principle [RZ25]. principles [CDM25]. prior [LLQ25]. prism [SL26]. prisoner [BHL25, LLMR25, LWH^{+}25, LW25b, LW25c, TS25, Tak26, YYW^{+}25, KH25]. privacy [HY25, QWWW26, ZLSQ26]. privacy-consensus [HY25]. privacy-preserving [ZLSQ26]. Probabilistic [LW25a, WSY25]. probabilities [PPC^{+}25, WWG^{+}25]. Probability [SL26, HHSW25, KK25a, LL25b, MCH25b]. problem [BK25, BM26, Cam26, CjZjHbX25, GED26, GMOS26, JZGV25, JDES25, LXM25, LL25c, MCH25a, Pal25, PBS25, TLSVGR25, WMZ25a, WMZ25b, XSB25, YWJ^{+}25, ZG26b]. problems [ACCO25, AK26, BSM^{+}26, BGI25, CWWT25, CCC^{+}26, DN25, GZL^{+}26, HRS25, HZ26a, ISC25, IAK25, KV25, LJ25, Li26, Mez25, RBT25, ST25, TFCL25, YY25a, ZYZ^{+}25, ZQS26, dOKA25]. procedure [ACC25, BGI25]. process [Hai26, Ruj25]. processes [BM25, Grz26]. product [Li26, LWML26]. production [RMS25]. programming [CF25]. project [TZL26]. projection [GKM25]. projects [ZHLL25]. prolong [XTGZ25]. promote [LF26, LW25b, LW25c]. promotes [DWL^{+}25, LLMR25, LWH^{+}25]. promoting [LSF^{+}25, YLR^{+}25]. promotion [LL25d]. propagation [GSB26]. Proper [WWL25]. properties [BEG^{+}25, BGV26, SSH25]. property [KMK26, XLLC25]. prostitution [DAG25]. protection [HY25]. Protocol [LL25b, CZGY25, CDYL25, JCW26, ZXS^{+}25, ZZS^{+}26, ZHZ^{+}26]. Protocol-based [LL25b]. proximal [HZF^{+}25]. pseudo [Abb26, XSB25, YYT25]. pseudo-Jacobi [XSB25]. pseudo-spectral [Abb26, YYT25]. public [CC26, GSL25, KH25, LCS26a, LHJ^{+}25, LLFS25, SIL25, WSY25, ZZZM25, Zha25, ZXP^{+}25]. PUM [CCD26]. punishment [GSL25, WSY25, ZZZM25]. punishment-reputation [ZZZM25]. pursuer [ZLSQ26]. PWA [XZSC26]. Python [ECD^{+}25].

Q [LHJ^{+}25]. Q-learning [LHJ^{+}25]. Qi [DS25]. Qt [LWML26]. Qt-product [LWML26]. quadcopter [TXJW26]. quadrant [CS25]. quadratic [CC25, JYLS25]. quadriculated [LWLZ26]. Qualitative [ZZT25]. quality [FWLZ25]. Quantifying [ACCO25]. quantization [CL25, LYD25, MCHB25]. Quantized [LDXL26, HWC25, WZWC25b]. Quantum [Yu25, Abb26, SKM26]. quarantine [FBK25]. Quasi [JCW26, LSP^{+}25, BGQR25, MW26, MR25]. quasi-geostrophic [BGQR25]. quasi-linear [MR25]. Quasi-synchronization [JCW26, LSP^{+}25]. quasi-Toeplitz [MW26]. Quaternion [KCLZ25, HYK^{+}25, LWML26]. quaternionic [JZGV25]. QUAV [SZSW25, ZLCL25]. queueing [ZXP^{+}25].

R&D [ZHLL25]. Radial [HS25, LD26, PMVA26, TNL25, WQLW25]. Radio [BVZ25]. radius [AdFS26, LL25a, ZZ25a]. Radon [MH25]. Rainbow [WJ25]. Ramsey [LMQZ25]. random [LSF^{+}25, MCHB25, QZ25, SW25]. randomization [Grz26]. randomized [LWML26]. range [QHZ25]. rank [HYK^{+}25, LLLZ25, XSB25]. rank-one [XSB25]. rate [WXT25]. rates [CA25]. rational [BBG^{+}26, LL26a]. RBF [SH25, VMM26]. RBF-WENO [SH25]. RBFNN [LHZL26a]. Reachable [SZY25, LZZ^{+}26]. reaction [JCHW25, LWY25, PSS^{+}25]. reaction-subdiffusion [JCHW25]. reactions [ADR^{+}26]. reactive [BFdOdS25]. realizations [AB25]. rearrangement [ZG26a]. reciprocity [LLMR25, QFNW25]. reconstructing [ALNP25]. Reconstruction [YNZ25]. recurrence [WX25]. recursion [ZZL^{+}25]. Recursive [RNL25, ZBM25]. Redefined [PP25]. Reduced [SZM25, FY25, KC26, LLLZ25, MP26]. Reduced-order [SZM25, FY25]. reduces [LCS26a]. reduction [COP^{+}25, GLD25, MeFW25]. reductions [BDST25]. redundant [YT25]. refueling [QAS26]. regarding [CC26]. region [WWS^{+}25]. regression [RNL25]. regressions [Hai26]. regular [ZHF25]. regularity [Tan25]. regularization [PDR25, YY25a]. Regularized [NS25, WMZ25a, WMZ25b, YY25a, YQYS25]. regularizer [HYK^{+}25]. regulation [ABC^{+}26, QWWW26, ZD25]. Reinforcement [JYLS25, LHZL25, XS26, CXZ26, GCW^{+}26, HCQ26, MZX25, WZLS26, WS26, YGY^{+}25, ZLH26]. rejection [ZZ25d]. relative [SL26]. relaxation [CG25a, KM25, YQYS25]. Relaxed [KL25]. Reliability [ZBM25, BQS25]. Remaining [LZ25b].

Rényi [MMNP25]. repeated [QLCW25, Ued25, YNZ25]. replanning [KNKR25]. replenishment [BK25]. representation [CRS25, HZ26a]. Representations [ZMM26]. Reproducing [QXZ25]. reputation [BHL25, DWL^{+}25, KH25, LF26, WL25, WY26b, YY25b, ZZZM25, ZZ25b, ZXP^{+}25]. reputation-based [DWL^{+}25, LF26]. reputational [WSY25]. Research [LWL^{+}25, WY26b, LZD^{+}25]. Reset [ZXS^{+}25]. residual [WLLL26a]. resilience [WY26b]. Resilient [EST^{+}25, LC25]. resistance [JY25b]. resolution [BS25]. Resolving [WMW^{+}25]. resource [CA25, GPH25, LZY^{+}25]. respiratory [FBK25]. response [LLWC26, NVGA^{+}26, ZLHS25]. responses [AS21, AS25]. restrained [GXY25]. restricted [YZM25]. restriction [ZBM25]. results [AMMT26, GKM25, KH26, LL25a, WHCB25, ZYY26, ZG26b]. retirement [AJK25]. retraction [CFS25]. reversibility [WMW^{+}25]. reversible [ADR^{+}26]. reversing [SZLG26]. reverting [CK25]. review [XS26]. reward [DWL^{+}25, IÖÜP25]. rewards [GLHS25, Wet25]. rhythmic [MBH^{+}25]. riders [Kis26]. Riemann [PBS25]. Riemannian [CFS25]. Riesz [CYZ26, TX25]. right [CF25]. right-hand [CF25]. ring [LLLZ25]. risk [CH26, RS25, SL26]. robin [CZGY25, ZHZ^{+}26]. robotic [FA26]. Robust [BBG^{+}26, GSB26, RPS^{+}25, VMM26, ZT25]. rock [BFdOdS25, LDT^{+}25]. rock-paper-scissors [BFdOdS25]. rod [MBF25]. Roesser [WWS^{+}25]. role [YLR^{+}25]. Roman [GXY25, GVW25]. root [BDTP23, BDTP25]. root-indices [BDTP23, BDTP25]. rotation [CC25, ZHL25]. round [CZGY25, ZHZ^{+}26]. round-robin [CZGY25, ZHZ^{+}26]. Routes [SKSC26]. row [Pal25]. rule [WSY25]. rules [LW25b, PDR25]. rumor [LHL^{+}26]. Runge [ML26, WY26a].

S [ZLJY25, ZLCS26, ZZ25d, ZHZ^{+}26]. SP [Grz26]. SABR [CK25]. saddle [AK26, Wu25]. salesman [BK25]. same [AdFS26]. sample [ZG26a]. Sampled [WLZ25a, CLWZ26, MLW^{+}25, WDMM25, WSX^{+}25, YFZ^{+}25]. sampled-data [CLWZ26, MLW^{+}25, WDMM25, WSX^{+}25, YFZ^{+}25]. Sampled-data-based [WLZ25a]. Sampling [DEMFRL26, LZZ^{+}26, QX25a, XTGZ25]. sampling-based [LZZ^{+}26, QX25a]. sanctioning [LCS26a]. satellite [LCW25]. saturating [CZGY25, LX25]. saturation [LLZL26, QX25a, SDH25, WZLS26]. scalar [CG25a, TNL25]. scale [MAMFBM^{+}25, PFZ25]. scaling [ZXW^{+}25]. scattering [ACCO25]. scheduling [JXZ^{+}25, RMS25]. scheme [AZD25, AKM25, BGV26, CTTN25, JCHW25, LLMT25, MR25, PDR25, SH25, WGeL25, WPTK26, XGL25, ZT25, ZYZ^{+}25, ZZS^{+}26]. schemes [CLLF25, CG25a, HCTS25, PSS^{+}25, QHZ25, SFL26, WMWY25]. Scholes [XGL25]. schools [Cro25]. Schrödinger [CLLF25, CYZ26, LTH^{+}25, WGeL25]. Schur [MMF26, ZMM26]. scissors [BFdOdS25, LDT^{+}25]. SCP [XC25]. SDEs [Hag26, WXT25]. search [Pal25]. secant [PSM25]. SECIR [WPTK26]. Second [CG25a, Tan25, Dib26, FLLR25, KH26, LT25, LZZ25, SZY25, SWW25, WZXY26]. second-derivative-free [LT25]. second-kind [FLLR25]. second-level [SWW25]. Second-order [CG25a, Tan25, Dib26, KH26, LZZ25, SZY25, WZXY26]. secondary [LC25]. Secure [BAT25, LHZL26b, RK25, ZC25, ZXW^{+}25]. security [XC25, ZKZ26]. seeking [Zha25]. seemingly [FGR25]. SEIAR [LHL^{+}26]. selection [LWH^{+}25, LZD^{+}25, LF26]. Self [FA26, LYZ^{+}26, ZXW^{+}25]. self-adaptive [LYZ^{+}26]. Self-triggered [FA26, ZXW^{+}25]. Semi [WLLL26a, CDYL25, LW25a, MW26, SW25, SSK^{+}26, TLW^{+}25, XC25, ZZ25c]. Semi-analytical [WLLL26a]. semi-finite [TLW^{+}25]. semi-infinite [MW26]. semi-Markov [CDYL25, LW25a, SW25, SSK^{+}26, XC25]. semi-Markovian [ZZ25c]. semilinear [ZT25]. semilocal [KGJ25]. sensitive [LCS26a]. sensitivity [ML26, NVGA^{+}26, PFZ25]. sensor [BAT25, GLY26, LYD25, SWLC25]. separable [WXD^{+}25]. separation [PC25, WSX^{+}25]. September [Ano25l, Ano25x]. sequence [GX25, YL26]. sequences [FGW25]. sequential [ALV25, ZXW^{+}25]. series [Han26, Mao25]. service [ZFFM25]. Set [MeFW25, ACCO25, KGJ25, LZZ^{+}26, SZY25]. sets [BFH^{+}25, HZ26b]. setting [JDES25]. shape [RPS^{+}25, VMM26]. shared [Wet25, ZFFM25]. shift [AK26]. shift-splitting [AK26]. shocks [WY26b]. shooting [BGI25]. shooting-Newton [BGI25]. short [PPH26, WWL25, YWR^{+}25, ZN25]. short-memory [PPH26]. shortest [DMT25]. Showcase [BAG^{+}25]. Showdown [BAG^{+}25]. side [CF25]. Sierpiński [GCV25, PJK25]. signal [CCD26, LLMT25]. Signed [GVW25, LYM25, SHY25]. signless [ZW25]. SIMD [ML26]. similarity [HCCH25, DCC^{+}25]. simplified [YGY^{+}25]. simulation [CLLF25, HRS25, HAD25, PSS^{+}25, WWQX25]. simulations [AT25, MBF25]. simultaneous [LCLY25, QX25b]. Sine [CYZ26, FLSS26]. Sine-transform-based [CYZ26, FLSS26]. single [Dib26, FZY^{+}25, MCHB25, ZLSQ26]. single- [MCHB25]. single-evader [ZLSQ26]. single-transistor [FZY^{+}25]. singular [AZD25, JYLS25, KXL^{+}25, LLLH25, LWML26, PTB25, SZM25, ZL25, ZKZ26]. singularities [JK25]. Singularity [SZSW25, GSB26, QX25a, ZT25]. Singularity-free [SZSW25, QX25a]. singularly [GSB26, RBT25, SSK^{+}26]. sixth [WGeL25]. sixth-order [WGeL25]. size [Wet25, XW25]. sketch [TZL26]. sketch-and-project [TZL26]. skew [ZG26a]. Sliding [ZHZ^{+}26, KNKR25, LLLH25, SSK^{+}26, WZWC25a, ZZT25]. slip [JH25]. small [LLQ25]. smart [ZLHS25]. Smith [AAV25, MLZ26]. smooth [YYMY26]. smooth-switching [YYMY26]. smoothed [RPD25]. smoothing [LLQ25, PSM25, THPA25]. snapshot [HCCH25]. snarks [ZHL25]. snowdrift [LLS^{+}26, Tak25]. social [FWLZ25, LYZ^{+}26, LdMAS25, MZX25, QTW25, QLCW25, WL25, YNZ25, ZWCT26]. sojourn [MCH25b]. solid [KM25]. solitary [CLLF25, FP25]. Soltés’ [Cam26, GED26]. Solution [Dib26, AD25, BDL25, BK25, Chr26, IÖÜP25, RMS25, TX25, WLLL26a]. solutions [CLLF25, CS25, EWH25, MBF25, PBS25, PMVA26, PTB25, SKM26, Tak26, TO25, ZL25, ZG26b]. solve [BKE25, CRHTV25, HS25, JT25, KCLZ25, KH25, WL25]. solver [BSM^{+}26, CWWT25, WWQX25]. solvers [CYZ26, KBM25]. Solving [Abb26, OGPFVV25, PPH26, AAS25, BGI25, BM26, CTTN25, EP25, HZF^{+}25, LXM25, LL26a, LD26, MCH25a, PP25, SST26, TZL26, WMWY25, YYT25]. Sombor [AD25]. Some [AMMT26, KK25b]. sort [DFKY26]. SOS [ZC25]. SOS-based [ZC25]. source [DN25, ISC25, QX25b, SG26, dOKA25]. space [CLLF25, DN25, ISC25, JLZ25, QX25b, QXZ25]. space-dependent [ISC25]. space-fractional [JLZ25, QX25b]. spacecraft [WS26]. spaces [SKM26]. span [DMT25]. spanning [ZLY26]. Sparse [CRS25, HYK^{+}25, LCLY25, TO25]. sparsity [LSF^{+}25]. Spatial [ZXP^{+}25, BHL25, GSL25, LCS26a, LWH^{+}25, LWY25, LYZ^{+}26, LW25b, LW25c, MLTL25, SIL25, ZZZM25, Zha25]. spatially [YQYS25, ZWL25]. spatio [WQLW25]. spatio-temporal [WQLW25]. special [FHX26]. spectra [AAV25, FHX26, ZW25]. Spectral [BEG^{+}25, LL25a, Abb26, AdFS26, GKM25, JLZ25, LZ25a, MLZ26, NVGA^{+}26, ST25, TKM25, YYT25, ZZ25a]. spherical [Som25]. spiking [ZLW^{+}26]. spins [BAG^{+}25]. spline [TX25]. splines [BD25, PP25, THPA25]. split [LCL25, WJ25]. split-star [LCL25]. splitting [AK26, LXM25, WLLL26b]. spreading [LHL^{+}26, SLX^{+}26]. square [NNW25, XWW^{+}26, YZH25]. squares [JZGV25, LXM25, MNE26, PJK25, TFCL25]. Stability [LZ26, NSX^{+}25, TNL25, WLLL26b, WY26a, CCLL25, EWH25, JXZ^{+}25, JT25, LZP25, LW25a, MMF26, MM25, MCKZ25, NNW25, SS24, SS25, WLZ25a, WYKS25, ZN25]. Stabilization [HLZ25, YFZ^{+}25, APJ25, CLWZ26, DZY^{+}25, LWY25, LW25a, MLW^{+}25, MeFW25, Mut26, SWXL25, WDMM25, WLZ25a]. stabilized [HJF25, LD25]. Stable [LJ25, AKM25, BGV26, WGeL25]. Stackelberg [DZML25]. stag [Tak25, Tak26]. stage [GZL^{+}26, PNN25, ZZJZ25]. stage-structured [PNN25]. standard [DEMFRL26]. star [DFKY26, LCL25, QM25, Sta25, ZZ26a]. star-free [Sta25]. state [APJ25, ALV25, CCLL25, EST^{+}25, GPL25, HZ26a, LWS^{+}25, LZP25, LW25d, SWXL25, ZYY26, ZHWW26]. state-constrained [GPL25]. state-delay [LW25d]. state-dependent [CCLL25]. state-overshooting [APJ25]. states [ABdF^{+}25, LHL^{+}26, SS24, SS25]. static [GX25]. stay [WPTK26]. steady [SS24, SS25]. stenosis [PPS25, RRR19, RRR20]. step [PPC^{+}25, ZZL^{+}25, ZLCS26]. stepping [SFL26]. Stewartson [HAD25]. stiff [FGJ25, WLLL26b]. stimuli [SP25a]. Stochastic [JXZ^{+}25, BDD25, CA25, CLWZ26, CK25, DAG25, HLZ25, LHJ^{+}25, LSP^{+}25, LYQZ26, LSH25, LZ26, MCH25b, MMTU25, NNW25, QZ25, RK25, TGS25, WL26, ZYL25, ZZ25c, ZFW25, ZC25, ZHWW26]. stock [ZCJL25]. Stokes [AA25, JLZ25, NS25, RR26]. storage [ZLCS26]. Störmer [WZXY26]. Störmer-Cowell [WZXY26]. strain [LGW^{+}26]. strategies [TS25, Ued25, ZCJL25]. strategy [AT25, FWLZ25, HWC25, LT25, LF26, LZZ^{+}26, LW25c, QFNW25, QLCW25, SLX^{+}26, Tak25, Tak26, WZWC25b, WYKS25, ZFFM25, ZJX25, ZXS^{+}25, ZZT25]. strength [FGW25]. strict [CL25, GCW^{+}26, GLY26, WZLS26]. strict-feedback [CL25, GCW^{+}26, GLY26]. strip [LDXL26]. strongly [ZHF25]. structural [FHX26, HC26]. Structure [CLLF25, KBM25, LTH^{+}25, WGeL25, YZM25, ZZJZ25, ZG26b]. Structure-preserving [CLLF25, LTH^{+}25, WGeL25]. structured [PSM25, PNN25, QLCW25]. structures [LLQ25]. studies [DCC^{+}25]. Study [lCXZ^{+}25, KC26, MZX25, NS25]. studying [MMF26]. Sturm [MK25, ST25, ZQS26]. sub [ZHLL25]. sub-fractional [ZHLL25]. subdiffusion [JCHW25, JT25]. subgraph [XW25]. subgraphs [BBV25]. subject [CZGY25, LYD25, QX25a, WZLS26]. subjected [CGS^{+}26]. subjective [YH25]. suboptimal [JT25]. subpath [KSSY26]. subspace [CCC^{+}26, LWML26]. subspace-orbit [LWML26]. substructure [YZM25]. Successive [BKE25]. successively [GK26]. sum [GPL25, YL26]. sum-connectivity [YL26]. summations [KL25]. Super [ZHY26]. supercloseness [XLLC25]. Superconvergence [GKM25, WF25, XLLC25]. Superconvergent [SST26]. superlinear [ZYL25]. Supervised [LLFS25]. supervisory [GLY26]. supply [YWJ^{+}25, ZFFM25]. support [MNE26]. sure [CMG25]. surface [TFCL25, YYMY26]. surrounding [TS25]. SVD [HYK^{+}25, Li26]. swaps [Ruj25]. switched [APJ25, CDYL25, CCLL25, HLT25, LZJ^{+}25, LSH25, LWS^{+}25, LZP25, MS26, NSX^{+}25, SZM25, WZP^{+}25, YBSC25, ZLJY25]. Switching [XWW^{+}26, AJK25, GX25, MCH25b, SDH25, WZWC25a, WLZ26, WZP^{+}25, YHGZ25, YYMY26, ZLJY25]. Sylvester [KCLZ25]. Sylvester-type [KCLZ25]. symbol [GKR^{+}25]. symmetric [CC25, MLZ26, WWLZ25]. Symmetrised [LLMT25]. Symmetry [BS25, SSH25]. Synchronization [PPN^{+}25a, PPN^{+}25b, WSZM25, YWR^{+}25, CDYL25, HWC25, JCW26, LSP^{+}25, LHZL26b, RK25, SP25a, SDH25, SW25, SHY25, WSX^{+}25, WZWC25b, XWW^{+}26]. synchronized [RMS25]. synergy [YK26, ZZZM25]. synthesis [GX25, LZZ^{+}26]. system [AKM25, BQS25, BEtM25, CG25a, CS25, DZY^{+}25, FBK25, HCQ26, KH26, LWY25, LDXL26, LW25a, QHZ25, SKM26, WL26, WS26, XTGZ25, ZLCS26, ZKZ26, ZZ25e, ZZJZ25]. systems [APJ25, AB25, BAT25, BRZS25, BM26, CWS^{+}25, CG25a, CZGY25, CL25, CGH^{+}25, CTTN25, CRHTV25, CCLL25, CXZ26, DZML25, EST^{+}25, EWH25, FY25, FLSS26, FA26, GX25, GCW^{+}26, GLY26, HLT25, HWC25, HY25, HLZ25, HZ26a, JXZ^{+}25, JYLS25, LZZ25, LXM25, LCW25, LHLL25, LZJ^{+}25, LSP^{+}25, LHZL25, LHZL26a, LLZL26, LMMG25, LSH25, LWS^{+}25, LLLH25, LZP25, LL25b, LZZ^{+}26, LW25d, MLW^{+}25, MM25, MCH25b, MS26, Mut26, NSX^{+}25, NNW25, PFZ25, PPC^{+}25, PPH26, QX25a, QMWY25, RHSZ26, RR26, SKSC26, SDH25, SZM25, SSK^{+}26, SZLG26, SWLC25, SWXL25, TO25, TXJW26, WDMM25, WZWC25a, WXD^{+}25, WLZ25a, WLZ26, WZLS26, WLLL26a, WWS^{+}25, WZJW25, WZP^{+}25, Wu25, WYKS25, XC25, XZSC26, YHGZ25, YBSC25, YGY^{+}25, YZ25, YHSZ26, YQQ^{+}26, ZWQC25, ZZ25c, ZLA25, ZFW25, ZLJY25, ZD25, ZYY26, ZLCS26, ZZD26, ZHWW26, ZXW^{+}25, ZZ25d, ZZS^{+}26, ZLL26, ZZ26b, ZZ25f, ZHZ^{+}26]. systems-based [LCW25]. Szeged [BDTP25, BDTP23].

T [ZLJY25, ZLCS26, ZZ25d, ZHZ^{+}26]. T-S [ZLJY25, ZLCS26, ZHZ^{+}26]. Tabu [PPN^{+}25a, PPN^{+}25b]. tapered [PPS25, RRR19, RRR20]. taxi [LWL^{+}25]. taxis [Che25]. Taylor [BSG^{+}25, FGJ25]. temporal [HCCH25, WQLW25]. tensor [LCLY25, LWML26, LLLZ25, TZL26]. tensors [Wan25]. term [MCH25a, TX25]. terminal [BGI25]. terms [BEtM25, YY25a]. tessellations [OCN^{+}26]. their [AAV25, BSG^{+}25]. theorem [Gao25]. Theory [ADZ25, CCC^{+}26, JZGV25, PDR25, PMVA26, SP25b, XS26]. theta [LLZ26, Mao25]. theta-free [LLZ26]. thin [PBS25]. three [DZY^{+}25, LXM25, LSQ25, LJ25, Tak25, Tak26, Wan25, WQLW25]. three-by-three [LXM25]. three-dimensional [LJ25, Wan25, WQLW25]. three-strategy [Tak25, Tak26]. threshold [KK25a, LZ25b, ZWL25]. thresholds [FBK25]. Tikhonov [DN25, PDR25]. Tikhonov-type [DN25]. tilings [LWLZ26]. Time [AT25, BQS25, LLZL26, LCRZ25, SGH^{+}25, AZD25, AAS25, ABdF^{+}25, ADPS25, BM26, CZGY25, CL25, CGH^{+}25, DZY^{+}25, DN25, EST^{+}25, EWH25, GCW^{+}26, HLT25, HWC25, HLZ25, ISC25, IJM25, JCHW25, JT25, LH25, LZJ^{+}25, LHZL25, LYQZ26, LZY^{+}25, LLLH25, LZP25, MLW^{+}25, MM25, MLTL25, Mut26, NSX^{+}25, PFZ25, PPC^{+}25, QX25a, RK25, SA26, SZM25, SFL26, SWW25, SZSW25, Tan25, WL25, WDMM25, WWQX25, WLZ26, WZLS26, WLLL26a, WZWC25b, WYKS25, XGL25, YHGZ25, YBSC25, YQYS25, ZT25, ZWQC25, ZN25, ZYY26, ZZ25d, ZZS^{+}26, ZKZ26, ZLL26, ZLH26, ZLCL25]. time-delay [CZGY25, PFZ25, SWW25]. time-delayed [SA26, WDMM25]. Time-dependent [AT25, BQS25]. time-discrete [AZD25]. time-domain [WLLL26a]. time-fractional [JCHW25, LYQZ26, Tan25, XGL25, ZT25]. time-space [DN25]. time-stepping [SFL26]. Time-varying [LLZL26, ABdF^{+}25, LH25, LHZL25, WL25, ZZ25d]. times [WPTK26]. timestamps [WWG^{+}25]. Toeplitz [BEG^{+}25, FLSS26, MW26]. Toeplitz-like [FLSS26]. tolerance [HY25, YZM25]. tolerant [CGH^{+}25, KNKR25, LCW25, PPC^{+}25, SZSW25, WS26, XZSC26, ZLCS26]. Topological [MAMFBM^{+}25, COP^{+}25]. topologies [WLZ26]. topology [WZWC25b, XTGZ25]. torus [LWLZ26]. total [CDD^{+}25, GVW25, JZGV25, KSFD25]. tough [HWS25]. tracing [SG26]. tracking [BAT25, CGS^{+}26, DZML25, JYLS25, LSH25, LC25, LCRZ25, Mut26, PSS^{+}25, SSK^{+}26, SZSW25, WS26, ZWQC25, ZLA25, ZD25, ZZ25d]. trade [YK26]. trading [QZ25]. Traditional [DCC^{+}25]. traffic [lCXZ^{+}25]. trajectory [KNKR25, ZLA25]. transfer [PSS^{+}25, WZP^{+}25, ZZ25c]. transform [CDM25, CYZ26, FLSS26, Gao25, Han26, MH25, RZ25]. transformation [SHY25]. transforms [JK25, KXL^{+}25, dOKA25]. transient [ABdF^{+}25, XLLC25]. transistor [FZY^{+}25]. transition [LL25b, PPC^{+}25]. transitive [GLD25, JLG25, LZCD25, ZHY26]. transmission [LGW^{+}26, SKSC26]. transport [CFS25, LL25c]. Trapezoid [SL26]. trapezoidal [SL26]. traveling [BK25]. Treating [WH26]. treatment [BDL25]. trees [AD25, BVZ25, LSWZ25, WJBL25, YT25, ZLY26]. triangles [PC25, Seb25, WJ25]. triangular [LXM25, RQWX25]. Tridiagonal [BFPS26]. trigger [LCW25]. triggered [APJ25, CWS^{+}25, CDYL25, DZY^{+}25, FA26, HLT25, HWC25, HCQ26, JCW26, LZJ^{+}25, LHZL25, LHZL26a, LLZL26, LC25, LZZ^{+}26, Mut26, PFZ25, PPC^{+}25, QX25a, QMWY25, SW25, SWW25, SZSW25, SWLC25, SWXL25, WYKS25, XWW^{+}26, YGY^{+}25, YZ25, ZWQC25, ZLCS26, ZXW^{+}25, ZXS^{+}25, ZZS^{+}26, ZZ25e, ZLCL25, ZHZ^{+}26]. triggering [WZP^{+}25, YBSC25]. tripartite [QWWW26]. truncated [CC25, Hag26, Mao25]. Trust [ABC^{+}26, SGH^{+}25]. TTCBLF [FA26]. tumor [BDL25]. tuning [LWS^{+}25]. Turán [LLX25]. Turkish [MBH^{+}25]. turn [LCW25]. turn-on [LCW25]. turn-on/off [LCW25]. twin [MNE26]. Two [DFKY26, IAK25, LCL25, MR25, WMW^{+}25, WMWY25, YQYS25, BVZ25, BGQR25, CG25b, GAYKS25, GZL^{+}26, HJF25, JY25b, KM25, LJ25, LGW^{+}26, PBS25, PPC^{+}25, RQWX25, SST26, Tan25, WWS^{+}25, YY26, ZT25, ZW25, ZLCS26]. two-branch [BVZ25]. two-dimensional [LJ25, PBS25, RQWX25, SST26, WWS^{+}25]. Two-disjoint-cycle-cover [DFKY26, LCL25]. two-field [HJF25]. two-grid [Tan25, ZT25]. two-layer [BGQR25, YY26]. Two-Level [IAK25, MR25]. two-phase [KM25]. Two-relaxation-time [YQYS25]. two-stage [GZL^{+}26]. two-step [PPC^{+}25, ZLCS26]. two-strain [LGW^{+}26]. type [BSG^{+}25, CDM25, CRHTV25, DN25, HWS25, HLZ25, KSR25, KCLZ25, LHZL26a, Liu25, LLLH25, MLZ26, PBS25, RBT25, RPD25, Yu25, ZZ25d]. type-2 [LHZL26a, LLLH25, ZZ25d]. types [TLW^{+}25].

UAV [KNKR25]. umbral [ADZ25]. unbalanced [LZY^{+}25]. unbounded [TKM25]. uncertain [BAT25, DZY^{+}25, FY25, LZZ25, LWS^{+}25, LZP25, LZ25b, PPC^{+}25, XZN^{+}25]. uncertainly [RZ25]. uncertainty [ACCO25, CDM25, LZ25b, YY26]. Unconditionally [WGeL25]. unconstrained [YY25a]. unequal [YLR^{+}25]. unicyclic [MP25, WJBL25]. unified [CK25, EST^{+}25]. uniform [JSZ25, PP25, PDV25, Tan25, ZZ25a]. union [PC25]. uniqueness [FHX26]. Univariate [BD25]. unknown [CWS^{+}25, CGS^{+}26, CGH^{+}25, JYLS25, LHLL25, LL25b, PFZ25, QX25a, QX25b, SZM25, YGY^{+}25, YQQ^{+}26, ZFW25, ZLH26, ZZ25f, ZLCL25]. unlimited [BAG^{+}25]. unstable [NSX^{+}25]. Unsteady [FGR25]. unstructured [RQWX25]. update [WL25, YY25b]. updating [LW25c]. upper [LXM25]. Urysohn [BMM25]. used [BKE25]. useful [LZ25b]. user [SZSW25]. user-defined [SZSW25]. users [ABC^{+}26]. using [BSG^{+}25, Chr26, HS25, JH25, LT25, MBF25, NVGA^{+}26, OGPFVV25, WS26, ZG26b].

Vaccination [DGD^{+}26, LGW^{+}26, LdMAS25]. vaccine [LGW^{+}26, LdMAS25, YK26]. validation [KBM25, RR26]. validity [LH25]. Valuation [ZHLL25, TLSVGR25, XZN^{+}25]. value [BSM^{+}26, BGI25, HRS25, LWML26, VL26, ZYZ^{+}25]. valued [XWW^{+}26, ZYL25]. values [CF25]. Variable [JZ25, BCCF25, BGV26, LCRZ25, PPH26, QXZ25]. Variable-coefficient [JZ25]. variable-order [PPH26]. variant [FGW25]. variants [LT25]. variation [LLMT25]. variational [HZF^{+}25]. Variety [KV25]. varying [ABdF^{+}25, LH25, LHZL25, LLZL26, WL25, ZZ25d]. Vector [Kov25, CFS25, MNE26, ZZT25]. vectorial [CTTN25]. vectorization [ML26]. vehicles [lCXZ^{+}25, LCRZ25, ZHLL25]. velocity [FA26]. verification [ABD25]. verifying [TO25]. version [JT25]. vertex [DFKY26, ET26, GAYKS25]. vertices [AD25, BBV25]. very [LD25]. vessels [RPS^{+}25]. via [BMM25, CLWZ26, CCLL25, EST^{+}25, HCQ26, HLZ25, JXZ^{+}25, JCW26, LLMT25, LHJ^{+}25, LSP^{+}25, LDXL26, LW25a, MBH^{+}25, Mut26, PPC^{+}25, SWXL25, WDMM25, WWS^{+}25, WSZM25, WZWC25b, WYKS25, YGY^{+}25, YWR^{+}25, YWJ^{+}25, ZXS^{+}25, ZHZ^{+}26]. Vicsek [ZZTL25]. view [ADZ25]. viewpoint [TF25]. Virtual [AZD25, EP25, HAD25]. viscoelastic [LSF^{+}25]. viscoplastic [FGR25]. viscous [WF25]. visibility [CDD^{+}25, GKR^{+}25, KV25, MBH^{+}25, RKS25]. Visualization [OCN^{+}26]. vital [ABC^{+}26]. VIX [Grz26]. Vlasov [LZ26, WXT25]. volatility [CK25, Ruj25, ZG26a]. voltage [LC25]. Volterra [BKE25, Che25, PTB25, YYT25, ZL25]. Volume [WMZ25a]. Vortical [FP25]. Vries [SFL26]. vs [DAG25].

walk [ZHF25]. walks [DMT25, MP26]. wave [ALC^{+}25, CLLF25, LCS^{+}26b, WF25, WGeL25]. wavelet [RZ25, ZYZ^{+}25]. wavelet-based [ZYZ^{+}25]. wavelets [TX25]. waves [CS25, FP25]. waving [LSP^{+}25]. Weak [MAS26, YZH25, ADR^{+}26, Tan25]. weakly [KXL^{+}25, PTB25, YM25]. weakly- [YM25]. WeiBo [TF25]. weight [LW25b]. Weighted [MWS25, GX25, HCTS25, Mez25, MS26, QTW25]. weights [XTGZ25]. well [BGV26, KCLZ25]. well-balanced [BGV26]. WENO [CSB25, SH25, WH26]. Weyl [CDM25]. WG [MAS26]. wheel [ZLW25].

Whitham [FP25]. whose [XW25]. width [LL26a]. wind [KH25]. windup [CZGY25]. wise [LZJ^{+}25]. within [TLW^{+}25]. without [BSG^{+}25, EJJ26, GCV25, LL25a, PC25, SHY25, WWG^{+}25, WWL25]. Words [KH25]. Wright [AA25]. written [XZN^{+}25].

Yeh [DS25].

Zernike [DEMFRL26, ECD^{+}25]. Zernikelike [DEMFRL26]. ZERNIPAX [ECD^{+}25]. Zero [ZJX25, ALNP25, GPL25, Ued25]. Zero-determinant [ZJX25, Ued25]. zeroorder [ALNP25]. zero-sum [GPL25]. zone [LHZL25, WZWC25a]. zooming [CSB25].

References

Askari:2025:SPM

[AA25] Hassan Askari and Alireza Ansari. Stokes phenomenon for the M-Wright function of order 1n. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005496.

Aghapour:2025:DTN

[AAS25] Ahmad Aghapour, Hamid Arian, and Luis Seco. Deeptime neural networks: an efficient approach for solving high-dimensional PDEs. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005782.

Abiad:2025:DGT

[AAV25] Aida Abiad, Carlos A. Alfaro, and Ralihe R. Villagrán. Distinguishing graphs by their spectra, Smith normal forms and complements. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006593.

Ayub:2025:NEL

[AB25] Muhammad Ayub and Saira Bano. Natural extension of a Lie algebra A_{4,18}^{b}, |b| ≤ 1, realizations, invariant systems and integrability. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000013.

Abbasbandy:2026:SLN

[Abb26] Saeid Abbasbandy. Solving linear and nonlinear Caputo fractional differential equations with a quantum pseudo-spectral approach. Applied Mathematics and Computation, 511 (??):??, February 15, 2026.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004515.

Alalawi:2026:TAR

[ABC^{+}26] Zainab Alalawi, Paolo Bova, Theodor Cimpeanu, Alessandro Di Stefano, Manh Hong Duong, Elias Fernández Domingos, The Anh Han, Marcus Krellner, Ndidi Bianca Ogbo, Simon T. Powers, and Filippo Zimmaro. Trust AI regulation? discerning users are vital to build trust and effective AI regulation. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003534.

Aksoy:2025:NVC

[ABD25] Birgül Aksoy, Taner Büyükköroglu, and Vakif Dzhafarov. On the numerical verification of a counterexample on parameter-dependent Lyapunov functions. Applied Mathematics and Computation, 492(??):??, May 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324007070.

Antonacci:2025:ETN

[ABdF^{+}25] Yuri Antonacci, Chiara Barà, Giulio de Felice, Antonino Sferlazza, Riccardo Pernice, and Luca Faes. Exploring transient neurophysiological states through local and time-varying measures of information dynamics. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500164X.

Arnal:2025:EPC

[ACC25] Ana Arnal, Fernando Casas, and Cristina Chiralt. An efficient procedure to compute the continuous Baker– Campbell–Hausdorff formula. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002899.

Abugattas:2025:QUI

[ACCO25] Carolina Abugattas, Ana Carpio, Elena Cebrián, and Gerardo Oleaga. Quantifying uncertainty in inverse scattering problems set in layered environments. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001808.

Ahmad:2025:CSM

[AD25] Sultan Ahmad and Kinkar Chandra Das. A complete solution for maximizing the general Sombor index of chemical trees with given number of pendant vertices. Applied Mathematics and Computation, 505 (??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002589.

Andelic:2026:EPB

[AdFS26] Milica Andelić, Carlos M.

da Fonseca, and Zoran Stanić. An equitable partition based construction of graphs with the same spectral radius. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004229.

Araujo:2025:ITG

[ADPS25] Julio Araujo, Mitre C.

Dourado, Fábio Protti, and Rudini Sampaio. The iteration time and the general position number in graph convexities. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005459.

Alam:2026:PCI

[ADR^{+}26] Khairul Habib Alam, Aswini Dolai, Yumnam Rohen, Sunil Panday, and Shaibal Mani. On Picard-CR iterations involving weak perturbative contraction operators and application to reversible chemical reactions. Applied Mathematics and Computation, 512(??): ??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004692.

Artioli:2025:THL

[ADZ25] M. Artioli, G. Dattoli, and U. Zainab. Theory of Hermite and Laguerre Bessel function from the umbral point of view. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005642.

Alfaro:2025:DID

[AHMSV25] Carlos A. Alfaro, Teresa I.

Hoekstra-Mendoza, Juan Pablo Serrano, and Ralihe R. Villagrán. Distance ideals of digraphs. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001572.

An:2025:OPR

[AJK25] Jongbong An, Junkee Jeon, and Takwon Kim. Optimal portfolio and retirement decisions with costly job switching options. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006763.

Ahmad:2026:CGS

[AK26] Sk. Safique Ahmad and Pinki Khatun. A class of generalized shift-splitting preconditioners for double saddle point problems. Applied Mathematics and Computation, 509(??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003844.

Arun:2025:APE

[AKM25] K. R. Arun, A. Krishnamurthy, and H. Maharna. An asymptotic preserving and energy stable scheme for the Euler system with congestion constraint. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000335.

Alcantara:2025:NAN

[ALC^{+}25] Adriano A. Alcântara, Juan B. Límaco, Bruno A. Carmo, Ronald R. Guardia, and Mauro A. Rincon. Numerical analysis for nonlinear wave equations with boundary conditions: Dirichlet, acoustics and impenetrability. Applied Mathematics and Computation, 484(??): ??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004703.

Abney:2025:CPA

[ALNP25] Ray Abney, Thuy T. Le, Loc H. Nguyen, and Cam Peters. A Carleman–Picard approach for reconstructing zero-order coefficients in parabolic equations with limited data. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500013X.

Aledo:2025:MSS

[ALV25] Juan A. Aledo, Jose P.

Llano, and Jose C. Valverde. Multi-state sequential Boolean networks: an analysis of its combinatorial dynamics. Applied Mathematics and Computation, 489(??): ??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005769.

Aghamollaei:2026:SRC

[AMMT26] Gholamreza Aghamollaei, Mahdiyeh Mortezaei, Dijana Mosić, and Néstor Thome. Some results on core EP Drazin matrices and partial isometries. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500373X.

Anonymous:2025:Aa

[Ano25a] Anonymous. 1 April 2025.

Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ac

[Ano25b]  Anonymous.  1 August

2025. Applied Mathematics and Computation, 498 (??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Da

[Ano25c] Anonymous. 1 December 2025. Applied Mathematics and Computation, 506 (??):??, December 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Fa

[Ano25d] Anonymous. 1 February 2025. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ja

[Ano25e]  Anonymous.  1 January

2025. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Je

[Ano25f] Anonymous. 1 July 2025. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Jc

[Ano25g] Anonymous. 1 June 2025.

Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ma

[Ano25h]  Anonymous.  1 March

2025. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Mc

[Ano25i] Anonymous. 1 May 2025.

Applied Mathematics and Computation, 492(??):??, May 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Na

[Ano25j] Anonymous. 1 November 2025. Applied Mathematics and Computation, 504 (??):??, November 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Oa

[Ano25k]  Anonymous.  1 October

2025. Applied Mathematics and Computation, 502 (??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Sa

[Ano25l] Anonymous. 1 September 2025. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ab

[Ano25m]  Anonymous.  15 April

2025. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ad

[Ano25n] Anonymous. 15 August 2025. Applied Mathematics and Computation, 499 (??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Db

[Ano25o] Anonymous. 15 December 2025. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Fb

[Ano25p] Anonymous. 15 February 2025. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Jb

[Ano25q] Anonymous. 15 January 2025. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Jf

[Ano25r] Anonymous. 15 July 2025.

Applied Mathematics and Computation, 497(??):??, July 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Jd

[Ano25s] Anonymous. 15 June 2025.

Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Mb

[Ano25t]  Anonymous.  15 March

2025. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Md

[Ano25u] Anonymous. 15 May 2025. Applied Mathematics and Computation, 493(??):??, May 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Nb

[Ano25v] Anonymous. 15 November 2025. Applied Mathematics and Computation, 505 (??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Ob

[Ano25w] Anonymous. 15 October 2025. Applied Mathematics and Computation, 503 (??):??, October 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Sb

[Ano25x] Anonymous. 15 September 2025. Applied Mathematics and Computation, 501 (??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:EBa

[Ano25y]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005101.

Anonymous:2025:EBb

[Ano25z]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005241.

Anonymous:2025:EBc

[Ano25-27]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006015.

Anonymous:2025:EBd

[Ano25-28]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006313.

Anonymous:2025:EBe

[Ano25-29]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006544.

Anonymous:2025:EBf

[Ano25-30]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006866.

Anonymous:2025:EBg

[Ano25-31]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400701X.

Anonymous:2025:EBh

[Ano25-32]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324007227.

Anonymous:2025:EBi

[Ano25-33]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 492 (??):??, May 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000189.

Anonymous:2025:EBj

[Ano25-34]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 493 (??):??, May 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000505.

Anonymous:2025:EBk

[Ano25-35]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 494 (??):??, June 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000785.

Anonymous:2025:EBl

[Ano25-36]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001006.

Anonymous:2025:EBm

[Ano25-37]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 496 (??):??, July 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001122.

Anonymous:2025:EBn

[Ano25-38]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 497 (??):??, July 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001493.

Anonymous:2025:EBo

[Ano25-39]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 498 (??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001869.

Anonymous:2025:EBp

[Ano25-40]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 499 (??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002085.

Anonymous:2025:EBq

[Ano25-41]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002164.

Anonymous:2025:EBr

[Ano25-42]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 501 (??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002401.

Anonymous:2025:EBs

[Ano25-43]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 502 (??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002644.

Anonymous:2025:EBt

[Ano25-44]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 503 (??):??, October 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002462.

Anonymous:2025:EBu

[Ano25-45]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 504 (??):??, November 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500298X.

Anonymous:2025:EBv

[Ano25-46]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 505 (??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002802.

Anonymous:2025:EBw

[Ano25-47]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 506 (??):??, December 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500308X.

Anonymous:2025:EBx

[Ano25-48]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003819.

Anonymous:2026:Ja

[Ano26a]  Anonymous.  1 January

2026. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2026:Ma

[Ano26b]  Anonymous.  1 March

2026. Applied Mathematics and Computation, 512(??): ??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2026:F

[Ano26c] Anonymous. 15 February 2026. Applied Mathematics and Computation, 511 (??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2026:Jb

[Ano26d] Anonymous. 15 January 2026. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2026:EBa

[Ano26e]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004564.

Anonymous:2026:EBb

[Ano26f]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004801.

Anonymous:2026:EBc

[Ano26g]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 511 (??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325005247.

Anonymous:2026:EBd

[Ano26h]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 512 (??):??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500551X.

Ai:2025:ETS

[APJ25] Zidong Ai, Lianghong Peng, and Yongzhi Jing. Eventtriggered stabilization for switched nonlinear systems under state-overshooting constraint. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300325002759.

Ackley:2021:LDN

[AS21] Matthew Ackley and Peter Stechlinski. Lexicographic derivatives of nonsmooth glucose-insulin kinetics under normal and artificial pancreatic responses. Applied Mathematics and Computation, 395(??):??, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300320308298.

See corrigendum [AS25].

Ackley:2025:CLD

[AS25] Matthew Ackley and Peter Stechlinski. Corrigendum to “Lexicographic derivatives of nonsmooth glucose– insulin kinetics under normal and artificial pancreatic responses” [Appl. Math. Compt. 395 (2021) 125876]. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/

                                        science/article/pii/S0096300324005940.

See [AS21].

Anwar:2025:TDS

[AT25] Muhammad Adnan Anwar and Jorge Tiago. Timedependent strategy for improving aortic blood flow simulations with boundary control and data assimilation. Applied Mathematics and Computation, 493(??): ??, May 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324007276.

Abbaszadeh:2025:VEA

[AZD25] Mostafa Abbaszadeh, Mahmoud A. Zaky, and Mehdi Dehghan. Virtual element approximation and BDF2 time-discrete scheme for a partial integro-differential equation with a singular Abel’s kernel. Applied Mathematics and Computation, 501(??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500178X.

An:2026:EMP

[AZZ^{+}26] Tianbo An, Ran Zhang, Jian Zhao, Huizhen Zhang, and Zhen Wang. Evolutionary multigame with perception competition between conformists and imitators. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004382.

Bayon:2025:NES

[BAG^{+}25] L. Bayón, P. Fortuy Ayuso, J. M. Grau, A. M. Oller-Marcén, and M. M. Ruiz. Nash equilibria in the Showcase Showdown game with unlimited spins. Applied Mathematics and Computation, 497(??):??, July 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300325000955.

Baghi:2025:SAO

[BAT25] Hamidreza Baghi, Farzaneh Abdollahi, and Heidar Ali Talebi. Secure adaptive output-feedback tracking control for a class of uncertain nonlinear systems in the presence of sensor and actuator faults under DoS and injection attacks. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400568X.

Bassik:2026:RPE

[BBG^{+}26] Oren Bassik, Yosef Berman, Soo Go, Hoon Hong, Ilia Ilmer, Alexey Ovchinnikov, Chris Rackauckas, Pedro Soto, and Chee Yap. Robust parameter estimation for rational ordinary differential equations. Applied Mathematics and Computation, 509(??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003649.

Baudon:2025:PVE

[BBV25] Olivier Baudon, Julien Bensmail, and Lyn Vayssieres. Partitioning vertices and edges of graphs into connected subgraphs. Applied Mathematics and Computation, 505(??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002577.

Bartoli:2025:MPM

[BCCF25] Laura Bartoli, Simone Cacace, Emiliano Cristiani, and Roberto Ferretti. A macroscopic pedestrian model with variable maximal density. Applied Mathematics and Computation, 499 (??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001316.

Bejancu:2025:UIC

[BD25] Aurelian Bejancu and Mohamed Dekhil. Univariate interpolation for a class of L-splines with adjoint natural end conditions. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001444.

Biscevic:2025:CSM

[BDD25] Helena Biscević, Raffaele D’Ambrosio, and Stefano Di Giovacchino. Contractivity of stochastic θ-methods under non-global Lipschitz conditions. Applied Mathematics and Computation, 505 (??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500253X.

Bulai:2025:NSM

[BDL25] I. M. Bulai, M. C. De Bonis, and C. Laurita. Numerical solution of metastatic tumor growth models with treatment. Applied Mathematics and Computation, 484(??):??, January 1, 2025.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004491.

Baraviera:2025:IRN

[BDST25] Alexandre Baraviera, Pedro Duarte, Longmei Shu, and Maria Joana Torres. Isospectral reductions of non-negative matrices. Applied Mathematics and Computation, 499(??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001298.

Brezovnik:2023:SMR

[BDTP23] Simon Brezovnik, Matthias Dehmer, Niko Tratnik, and Petra Žigert Pleteršek. Szeged and Mostar rootindices of graphs. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322008049. See

corrigendum [BDTP25].

Brezovnik:2025:CSM

[BDTP25] Simon Brezovnik, Matthias Dehmer, Niko Tratnik, and Petra Žigert Pleteršek. Corrigendum to “Szeged and Mostar root-indices of graphs”. Applied Mathematics and Computation, 499 (??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001717.

See [BDTP23].

Barbarino:2025:SPF

[BEG^{+}25] Giovanni Barbarino, Sven-Erik Ekström, Carlo Garoni, David Meadon, Stefano Serra-Capizzano, and Paris Vassalos. Spectral properties of flipped Toeplitz matrices and computational applications. Applied Mathematics and Computation, 499(??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001353.

Boutaayamou:2025:NCO

[BEtM25] Idriss Boutaayamou, Fouad Et-tahri, and Lahcen Maniar. Null controllability of an ODE-heat system with coupled boundary and internal terms. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500030X.

Bazeia:2025:MBH

[BFdOdS25] D. Bazeia, M. J. B. Ferreira, B. F. de Oliveira, and W. A. dos Santos. Monitoring biodiversity on highly reactive rock-paper-scissors models. Applied Mathematics and Computation, 502 (??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002000.

Bresar:2025:PCD

[BFH^{+}25] Bostjan Bresar, Jasmina Ferme, Premysl Holub, Marko Jakovac, and Petra Melicharová. S-packing colorings of distance graphs with distance sets of cardinality 2. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006611.

Bueno:2026:TMM

[BFPS26] M. I. Bueno, Susana Furtado, K. Kranthi Priya, and K. C. Sivakumar. Tridiagonal M-matrices with tridiagonal Moore–Penrose inverse. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003662.

Brugnano:2025:SNP

[BGI25] Luigi Brugnano, Gianmarco Gurioli, and Felice Iavernaro. A shooting-Newton procedure for solving fractional terminal value problems. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006258.

Besabe:2025:LNF

[BGQR25] Lander Besabe, Michele Girfoglio, Annalisa Quaini, and Gianluigi Rozza. Linear and nonlinear filtering for a two-layer quasi-geostrophic ocean model. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005824.

Burger:2026:ESW

[BGV26] Raimund Bürger, Andrés Guerra, and Carlos A. Vega. An entropy stable and well-balanced scheme for an augmented blood flow model with variable geometrical and mechanical properties. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500445X.

Bi:2025:IRL

[BHL25] Yan Bi, Qingyi Hao, and Kui Liu. The influence of reputation levels and disparities on the evolution of cooperation in spatial prisoner’s dilemma game. Applied Mathematics and Computation, 501 (??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001997.

Bulbul:2025:NMM

[BK25] K. Gulnaz Bulbul and Refail Kasimbeyli. A new mathematical model and solution method for the asymmetric traveling salesman problem with replenishment arcs. Applied Mathematics and Computation, 494 (??):??, June 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000050.

Behroozi:2025:SAM

[BKE25] Hasan Behroozi, Manochehr Kazemi, and Reza Ezzati. Successive approximations method is used to solve nonlinear Volterra delay integral equations. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002735.

Buckwar:2025:NAC

[BM25] Evelyn Buckwar and Amira Meddah. Numerical approximations and convergence analysis of piecewise diffusion Markov processes, with application to glioma cell migration. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006945.

Buzikov:2026:NMS

[BM26] Maksim Buzikov and Alina Mayer. Numerical methods for solving minimumtime problem for linear systems. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003601.

Bhat:2025:CAN

[BMM25] Imtiyaz Ahmad Bhat, Lakshmi Narayan Mishra, and Vishnu Narayan Mishra. Comparative analysis of nonlinear Urysohn functional integral equations via Nyström method. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300325000141.

Benko:2025:EGC

[BPL^{+}25] Tina P. Benko, Bin Pi, Qin Li, Minyu Feng, Matjaž Perc, and Helena Blažun Vošner. Evolutionary games for cooperation in open data management. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000918.

Bilfaqih:2025:TDR

[BQS25] Yusuf Bilfaqih, Mochamad Nur Qomarudin, and Mochammad Sahal. Time-dependent reliability computation of system with multistate components. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005435.

Brugnoli:2025:FEH

[BRZS25] Andrea Brugnoli, Ramy Rashad, Yi Zhang, and Stefano Stramigioli. Finite element hybridization of port-Hamiltonian systems. Applied Mathematics and Computation, 498 (??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001043.

Braverman:2025:SBL

[BS25] Elena Braverman and John Ioannis Stavroulakis. Symmetry and the Buchanan– Lillo conjecture: a resolution of the mixed feedback case. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001031.

Bate:2025:JTM

[BSG^{+}25] Indra Bate, Kedarnath Senapati, Santhosh George, Muniyasamy M., and Chandhini G. Jarratt-type methods and their convergence analysis without using Taylor expansion. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005733.

Bourantas:2026:MPC

[BSM^{+}26] G. C. Bourantas, A. Sakellarios, N. Malamos, V. C. Loukopoulos, V. N. Burganos, D. I. Fotiadis, and V. M. Calo. A meshless point collocation solver for elliptic boundary value problems. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003996.

Bantva:2025:RLT

[BVZ25] Devsi Bantva, Samir Vaidya, and Sanming Zhou. Radio labelling of two-branch trees. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005587.

Capitan:2025:EIF

[CA25] José A. Capitán and David Alonso. Out-of-equilibrium inference of feeding rates through population data from generic consumerresource stochastic dynamics. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001614.

Cambie:2026:TES

[Cam26]  Stijn Cambie.  Towards

the essence of Soltés’ problem. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003479.

Cusick:2025:EQT

[CC25] Thomas W. Cusick and Younhwan Cheon. Extended quadratic truncated rotation symmetric Boolean functions. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003273.

Chen:2026:IBP

[CC26] Hao Chen and Weicheng Cui. Intergroup-based public goods games in complex networks with leaderdriven other-regarding preferences. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003492.

Cheung:2026:TNS

[CCC^{+}26] Siu Wun Cheung, Youngsoo Choi, Seung Whan Chung, Jean-Luc Fattebert, Coleman Kendrick, and Daniel Osei-Kuffuor. Theory and numerics of subspace approximation of eigenvalue problems. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004448.

Cavoretto:2026:CDM

[CCD26] Roberto Cavoretto, Chiara Comoglio, and Alessandra De Rossi. Community detection methods for GBF-PUM signal approximation on graphs. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500428X.

Cui:2025:ESS

[CCLL25] Qian Cui, Jinde Cao, Lulu Li, and Yang Liu. Exponential stability of switched systems with state-dependent delayed impulses via Bequivalent method. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004922.

Cicerone:2025:MTM

[CDD^{+}25] Serafino Cicerone, Alessia Di Fonso, Gabriele Di Stefano, Alfredo Navarra, and Francesco Piselli. Mutual and total mutual visibility in hypercube-like graphs. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006775.

Chen:2025:THP

[CDM25] Xuan Chen, Pei Dang, and Weixiong Mai. L^{p}- type Heisenberg–Pauli–Weyl uncertainty principles for fractional Fourier transform. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006970.

Chen:2025:MBT

[CDYL25] Mengshen Chen, Derui Ding, Huaicheng Yan, and Kun Liang. Mode-based triggered protocol for synchronization of switched dynamic networks with semi-Markov kernel and its application. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000633.

Cococcioni:2025:LPI

[CF25] Marco Cococcioni and Lorenzo Fiaschi. Linear programming with infinite, finite, and infinitesimal values in the right-hand side. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005058.

Chen:2025:NCG

[CFS25] Kangming Chen, Ellen Hidemi Fukuda, and Hiroyuki Sato. Nonlinear conjugate gradient method for vector optimization on Riemannian manifolds with retraction and vector transport. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004624.

Chalons:2025:SOS

[CG25a] Christophe Chalons and Mathieu Girardin. Secondorder schemes based on a relaxation system for scalar and systems of conservation laws. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006349.

Cheng:2025:EFT

[CG25b] Panpan Cheng and Yunshu Gao. Extremal function of two independent chorded cycles in a bipartite graph. Applied Mathematics and Computation, 493(??):??, May 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007148.

Cheng:2025:FTA

[CGH^{+}25] Xuanrui Cheng, Ming Gao, Wuxiang Huai, Yichun Niu, and Li Sheng. Fixedtime active fault-tolerant control for dynamical systems with intermittent faults and unknown disturbances. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005150.

Chai:2026:IBO

[CGS^{+}26] Yuanjun Chai, Chao Ge, Yanpeng Shi, Zhinan Peng, and Huiyan Zhang. Impulsebased output consensus tracking of heterogeneous clustered networks subjected to unknown bounded cyberattacks. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003868.

Chen:2026:CRP

[CH26] Yahong Chen and He Huang. Collaborated risk perception against multiple epidemics in a multiplayer network. Applied Mathematics and Computation, 511 (??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004631.

Chen:2025:PDL

[Che25] Mengxin Chen. Pattern dynamics of a Lotka–Volterra model with taxis mechanism. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004788.

Christou:2026:NSE

[Chr26] Katerina Christou. The numerical solution of epidemic models on moving domains using combined masses. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004175.

Chen:2025:NNI

[CjZjHbX25] Tao Chen, Yao jia Zhang, Nan jing Huang, and Yi bin Xiao. A new nonlocal impulsive fractional differential hemivariational inclusions with an application to a frictional contact problem. Applied Mathematics and Computation, 490(??): ??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006726.

Choi:2025:UMS

[CK25] Sun-Yong Choi and Jeong-Hoon Kim. A unified model of SABR and meanreverting stochastic volatility for derivative pricing. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500325X.

Chen:2025:CFB

[CL25] Chen Chen and Jinghao Li. A composite filterbased approach to adaptive prescribed-time outputfeedback control of strictfeedback nonlinear systems with output and input quantization. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000062.

Chen:2025:MEE

[CLLC25] Xiaofeng Chen, Dongyuan Lin, Hua Li, and Zhi Cheng. Minimum error entropy high-order extend Kalman filter with fiducial points. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005745.

Chai:2025:SPC

[CLLF25] Li Chai, Yang Liu, Hong Li, and Zhichao Fang. Structure-preserving compact ADI schemes for the space fractional Klein– Gordon–Schrödinger equations and the dynamic simulation of solitary wave solutions. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001705.

Chen:2026:LSB

[CLWZ26] Bingquan Chen, Bowen Li, Tao Wu, and Yanling Zheng. Local stabilization of Boolean control networks via stochastic sampled-data control. Applied Mathematics and Computation, 511 (??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004710.

Canovas:2025:DLH

[CMG25] J. S. Cánovas and M. Muñoz-Guillermo. On the dynamics of a linear-hyperbolic population model with Allee effect and almost sure extinction. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004661.

Choi:2025:EDR

[COP^{+}25] Seonmi Choi, Jinseok Oh, Jeong Rye Park, Seung Yeop Yang, and Hongdae Yun. Effective data reduction algorithm for topological data analysis. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000293.

Conte:2026:HON

[CPR26] D. Conte, G. Pagano, and T. Roldán. High order nonstandard finitedifference methods. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004072.

Cordero:2025:MED

[CRHTV25] Alicia Cordero, Renso V. Rojas-Hiciano, Juan R. Torregrosa, and Maria P. Vassileva. Maximally efficient damped composed Newtontype methods to solve nonlinear systems of equations. Applied Mathematics and Computation, 492 (??):??, May 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006921.

Crokidakis:2025:MMB

[Cro25] Nuno Crokidakis. A mathematical model for the bullying dynamics in schools. Applied Mathematics and Computation, 492(??):??, May 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400715X.

Cotronei:2025:SIR

[CRS25] Mariantonia Cotronei, Dörte Rüweler, and Tomas Sauer. Sparse image representation through multiple multiresolution analysis. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001675.

Chhatria:2025:ATN

[CS25] Balakrishna Chhatria and T. Raja Sekhar. Asymptotics towards nonlinear diffusion waves for the solutions of a hyperbolic system with linear damping on quadrant. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003030.

Crnkovic:2025:WBA

[CSB25] Bojan Crnković, Jerko Skifić, and Tina Bosner. WENO based adaptive image zooming algorithm. Applied Mathematics and Computation, 492(??):??, May 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006891.

Cordero:2025:FOV

[CTTN25] Alicia Cordero, Juan R. Torregrosa, and Paula Triguero-Navarro. First optimal vectorial eighth-order iterative scheme for solving non-linear systems. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001286.

Cai:2025:BLF

[CWS^{+}25] Yuliang Cai, Youtong Wang, Hanguang Su, Lianyan Fu, and Qiang He. Bipartite leader-following consensus of linear multi-agent systems with unknown disturbances under directed graphs by double dynamic event-triggered mechanism. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000748.

Chen:2025:FBD

[CWWT25] Jianjun Chen, Yuxuan Wang, Weiyi Wang, and Zhijun Tan. A FFT-based DDSIIM solver for elliptic interface problems with discontinuous coefficients on arbitrary domains and its error analysis. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005472.

Cui:2026:DRL

[CXZ26] Lili Cui, Xiangpeng Xie, and Yong Zhang. Distributed reinforcement learning-based FTC for nonlinear multiagent systems with performance guarantees. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300325004916.

Chen:2026:STB

[CYZ26] Chao Chen, Xi Yang, and Fei-Yan Zhang. Sinetransform-based fast solvers for Riesz fractional nonlinear Schrödinger equations with attractive nonlinearities. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500400X.

Chen:2025:AWD

[CZGY25] Yonggang Chen, Yaxue Zhao, Zhou Gu, and Xinfen Yang. Anti-windup design for networked timedelay systems subject to saturating actuators under round-robin protocol. Applied Mathematics and Computation, 499(??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001407.

Divya:2025:MPP

[DAG25] G. Divya, S. Athithan, and Mini Ghosh. Modelling the prevalence of prostitution under the influence of poverty: a deterministic vs. stochastic approach. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004910.

Ding:2025:TCM

[DCC^{+}25] Rui Ding, Shujuan Cao, Binying Cai, Yongming Zou, and Fang xiang Wu. Traditional Chinese medicine studies for AD based on Logistic Matrix Factorization and Similarity Network Fusion. Applied Mathematics and Computation, 496 (??):??, July 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000736.

Diaz-Elbal:2026:SPZ

[DEMFRL26] S. Díaz-Elbal, A. Martínez-Finkelshtein, and D. Ramos-López. Sampling patterns for Zernike-like bases in nonstandard geometries. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004527.

Dai:2026:TDC

[DFKY26] Hanle Dai, Yan-Quan Feng, Young Soo Kwon, and Da-Wei Yang. Two-disjoint-cycle-cover vertex bipancyclicity of bubble-sort star graphs. Applied Mathematics and Computation, 512 (??):??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325005120.

Das:2026:VDM

[DGD^{+}26] Susanta Kumar Das, Adrita Ghosh, Parthasakha Das, K. M. Ariful Kabir, Pritha Das, and Dibakar Ghosh. Vaccination dilemmas in mitigating monkeypox outbreaks: an imitation dynamics game model approach. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003674.

Diblik:2026:SDM

[Dib26] Josef Diblík. Solution of a discrete matrix equation with second-order difference and single delay. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003339.

Dravec:2025:GSS

[DMT25] Tanja Dravec, Mirjana Mikalacki, and Andrej Taranenko. Graphs with span 1 and shortest optimal walks. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001602.

Duc:2025:TTM

[DN25] Nguyen Van Duc and Thi-Phong Nguyen. A Tikhonov-type method for inverse source problems for time-space fractional parabolic equations. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002930.

deOliveira:2025:ITE

[dOKA25] André J. P. de Oliveira, Diego C. Knupp, and Luiz A. S. Abreu. Integral transforms for explicit source estimation in non-linear advection–diffusion problems. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005538.

Dou:2025:CBM

[DS25] Donna Quanjie Dou and Lisa Hui Sun. A conjecture on Boros–Moll polynomials due to Ma, Qi, Yeh and Yeh. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006477.

Deng:2025:RBR

[DWL^{+}25] Lili Deng, Rugen Wang, Ying Liao, Ronghua Xu, and Cheng Wang. The reputation-based reward mechanism promotes the evolution of fairness. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005034.

Dong:2025:OFH

[DZML25] Xu Dong, Huaguang Zhang, Zhongyang Ming, and Yanhong Luo. Optimal finitehorizon tracking control in affine nonlinear systems: a Stackelberg game approach with H_{2}/H_{∞} framework. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000037.

Du:2025:DTE

[DZY^{+}25] Zhaoping Du, Chi Zhang, Xiaofei Yang, Hui Ye, and Jianzhen Li. Discrete-time event-triggered h-infinity stabilization for three closedloop cyber-physical system with uncertain delay. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005885.

Elmacioglu:2025:ZFA

[ECD^{+}25] Yigit Gunsur Elmacioglu, Rory Conlin, Daniel W. Dudt, Dario Panici, and Egemen Kolemen. ZERNI-PAX: a fast and accurate Zernike polynomial calculator in Python. Applied Mathematics and Computation, 505(??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002607.

Eze:2026:GC

[EJJ26] Leonard Chidiebere Eze, Robert Jajcay, and Jorik Jooken. On (k, g)-graphs without (g + 1)-cycles. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003716.

Enabe:2025:HVE

[EP25] Paulo Akira F. Enabe and Rodrigo Provasi. A hybrid virtual element method and deep learning approach for solving one-dimensional Euler–Bernoulli beams. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003261.

Elayabharath:2025:ROB

[EST^{+}25] V. T. Elayabharath, P. Sozhaeswari, N. Tatar, R. Sakthivel, and T. Satheesh. Resilient observer-based unified state and fault estimation for nonlinear parabolic PDE systems via fuzzy approach over finite-time interval. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005861.

Erves:2026:ICV

[ET26] Rija Erves and Aleksandra Tepeh. Induced cycles vertex number and (1,2)-domination in cubic graphs. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004266.

Elshenhab:2025:ESF

[EWH25] Ahmed M. Elshenhab, Xing Tao Wang, and Mohamed Hosny. Explicit solutions and finitetime stability for fractional delay systems. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001158.

Fu:2026:STC

[FA26] Longbin Fu and Liwei An.

Self-triggered control of robotic systems with obstacle avoidance and velocity constraints: a double integral TTCBLF approach. Applied Mathematics and Computation, 511 (??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004643.

Fome:2025:ACR

[FBK25] Anna Daniel Fome, Wolfgang Bock, and Axel Klar. Analysis of a competitive respiratory disease system with quarantine: Epidemic thresholds and crossimmunity effects. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004296.

Forrier:2025:NLB

[FGJ25] Philip P. Forrier, Joan Gimeno, and Àngel Jorba. A note on the local behavior of the Taylor method for stiff ODEs. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000712.

Fusi:2025:UOD

[FGR25] L. Fusi, A. Giovinetto, and K. R. Rajagopal. Unsteady one dimensional motions of a new class of seemingly viscoplastic materials. Applied Mathematics and Computation, 497(??): ??, July 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000931.

Flaszczynska:2025:NSV

[FGW25] Anna Flaszczyńska, Aleksandra Gorzkowska, and Mariusz Woźniak. A note on sequences variant of irregularity strength for hypercubes. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000396.

Farhad:2026:CSS

[FHX26] Qaisar Farhad, Mumtaz Hussain, and Shou-Jun Xu. Closeness spectra and structural uniqueness of special graph classes. Applied Mathematics and Computation, 509(??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003923.

Fermo:2025:GAM

[FLLR25] Luisa Fermo, Anna Lucia Laguardia, Concetta Laurita, and Maria Grazia Russo. A global approximation method for secondkind nonlinear integral equations. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005551.

Fang:2026:IAS

[FLSS26] Zhi-Wei Fang, Tian-Yi Li, Hai-Wei Sun, and Tao Sun. Improved analysis of sine-transform-based approximate inverse preconditioner for Toeplitz-like systems. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003352.

Flamarion:2025:FIP

[FP25] Marcelo V. Flamarion and Efim Pelinovsky. Features of the interaction of paired solitary waves with the Cubic Vortical Whitham equation. Applied Mathematics and Computation, 493(??):

??, May 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007264.

Fang:2025:NCH

[FT25] Yu Fang and Yingzhi Tian.

Neighbor connectivity of hierarchical pancake networks. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003297.

Fan:2025:ECS

[FWLZ25] Xue Fan, Lidong Wang, Xuesong Liu, and Anhao Zheng. Exemplary cooperation strategy with positive influence enhances cooperation quality in social dilemma. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001249.

Fan:2025:IOR

[FY25] Ziyao Fan and Chen Yang.

Interval-oriented reducedorder model for uncertain control systems. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300325000268.

Fu:2025:MEC

[FZY^{+}25] Longxiang Fu, Wanting Zhu, Bo Yu, Yaoyao Zhang, Pedro Antonio Valdes-Sosa, Chunbiao Li, Leonardo Ricci, Mattia Frasca, and Ludovico Minati. Modeling and experimental circuit implementation of fractional single-transistor chaotic oscillators. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001651.

Gao:2025:CTA

[Gao25] Wen-Biao Gao. Correlation theorem and applications associated with the fractional Fourier transform in polar coordinates. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000645.

Ghebleh:2025:GHT

[GAYKS25] Mohammad Ghebleh, Salem Al-Yakoob, Ali Kanso, and Dragan Stevanović. Graphs having two main eigenvalues and arbitrarily many distinct vertex degrees. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000384.

Gurubachan:2025:AFF

[GCV25] Gurubachan, V. V. M. S.

Chandramouli, and S. Verma. Analysis of α-fractal functions without boundary point conditions on the Sierpiński gasket. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005332.

Gao:2026:FTO

[GCW^{+}26] Dong-Xiang Gao, Wen-Hua Cui, Li-Bing Wu, Yu-Jun Zhang, and Ye Tao. Fixed-time optimized control for nonlinear strictfeedback systems based on reinforcement learning and disturbance observer. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003546.

Gottwald:2026:SPK

[GED26] Kurt Klement Gottwald, Snjezana Majstorović Ergotić, and Tomislav Doslić. Soltés problem for the Kirchhoff index of a graph. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004205.

Grad:2026:ENM

[GK26] Simon Grad and Sandi Klavzar. Enumerating the number of k-matchings in successively amalgamated graphs. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004291.

Gupta:2025:SRH

[GKM25] Saloni Gupta, Arnab Kayal, and Moumita Mandal. Superconvergence results for hypersingular integral equation of first kind by Chebyshev spectral projection methods. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032400554X.

Gartner:2025:OSV

[GKR^{+}25] Bernd Gärtner, Vishwas Kalani, Meghana M. Reddy, Wouter Meulemans, Bettina Speckmann, and Miloš Stojaković. Optimizing symbol visibility through displacement. Applied Mathematics and Computation, 505 (??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002553.

Gong:2025:RBT

[GLD25] Luozhong Gong, Weijun Liu, and Shaojun Dai. The reduction of blocktransitive 3 − (v, k, 2) designs. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001626.

Geng:2025:ECA

[GLHS25] Yini Geng, Yikang Lu, Lijun Hong, and Lei Shi. Evolution of cooperation with asymmetric rewards. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005368.

Guo:2026:DNN

[GLY26] Shanshan Guo, Jinghao Li, and Guang-Hong Yang. Deep neural network-based adaptive supervisory control for strict-feedback nonlinear systems with sensor and actuator faults. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004709.

Grzelec:2026:NPA

[GMOS26] Igor Grzelec, Tomás Madaras, Alfréd Onderko, and Roman Soták. On a new problem about the local irregularity of graphs. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004886.

Gao:2025:EEC

[GPH25] Liyan Gao, Qiuhui Pan, and Mingfeng He. The effect of ecological compensation on a coevolutionary common pool resource game. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006696.

Guo:2025:DED

[GPL25] Siyu Guo, Yingnan Pan, and Hongyi Li. Dynamic event-driven optimal consensus control for stateconstrained multiagent zerosum differential graphical games. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004405.

Grzelak:2026:RAD

[Grz26] Lech A. Grzelak. On randomization of affine diffusion processes with application to pricing of options on VIX and SP 500. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003248.

Garima:2026:RNA

[GSB26] Garima, Kapil K. Sharma, and Komal Bansal. Robust numerical analysis for 2D singularly perturbed elliptic delay differential equations exhibiting interior layers due to singularity propagation. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004953.

Gao:2025:CUE

[GSL25] Shiping Gao, Jinghui Suo, and Nan Li. Cooperation under endogenous punishment in the spatial public goods game. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006179.

Guo:2025:STR

[GVW25] Yubao Guo, Lutz Volkmann, and Yun Wang. Signed total Roman domination and domatic numbers in graphs. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005356.

Gao:2025:WSS

[GX25] Jingjing Gao and Xiangpeng Xie. A weighted switching sequence optimization algorithm for static output feedback control synthesis of nonlinear systems. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006131.

Gao:2025:RDR

[GXY25] Zhipeng Gao, Changqing Xi, and Jun Yue. The restrained double Roman domination and graph operations. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000426.

Guo:2026:TSP

[GZL^{+}26] Wenxiu Guo, Hua Zheng, Xiaoping Lu, Yongxiong Zhang, and Seakweng Vong. A two-stage parallel iteration method for nonlinear complementarity problems. Applied Mathematics and Computation, 512(??): ??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325005028.

Hooshyarfarzin:2025:NSE

[HAD25] Baharak Hooshyarfarzin, Mostafa Abbaszadeh, and Mehdi Dehghan. Numerical simulation and error estimation of the Davey– Stewartson equations with virtual element method. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006076.

Haghighi:2026:CPP

[Hag26] Amir Haghighi. Convergence of a positivity preserving logarithmic truncated EM method for SDEs with discontinuous drift coefficients. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004308.

Hainaut:2026:AOP

[Hai26] Donatien Hainaut. American option pricing with model constrained Gaussian process regressions. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004722.

Hanna:2026:MBE

[Han26] Magdy Tawfik Hanna. Matrixbased evaluation of the fractional Hankel transform by Bessel series expansion. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003972.

Huang:2026:IED

[HC26] Yijie Huang and Yanhong Chen. Investigating the evolutionary dynamics of cooperation on a dynamic network: the influence of aspiration-driven structural adaptations. Applied Mathematics and Computation, 509(??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003960.

Huang:2025:CPD

[HCCH25] Xianbin Huang, Liming Chen, Wangyong Chen, and Yao Hu. Change point detection in temporal networks based on graph snapshot similarity measures. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400626X.

Hu:2026:ETP

[HCQ26] Yuanyang Hu, Jiaqi Chen, and Chunbin Qin. Eventtriggered prescribed performance control of the multiplayer game nonlinear system via integral reinforcement learning. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004424.

Hao:2025:NIB

[HCTS25] Tianchu Hao, Yaming Chen, Lingyan Tang, and Songhe Song. A nonlinear immersed boundary method for weighted compact nonlinear schemes. Applied Mathematics and Computation, 499 (??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001377.

Hauenstein:2025:BPH

[HHSW25] Jonathan D. Hauenstein, Caroline Hills, Andrew J. Sommese, and Charles W. Wampler. Branch points of homotopies: Distribution and probability of failure. Applied Mathematics and Computation, 493(??): ??, May 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007343.

He:2025:NSM

[HJF25] Linshuang He, Luru Jing, and Minfu Feng. New stabilized mixed finite element methods for two-field poroelasticity with low permeability. Applied Mathematics and Computation, 494 (??):??, June 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000128.

Hou:2025:ETF

[HLT25] Yingxue Hou, Yan-Jun Liu, and Shaocheng Tong. Eventtriggered fault-compensation-based fuzzy finite-time FTC for MIMO switched nonlinear systems. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000438.

Huang:2025:SHS

[HLZ25] Yueli Huang, Yuan Lv, and Quanxin Zhu. Stabilization of hybrid stochastic neutral-type systems with non-differentiable delays under high nonlinearity via discrete-time feedback control. Applied Mathematics and Computation, 495(??): ??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000608.

Han:2025:ESM

[HRS25] Qiansheng Han, Antti Rasila, and Tommi Sottinen. Efficient simulation of mixed boundary value problems and conformal mappings. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005800.

Hosseinzadeh:2025:RBE

[HS25] Hossein Hosseinzadeh and Zeinab Sedaghatjoo. Radial boundary elements method, a new approach on using radial basis functions to solve partial differential equations, efficiently. Applied Mathematics and Computation, 493(??):??, May 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007136.

Hou:2025:PFT

[HWC25] Xinzhao Hou, Huaiqin Wu, and Jinde Cao. Practical finite-time synchronization for Lur’e systems with performance constraint and actuator faults: a memorybased quantized dynamic event-triggered control strategy. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005691.

Hu:2025:FTC

[HWS25] Zhiquan Hu, Jie Wang, and Changlong Shen. A Fan-type condition for cycles in 1-tough and kconnected (P_{2} ∪ kP_{1})-free graphs. Applied Mathematics and Computation, 494 (??):??, June 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500027X.

Hu:2025:EBP

[HY25] Xiaoyan Hu and Dan Ye. Event-based privacyconsensus control for multiagent systems: an enhanced framework for faulttolerance and protection against eavesdropping. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001730.

Han:2025:LRQ

[HYK^{+}25] Juan Han, Liqiao Yang, Kit Ian Kou, Jifei Miao, and Lizhi Liu. Low-rank quaternion matrix completion based on approximate quaternion SVD and sparse regularizer. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400691X.

Huang:2025:EII

[HZ25] Jing Huang and Minjie Zhang. On the eccentricity inertia indices of chain graphs. Applied Mathematics and Computation, 493 (??):??, May 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400732X.

Huang:2026:CSF

[HZ26a] Jianping Huang and Huacheng Zhou. Causal state feedback representation for infinite horizon LQ problems of fractional systems. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004187.

Huang:2026:MNM

[HZ26b] Zejun Huang and Xinwei Zhang. The maximum number of maximum dissociation sets in potted graphs. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003443.

Huang:2025:PBA

[HZF^{+}25] Ming Huang, Si Qi Zhang, Yong Xiu Feng, Jin Long Yuan, and Hong Han Bei. A proximal bundle approach for solving the generalized variational inequalities with inexact data. Applied Mathematics and Computation, 491(??):??, April 15, 2025.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005484.

Iqbal:2025:TLM

[IAK25] Azhar Iqbal, Shahbaz Ahmad, and Junseok Kim. Two-level method for blind image deblurring problems. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004697.

Iman:2025:DDP

[IJM25] Sara Iman and Mohammad-Reza Jahed-Motlagh. Datadriven predictor of controlaffine nonlinear dynamics: Finite discrete-time bilinear approximation of Koopman operator. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005290.

Izgi:2025:HMN

[IÖÜP25] Burhaneddin Izgi, Murat Özkaya, Nazim Kemal Üre, and Matjaz Perc. A holistic matrix norm-based alternative solution method for Markov reward games. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400585X.

Ilyas:2025:ISP

[ISC25] Asim Ilyas and Stefano Serra-Capizzano. Inverse source problems for identifying time and spacedependent coefficients in a 2D generalized diffusion equation. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003236.

Jiang:2025:FDA

[JCHW25] Yubing Jiang, Hu Chen, Chaobao Huang, and Jian Wang. A fully discrete GL-ADI scheme for 2D time-fractional reactionsubdiffusion equation. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324006088.

Jin:2026:QSM

[JCW26] Xin Jin, Xiaojie Chen, and Zhengxin Wang. Quasisynchronization of multiplex networks under deception attacks via eventtriggered impulsive protocol. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003613.

Jodlbauer:2025:AOD

[JDES25] Herbert Jodlbauer, Matthias Dehmer, and Frank Emmert-Streib. Aspects on the optimal decoupling point setting problem. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000700.

Jarolimova:2025:DNS

[JH25] Alena Jarolímová and Jaroslav Hron. Determination of Navier’s slip parameter using data assimilation. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006180.

Jia:2025:NMO

[JK25] Yingying Jia and Hongchao Kang. Numerical methods of oscillatory Bessel transforms with algebraic and Cauchy singularities. Applied Mathematics and Computation, 505(??):??, November 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002498.

Ju:2025:CAT

[JLG25] Ran Ju, Jing Jian Li, and Yang Gao. Cubic s-arc-transitive bi-Cayley graphs. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001274.

Jia:2026:LHC

[JLJZ26] Xuanhe Jia, Fengxia Liu, Baiyun Ji, and Zihang Zhao. The list r-hued coloring of P_{5}-free graph. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004679.

Jiao:2025:JSC

[JLZ25] Yujian Jiao, Tingting Li, and Zhongqiang Zhang. Jacobi spectral collocation method of space-fractional Navier–Stokes equations. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005721.

Jing:2025:OOD

[JS25] Ruixiao Jing and Yuefang Sun. One-to-one disjoint path covers in digraphs with faulty edges. Applied Mathematics and Computation, 493(??):??, May 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324007318.

Jiang:2025:NFN

[JSZ25] Taijiang Jiang, Qiang Sun, and Chao Zhang. A note on the 1-factorization of nonuniform complete hypergraph. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004685.

Jiang:2025:LTN

[JT25] Huifa Jiang and Emran Tohidi. Long time H^{1}- norm stability and suboptimal convergence of a fully discrete L2-version compact difference method on nonuniform mesh grids to solve fourth-order subdiffusion equations. Applied Mathematics and Computation, 501(??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001924.

Ji:2025:SSN

[JXZ^{+}25] Huihui Ji, Shengyuan Xu, He Zhang, Chenlong Li, and Yuhua Xu. Stochastic stability of networked control systems with packet disordering via a gain-scheduling controller. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000098.

Jain:2025:HDG

[JY25a] Riya Jain and Sangita Yadav. Hybridizable discontinuous Galerkin method for nonlinear hyperbolic integro-differential equations. Applied Mathematics and Computation, 498 (??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001201.

Jiang:2025:CRD

[JY25b] Yaxin Jiang and Yujun Yang. Computation of resistance distances and Kirchhoff indices for two classes of graphs. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000815.

Jia:2025:RLB

[JYLS25] Guolong Jia, Qing Yang, Jinxu Liu, and Hao Shen. Reinforcement learning-based linear quadratic tracking control for partially unknown Markov jump singular interconnected systems. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006908.

Jin:2025:VCB

[JZ25] Zhixiang Jin and Chengjian Zhang. Variable-coefficient BDF methods with fullygeometric grid for linear nonhomogeneous neutral pantograph equations. Applied Mathematics and Computation, 499(??):??, August 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001390.

Jiang:2025:AAT

[JZGV25] Tongsong Jiang, Dong Zhang, Zhenwei Guo, and V. I. Vasil’ev. An algebraic algorithm for the total least squares problem in commutative quaternionic theory. Applied Mathematics and Computation, 494(??): ??, June 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400729X.

Kabir:2025:MIP

[Kab25] K. M. Ariful Kabir. Mitigating infection in prey through cooperative behavior under fear-induced defense. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000451.

Kamyczura:2026:IMA

[Kam26] Mateusz Kamyczura. Introduction to majority additive coloring. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500339X.

Kunstek:2025:MCV

[KBM25] Petar Kunstek, Martina Bukac, and Boris Muha. Mass conservation in the validation of fluid-poroelastic structure interaction solvers. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005423.

Kumar:2026:SRC

[KC26] Ajay Kumar and V. V. M. S. Chandramouli. Study of reduced Chialvo map with electromagnetic flux: Dynamics and network behavior. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003765.

Ke:2025:QMC

[KCLZ25] Yifen Ke, Xiaomin Cai, Riwei Liao, and Huai Zhang. Quaternion modified conjugate gradient algorithm to solve Sylvester-type quaternion matrix equations with generalized coupled form as well as application. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000578.

Kunnarath:2025:LSA

[KGJ25] Ajil Kunnarath, Santhosh George, and P. Jidesh. Local and semilocal analysis of a class of fourth order methods under common set of assumptions. Applied Mathematics and Computation, 505(??):??, November 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002528.

Krellner:2025:WWH

[KH25] Marcus Krellner and The Anh Han. Words are not wind

— how public joint commitment and reputation solve the Prisoner’s Dilemma. Applied Mathematics and Computation, 506(??):??, December 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002565.

Kattan:2026:HIC

[KH26] Doha A. Kattan and Hasanen A. Hammad. Hemivariational inequalities and controllability results for second-order non-autonomous evolution system with impulsive effects. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300325004758.

Kishor:2026:FCA

[Kis26] Avadh Kishor. On the fairness and cooperation among free-riders. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004497.

Keng:2025:CPL

[KK25a] Ying Ying Keng and Kiam Heong Kwa. Contagion probability in linear threshold model. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005514.

Kim:2025:SID

[KK25b] Taekyun Kim and Dae San Kim. Some identities on degenerate harmonic and degenerate higher-order harmonic numbers. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400506X.

Kim:2025:RCP

[KL25] Do Wan Kim and Donghwan Lee. Relaxed conditions for parameterized linear matrix inequality in the form of nested fuzzy summations. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400540X.

Kolbe:2025:RAC

[KM25] Niklas Kolbe and Siegfried Müller. A relaxation approach to the coupling of a two-phase fluid with a linear-elastic solid. Applied Mathematics and Computation, 504(??):??, November 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002292.

Kim:2026:NMM

[KMK26] Soo Hyun Kim, Hwan Pyo Moon, and Song-Hwa Kwon. A novel method for manipulating polynomial curves by the Gauss–Legendre control polygon with points interpolating property. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300325004850.

Khodaverdian:2025:FTM

[KNKR25] Maria Khodaverdian, Majdeddin Najafi, Omid Kazemifar, and Shahabuddin Rahmanian. Fault-tolerant model predictive sliding mode control for trajectory replanning of multi-UAV formation flight. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005344.

Kovacevic:2025:VMM

[Kov25] Mladen Kovacević. Vector multispaces and multispace codes. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005022.

Knor:2025:EAG

[KSFD25] Martin Knor, Riste Skrekovski, Slobodan Filipovski, and Darko Dimitrov. Extremizing antiregular graphs by modifying total σirregularity. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400660X.

Kassid:2025:CAN

[KSR25] Qusay Abdulraheem Kassid, Saeed Sohrabi, and Hamid Ranjbar. Convergence analysis of a Nyströmtype method for a class of nonlinear integral equations with highly oscillatory kernels. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001778.

Knor:2026:ISN

[KSSY26] Martin Knor, Jelena Sedlar, Riste Skrekovski, and Yu Yang. Invitation to the subpath number. Applied Mathematics and Computation, 509(??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003728.

Korze:2025:VMV

[KV25] Danilo Korze and Aleksander Vesel. Variety of mutual-visibility problems in hypercubes. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006799.

Kang:2025:FAA

[KXL^{+}25] Hongchao Kang, Chunzhi Xiang, Guidong Liu, Ao Liu, and Xinrui Hou. Fast algorithms and analysis of oscillatory and weakly singular generalized Bessel transforms. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006672.

Liu:2025:NMO

[LBLZ25] Jie Liu, Driss Boutat, Da-Yan Liu, and Xue-Feng Zhang. Nonlinear MIMO observable normal forms with output injection and output diffeomorphism. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006350.

Liu:2025:FDE

[LC25] Xin-Yu Liu and Wei-Wei Che. Fully distributed eventtriggered secondary voltage resilient tracking control for microgrids. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006453.

Li:2025:TDC

[LCL25] Hao Li, Liting Chen, and Mei Lu. Two-disjoint-cycle-cover pancyclicity of splitstar networks. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005460.

Li:2025:SOA

[LCLY25] Xinying Li, Chao Chang, Jianze Li, and Yuning Yang. A sparse optimization approach for simultaneous orthogonal tensor diagonalization. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006647.

Liu:2025:TVG

[LCRZ25] Linlin Liu, Liang Cao, Hongru Ren, and Meng Zhao. Time variable gain ESObased fixed-time platoon tracking control for connected vehicles. Applied Mathematics and Computation, 502(??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002036.

Lee:2026:CSN

[LCS26a] Hsuan-Wei Lee, Colin Cleveland, and Attila Szolnoki. Context-sensitive norm enforcement reduces sanctioning costs in spatial public goods games. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003522.

Li:2026:ECK

[LCS^{+}26b] Xiaobin Li, Meng Chen, Zhengjie Sun, Leevan Ling, and Siqing Li. Energyconserving Kansa methods for Hamiltonian wave equations. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004084.

Li:2025:DAS

[LCW25] Baomin Li, Mou Chen, and Min Wan. Dual auxiliary systems-based turn-on/off trigger fault-tolerant control for input-constrained satellite. Applied Mathematics and Computation, 506 (??):??, December 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002747.

Cen:2025:STF

[lCXZ^{+}25] Bing ling Cen, Yu Xue, Kun Zhang, Lisi Jia, and Hong di He. Study on traffic flows with connected vehicles and human-driven vehicles. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400643X.

Li:2025:ASE

[LD25] Xiaolin Li and Haiyun Dong. Analysis of a stabilized element-free Galerkin method for magnetohydrodynamic flow at very large Hartmann numbers. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500061X.

Liu:2026:RBF

[LD26] Tao Liu and Bolin Ding. A radial basis function neural network approach for solving a diffusion partial differential equation efficiently. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003777.

Lu:2025:ISP

[LdMAS25] Yikang Lu, Alfonso de Miguel-Arribas, and Lei Shi. Influence of social peers on vaccine hesitancy under imperfect vaccination.

Applied Mathematics and Computation, 491(??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006751.

Lu:2025:EPR

[LDT^{+}25] Yikang Lu, Hui Dai, Huaiyu Tan, Xiaofang Duan, Lei Shi, and Junpyo Park. Enhancement of persistence in the rock–paper–scissors dynamics through higherorder interactions. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005447.

Li:2026:QDC

[LDXL26] Teng-Fei Li, Liming Ding, Jun Xiong, and Haibing Li. Quantized dissipative control for fuzzy hot strip mill cooling system under input constraint via dynamic output feedback approach. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003418.

Liang:2026:ERB

[LF26] Rongqian Liang and Suohai Fan. Environmental reputation-based preference selection and strategy learning mechanisms promote cooperation. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004862.

Liu:2026:DTS

[LGW^{+}26] Lun Liu, Shaorui Geng, Lilin Wei, Zhenyong Lu, and Yinghong Ma. Dynamics of two-strain disease transmission with negative vaccine information and vaccination on multiplex networks. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004278.

Li:2025:ETV

[LH25] Ming Li and Liang’an Huo. Effect of time-varying validity of individual interaction on co-evolution of awareness and epidemics in a multiplex high-order network. Applied Mathematics and Computation, 499 (??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001237.

Li:2025:EIP

[LHJ^{+}25] Yipeng Li, Xiangyue Hu, Xing Jin, Huizhen Zhang, Jiajia Yang, and Zhen Wang. Environmental information perception enhances cooperation in stochastic public goods games via Qlearning. Applied Mathematics and Computation, 504 (??):??, November 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002310.

Li:2026:SRS

[LHL^{+}26] Peng-Yue Li, Feng Hu, Fa-Xu Li, You-Feng Zhao, and Yu-Rong Song. SEIAR rumor spreading model with antagonistic states in hypernetworks. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003637.

Li:2025:EBG

[LHLL25] Qidong Li, Changchun Hua, Kuo Li, and Hao Li. Eventbased global asymmetric constraint control for pnormal systems with unknown control coefficients. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000025.

Li:2025:RLB

[LHZL25] Xin Li, Dakuo He, Qiang Zhang, and Hailong Liu. Reinforcement learning-based adaptive event-triggered control of multi-agent systems with time-varying deadzone. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005204.

Li:2026:DME

[LHZL26a] Xin Li, Dakuo He, Qiang Zhang, and Hailong Liu. Dynamic memory eventtriggered interval type-2 fuzzy RBFNN control for multi-agent systems under false data injection attacks. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003200.

Li:2026:SOD

[LHZL26b] Yan Li, Qirun Hang, Lijuan Zha, and Jinliang Liu. Secure observer-driven synchronization for complex networks constrained by partial information exchange and limited communication bandwidth. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004199.

Li:2026:GSM

[Li26]  Haibo Li.  Generalizing

the SVD of a matrix under nonstandard inner product and its applications to linear ill-posed problems. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003340.

Liu:2025:ECN

[Liu25] Qiong Liu. The equivalent conditions for norm of a Hilbert-type integral operator with a combination kernel and its applications. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400537X.

Li:2025:SGF

[LJ25] Jiajun Li and Ying Jiang. Stable generalized finite element for two-dimensional and three-dimensional nonhomogeneous interface problems. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005174.

Lei:2025:SER

[LL25a] Xingyu Lei and Shuchao Li. Spectral extremal results on the A_{α}-spectral radius of graphs without K_{a,b}- minor. Applied Mathematics and Computation, 492 (??):??, May 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006933.

Liu:2025:PBE

[LL25b] Guixiu Liu and Bing Li. Protocol-based H_{∞} estimation for Markovian jumping delayed systems with partially unknown transition probability. Applied Mathematics and Computation, 492(??):??, May 1, 2025.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007082.

Liu:2025:DBP

[LL25c] Wei Liu and Kai Li. Decoupled bound-preserving algorithms for compressible Darcy–Brinkman flow with advection–diffusion transport problem in fractured media. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000256.

Liu:2025:PEP

[LL25d] Yandi Liu and Yonghui Li. A payoff equality perspective for evolutionary games: Mental accounting and cooperation promotion. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005009.

Li:2026:LBR

[LL26a] Zongcheng Li and Jin Li.

Linear barycentric rational collocation method for solving equal width equation. Applied Mathematics and Computation, 512(??): ??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500493X.

Lu:2026:IEE

[LL26b] Wen Lu and Shu Liang. Integrating emotion and expectation improves cooperation. Applied Mathematics and Computation, 511 (??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004618.

Ling:2025:SCI

[LLFS25] Ting Ling, Zhang Li, Minyu Feng, and Attila Szolnoki. Supervised cooperation on interdependent public goods games. Applied Mathematics and Computation, 492 (??):??, May 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324007100.

Liu:2025:EDA

[LLLH25] Fangyuan Liu, Guobao Liu, Huai Liu, and Shicheng Huo. Extended dissipative analysis of integral sliding mode control for singular interval type-2 fuzzy systems with time delay. Applied Mathematics and Computation, 493(??):??, May 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007124.

Luo:2025:LRR

[LLLZ25] Hui Luo, Xin Liu, Wei Liu, and Yang Zhang. Low-rank reduced biquaternion tensor ring decomposition and tensor completion. Applied Mathematics and Computation, 504 (??):??, November 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002309.

Li:2025:CRP

[LLMR25] Lin Li, Jingyu Lv, Lili Ma, and Jing Ruan. Community reciprocity promotes cooperation in a network evolutionary prisoner’s dilemma game. Applied Mathematics and Computation, 502 (??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002231.

Lanza:2025:SFV

[LLMT25] Alessandro Lanza, Antonio Leaci, Serena Morigi, and Franco Tomarelli. Symmetrised fractional variation with L^{1} fidelity for signal denoising via Grünwald– Letnikov scheme. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001560.

Li:2025:PGP

[LLQ25] Tingting Li, Fang Li, and Huiqing Qi. A piecewise gradient prior for small structures and contrast preserving image smoothing. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002838.

Lv:2026:EGB

[LLS^{+}26] Rongjian Lv, Hua Li, Qiubai Sun, Qiyue Gao, and Bowen Li. Evolution of group behaviour in the ecological snowdrift game-level-headed and impatient delayed epidemic model under emergencies. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003947.

Long:2026:CNF

[LLW26] Jiali Long, Kaiyang Lan, and Yan Wang. The chromatic number of {P_{2} ∪ P_{3}, banner}-free graphs. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004333.

Liu:2026:IBC

[LLWC26] Aixin Liu, Haitao Li, Lin Wang, and Guanrong Chen. Incentive-based control of networked evolutionary coordination games under asynchronous bestresponse dynamics. Applied Mathematics and Computation, 509(??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004102.

Li:2025:OTN

[LLX25] Jin Li, Yongxin Lan, and Changqing Xu. The outerplanar Turán number of disjoint copies of paths. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000232.

Liu:2026:AHF

[LLZ26] Shuqi Liu, Huiqing Liu, and Qin Zhao. Analogy of hybrid fault diagnosis capability on theta-free connected graphs. Applied Mathematics and Computation, 512 (??):??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325005004.

Li:2026:TVD

[LLZL26] Meilin Li, Yue Long, Kai Zhang, and Tieshan Li. Time-varying dynamic eventtriggered-based prescribedtime leader-following consensus for multi-agent systems with input saturation. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500356X.

Lotstedt:2025:NIM

[LM25] Per Lötstedt and Sonja Mathias. Numerical integration of mechanical forces in center-based models for biological cell populations. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005307.

Liang:2025:HML

[LMMG25] Ce Liang, Weiyuan Ma, Chenjun Ma, and Ling Guo. Harnessing machine learning for identifying parameters in fractional chaotic systems. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500181X.

Li:2025:NAR

[LMQZ25] Rui Li, Yuede Ma, Zhongmei Qin, and Yingping Zhao. Note on the anti-Ramsey number for matching in hypercubes. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006155.

Li:2025:HDM

[LSF^{+}25] Xuan Li, Jin Su, Jin-Qian Feng, Xiong Lei, and Rui-Bo Zhang. A hybrid dynamic mode decomposition algorithm combining random and sparsity promoting and its application to viscoelastic flow around circular cylinder. Applied Mathematics and Computation, 505 (??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002346.

Liu:2025:AFC

[LSH25] Yanli Liu, Yihua Sun, and Li-Ying Hao. Adaptive FTPP control of switched stochastic nonlinearly parameterized systems with asymptotic tracking performance. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000323.

Li:2025:QSS

[LSP^{+}25] Hanfei Li, Jiankun Sun, Dong Pan, Huchen Luo, and Sen Li. Quasisynchronization of stochastic delayed multi-agent systems via gain-waving irregular intermittent control with its application in circuits. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001729.

Li:2025:GBN

[LSQ25] Yinkui Li, Guiyu Shi, and Xiaoxiao Qin. Graphs with burning number three. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005617.

Liu:2025:NCE

[LSWZ25] Fenjin Liu, Ke Su, Wei Wang, and Hao Zhang. Non-cospectral equienergetic trees of diameter at most four. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005654.

Laadhari:2025:EFE

[LT25] Aymen Laadhari and Helmi Temimi. Efficient finite element strategy using enhanced high-order and second-derivative-free variants of Newton’s method. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005198.

Luo:2025:LSP

[LTH^{+}25] Fangwen Luo, Qiong Tang, Yiting Huang, Yanhui Ding, and Sijia Tang. Local structure-preserving algorithms for the nonlinear Schrödinger equation with power law nonlinearity. Applied Mathematics and Computation, 484(??): ??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004478.

Liu:2025:PSS

[LW25a] Yang-Fan Liu and Huai-Ning Wu. Probabilistic stability and stabilization of humanmachine system via hidden semi-Markov modeling approach. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006143.

Lu:2025:EDL

[LW25b] Shounan Lu and Yang Wang. Experience-driven learning and interactive rules under link weight adjustment promote cooperation in spatial prisoner’s dilemma game. Applied Mathematics and Computation, 497(??):??, July 15, 2025. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300325001080.

Lu:2025:PPD

[LW25c] Shounan Lu and Yang Wang. Past-performance-driven strategy updating promote cooperation in the spatial prisoner’s dilemma game. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006817.

Luo:2025:CPL

[LW25d] Huiping Luo and JinRong Wang. Controllability of piecewise linear state-delay systems. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000086.

Li:2025:EFP

[LWH^{+}25] Minlan Li, Chao Wang, Yanyan Han, Si-Yi Wang, and Ruiwu Wang. Environmental feedback promotes cooperation in a spatial prisoner’s dilemma game with preferential selection. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500058X.

Li:2025:RMM

[LWL^{+}25] Kun Li, Xia Wang, Fei Lu, Zhe Zhang, and Xiaodi Sun. Research on monitoring mechanism of autonomous taxi: an evolutionary game approach. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003224.

Liu:2026:CFG

[LWLZ26] Qianqian Liu, Jingfeng Wang, Chunmei Li, and Heping Zhang. Components of flip graphs of domino tilings in quadriculated cylinder and torus. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004230.

Liu:2026:SOR

[LWML26] Yonghe Liu, Fengsheng Wu, Jifei Miao, and Chaoqian Li. A subspace-orbit randomized algorithm for quaternion tensor singular value decomposition based on Qtproduct. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325005053.

Liu:2025:DOB

[LWS^{+}25] Ming-Rui Liu, Li-Bing Wu, Hong Sang, Liang-Dong Guo, and Sheng-Juan Huang. Disturbance observer-based fuzzy adaptive switched tuning control of uncertain nonlinear systems with full state constraints. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400571X.

Li:2025:ESS

[LWY25] Xing-Yu Li, Kai-Ning Wu, and Zhan-Wen Yang. Exponential stabilization for spatial multiple-fractional advection–diffusion–reaction system. Applied Mathematics and Computation, 499 (??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001365.

Li:2025:NSE

[LX25] Yuying Li and Kexiang Xu. On the number of F_{k,4}-saturating edges. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006234.

Li:2025:BUT

[LXM25] Jun Li, Kailiang Xin, and Lingsheng Meng. A block upper triangular splitting method for solving block three-by-three linear systems arising from the large indefinite least squares problem. Applied Mathematics and Computation, 505 (??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002723.

Lv:2025:DAM

[LYD25] Yuan-Wei Lv, Guang-Hong Yang, and Georgi Marko Dimirovski. Distributed adaptive moving horizon estimation for multi-sensor networks subject to quantization effects. Applied Mathematics and Computation, 488(??):??, March 1, 2025.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005873.

Liu:2025:FDF

[LYLS25] Xuelong Liu, Guoju Ye, Wei Liu, and Fangfang Shi. Fuzzy discrete fractional granular calculus and its application to fractional cobweb models. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006374.

Li:2025:MLL

[LYM25] Dan Li, Minghui Yan, and Jixiang Meng. Minimizing the least Laplacian eigenvalue of signed complete graphs. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004636.

Li:2026:NAS

[LYQZ26] Qimin Li, Yubin Yan, Leijie Qiao, and Yu Zhang. Numerical approximation for a stochastic time-fractional cable equation. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004357.

Li:2026:ECS

[LYZ^{+}26] Xiaopeng Li, Xuexue Yang, Xiuli Zhang, Weiwei Han, and Yan Xu. Evolution of cooperation in spatial heterogeneous social dilemmas with dynamic self-adaptive matching. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003327.

Li:2025:SEG

[LZ25a] Shuchao Li and Zihan Zhou. On spectral extrema of graphs with given order and generalized 4-independence number. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400479X.

Liu:2025:RUL

[LZ25b] Zhe Liu and Qingyuan Zhang. Remaining useful life estimation considering threshold epistemic uncertainty with uncertain differential equation. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000670.

Liu:2026:SAS

[LZ26] Yicheng Liu and Quanxin Zhu. Stability analysis of stochastic McKean– Vlasov equations with discrete observation. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004655.

Li:2025:NBT

[LZCD25] Shangzhao Li, Xiran Zhang, Jing Chen, and Shaojun Dai. A note on blocktransitive 2 − (v, k, 1) designs. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000414.

Li:2025:OHB

[LZD^{+}25] Kun Li, PeiYun Zhong, Li Dong, LingMin Wang, and Luo-Luo Jiang. OP-HHO based feature selection improves the performance of depression classification framework: a gender biased multiband research. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500044X.

Li:2025:LBM

[LZJ^{+}25] Shuo Li, Yiyun Zhang, Yongqi Jia, Zhengrong Xiang, and Jinling Wang. LP-based mode-dependent piece-wise event-triggered finite-time L_{1} control for switched positive systems under MDMDT constraint. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000591.

Liu:2025:FTS

[LZP25] Shuo Liu, Huaguang Zhang, and Hongbo Pang. Fixedtime stability criteria and adaptive state constrained control for uncertain switched nonlinear systems. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000724.

Lian:2025:PTD

[LZY^{+}25] Yuxiao Lian, Baoyong Zhang, Deming Yuan, Yao Yao, and Bo Song. Prescribedtime distributed resource allocation algorithm for heterogeneous linear multiagent networks with unbalanced directed communication. Applied Mathematics and Computation, 503 (??):??, October 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002243.

Li:2025:BCS

[LZZ25] Chaoyang Li, Shidong Zhai, and Yuanshi Zheng. Bounded consensus in second-order uncertain nonlinear multiagent systems: a distributed neural network control approach. Applied Mathematics and Computation, 498 (??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000967.

Liu:2026:OBR

[LZZ^{+}26] Mingde Liu, Liang Zhang, Ning Zhao, Xinjun Wang, and Yongchao Liu. Observerbased reachable set synthesis for multi-agent systems: a sampling-based distributed event-triggered strategy. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004096.

Molina-Abril:2025:TSF

[MAMFBM^{+}25] H. Molina-Abril, M. J. Morón-Fernández, M. Benito-Marimón, F. Díaz del Río, and P. Real. Topological scale framework for hypergraphs. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004508.

Mao:2025:ACT

[Mao25]  Renrong Mao.  Asymp-

totics for the coefficients of the truncated theta series. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003182.

Mosic:2026:WWI

[MAS26] Dijana Mosić, Mohamed Amine Aouichaoui, and Predrag S. Stanimirović. Weak m-WG inverse. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003856.

Meier:2025:MFS

[MBF25] Andreas Meier, Eberhard Bänsch, and Florian Frank. Multiscale flow simulations of dilute polymeric solutions with bead-rod chains using Brownian configuration fields. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005526.

Mirza:2025:DRC

[MBH^{+}25] Fuat Kaan Mirza, Tunçer Baykas, Mustafa Hekimoglu, Önder Pekcan, and Gönül Paçaci Tunçay. Decoding rhythmic complexity: a nonlinear dynamics approach via visibility graphs for classifying asymmetrical rhythmic frameworks of Turkish classical music. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002917.

Ma:2025:NMM

[MCH25a] Y. Ma, C. S. Chen, and Y. C. Hon. A novel meshless method for solving long-term evolution problem on irregular domain. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006702.

Mei:2025:ACA

[MCH25b] Xu Mei, Jun Cheng, and Wentao Huang. Attackcompensated asynchronous output feedback control for stochastic switching systems with sojourn probability. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004855.

Ma:2025:IAO

[MCHB25] Liyang Ma, Daniel Conus, Wei-Min Huang, and Paolo Bocchini. Improved algorithm for the optimal quantization of single- and multivariate random functions. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004892.

Munir:2025:ASE

[MCKZ25] Taj Munir, Hongchu Chen, Can Kang, and Hussan Zeb. Analysis of stability, energy conservation, and convergence of coupled advection–diffusion models for ocean–atmosphere interactions. Applied Mathematics and Computation, 502 (??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500222X.

Mu:2025:SSB

[MeFW25] Tiantian Mu, Jun e Feng, and Biao Wang. Set stabilization of Boolean control networks based on bisimulations: a dimensionality reduction approach. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000657.

Mezzadri:2025:MBF

[Mez25]  Francesco Mezzadri.  A

modulus-based framework for weighted horizontal linear complementarity problems. Applied Mathematics and Computation, 495(??): ??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000402.

Moon:2025:IFA

[MH25] Sunghwan Moon and Markus Haltmeier. Inversion formulas for the attenuated conical Radon transform: Plane and cylinder case. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006209.

Mosazadeh:2025:DSL

[MK25] Seyfollah Mosazadeh and Hikmet Koyunbakan. Discrete Sturm–Liouville operators having the hydrogen atom potential. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001547.

Martins:2026:CID

[ML26] Rui Martins and Evgeny Lakshtanov. A C++ implementation of the discrete adjoint sensitivity analysis method for explicit adaptive Runge–Kutta methods enabled by automatic adjoint differentiation and SIMD vectorization. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004254.

Mu:2025:HCP

[MLTL25] Yu Mu, Wing-Cheong Lo, Yuanshun Tan, and Zijian Liu. Hybrid control for the prey in a spatial preypredator model with cooperative hunting and fear effect time lag. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006787.

Mao:2025:ANF

[MLW^{+}25] Jun Mao, Qiang Li, Ronghao Wang, Wencheng Zou, and Zhengrong Xiang. Adaptive neural fixed-time sampleddata output-feedback stabilization for a class of nonlinear systems. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002760.

Miao:2026:SNF

[MLZ26] Shujing Miao, Shuchao Li, and Licheng Zhang. Smith normal form and the generalized spectral characterization of a type of symmetric matrices. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004898.

Mapui:2025:LLC

[MM25] Arnab Mapui and Santwana Mukhopadhyay. Lyapunovlike conditions for prescribedtime stability of perturbed impulsive systems. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324006489.

Ma:2026:ASC

[MMF26] Jian Ma, Yixue Ma, and Qiuxia Fu. Applications of Schur–Cohn matrix and matrix pencil methods in studying the stability of high-dimensional neutral delay differential equations. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003650.

Munoz:2025:ELL

[MMNP25] Francisco J. Muñoz, Luca Meacci, Juan Carlos Nuño, and Mario Primicerio. Exploring the limits of the law of mass action in the mean field description of epidemics on Erdős–Rényi networks. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004806.

Mondal:2025:HPH

[MMTU25] Bapin Mondal, Sayan Mandal, Pankaj Kumar Tiwari, and Ranjit Kumar Upadhyay. How predator harvesting affects preypredator dynamics in deterministic and stochastic environments? Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001079.

Mohammadi:2026:GLS

[MNE26] Abdullah Mohammadi, Jalal A. Nasiri, and Sohrab Effati. Gravitational least squares twin support vector machine based on optimal angle for class imbalance learning. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500431X.

Moon:2025:LED

[MP25] Sunyo Moon and Seungkook Park. Laplacian eigenvalue distribution for unicyclic graphs. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004831.

Malnic:2026:RHW

[MP26] Aleksander Malnic and Rok Pozar. On reduced Hamilton walks. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004217.

Mittal:2025:TLI

[MR25] Kajal Mittal and Rajendra K. Ray. Two-level implicit high-order compact scheme in exponential form for 3D quasilinear parabolic equations. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004454.

Miao:2026:CDO

[MS26] Suoxia Miao and Housheng Su. Consensus of differentorder switched multi-agent systems on matrix-weighted networks. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003364.

Muthusamy:2026:AIO

[Mut26] Vijayakumar Muthusamy.

Augmented input-output finite time stabilization and tracking control for networked control systems via adaptive event-triggered mechanism. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003509.

Meng:2026:MEA

[MW26] Jie Meng and Yuezhi Wang. On the matrix equation X + AX^{−1}B = Q with semiinfinite quasi-Toeplitz coefficients. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003741.

Ma:2025:WMP

[MWS25] Haifeng Ma, Wen Wang, and Predrag S. Stanimirović. Weighted Moore–Penrose inverses for dual matrices and its applications. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006064.

Ma:2025:SMI

[MZX25] Yuanyuan Ma, Qiannan Zhang, and Leilei Xie. A study on multi-information diffusion model considering dual social reinforcement effect from the perspective of evolutionary game. Applied Mathematics and Computation, 492 (??):??, May 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007161.

Nowak:2025:NNA

[NKSB25] Anna Nowak, Dominika Kustal, HongGuang Sun, and Tomasz Blaszczyk. Neural network approximation of the composition of fractional operators and its application to the fractional Euler–Bernoulli beam equation. Applied Mathematics and Computation, 501 (??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002012.

Niu:2025:CSM

[NNW25] Mengyao Niu, Yuanling Niu, and Jiaxin Wei. Convergence and stability in mean square of the stochastic θ-methods for systems of NSDDEs under a coupled monotonicity condition. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001225.

Nogueira:2025:RDD

[NS25] Pedro Nogueira and Ana L.

Silvestre. Regularized directional do-nothing boundary conditions for the Navier– Stokes equations: Analytical and numerical study. Applied Mathematics and Computation, 499(??):??, August 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001250.

Ni:2025:SPU

[NSX^{+}25] Jinyuan Ni, Yin Sheng, Qiang Xiao, Zhigang Zeng, and Tingwen Huang. Stability of partially unstable discrete-time switched positive nonlinear systems with delays. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003170.

Navarro:2026:AOS

[NVGA^{+}26] Juan David Navarro, Juan C. Velasquez-Gonzalez, Mauricio Aristizabal, Arturo Montoya, Harry R. Millwater, and David Restrepo. Arbitrary-order sensitivity analysis of frequency response functions using hypercomplex automatic differentiation and spectral finite elements. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004035.

Ouyang:2026:VEL

[OCN^{+}26] Peichang Ouyang, Kwok Wai Chung, Alain Nicolas, Robert W. Fathauer, David Bailey, and Krzysztof Gdawiec. Visualization of Escher-like hyperbolic tessellations. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004369.

Orizaga:2025:SAC

[OGPFVV25] Saulo Orizaga, Gilberto González-Parra, Logan Forman, and Jesus Villegas-Villanueva. Solving Allen– Cahn equations with periodic and nonperiodic boundary conditions using mimetic finite-difference operators. Applied Mathematics and Computation, 484(??):??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004545.

Palubeckis:2025:FLS

[Pal25] Gintaras Palubeckis. A fast local search based memetic algorithm for the parallel row ordering problem. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005010.

Pandey:2025:CSR

[PBS25] Anamika Pandey, Rahul Barthwal, and T. Raja Sekhar. Construction of solutions of the Riemann problem for a two-dimensional Keyfitz–Kranzer type model governing a thin film flow. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001055.

Pang:2025:CUS

[PC25] Xinhong Pang and Min Chen. Choosability with union separation of planar graphs without intersecting triangles. Applied Mathematics and Computation, 502(??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500219X.

P:2025:CPC

[PDR25] Sreepriya P., Denny K. D., and G. D. Reddy. A class of parameter choice rules for fractional Tikhonov regularization scheme in learning theory. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001742.

Portugal:2025:IUH

[PDV25] Lucas Portugal and Renata Del-Vecchio. Integral uniform hypergraphs. Applied Mathematics and Computation, 504(??):??, November 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002334.

Pan:2025:ETO

[PFZ25] Weihao Pan, Yiyu Feng, and Xianfu Zhang. Eventtriggered output feedback impulsive control for largescale nonlinear time-delay systems with unknown measurement sensitivity. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003315.

Prabhu:2025:MDG

[PJK25] S. Prabhu, T. Jenifer Janany, and Sandi Klavzar. Metric dimensions of generalized Sierpiński graphs over squares. Applied Mathematics and Computation, 505 (??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002541.

Pandit:2026:ENE

[PMVA26] Biswajit Pandit, Pratikshya Mainini, Amit K. Verma, and Ravi P. Agarwal. Existence and non-existence of radial solutions for a class of fourth order elliptic PDE arising in epitaxial growth theory. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003686.

Phan:2025:MME

[PNN25] Kien Trung Phan, Anh Duc Nguyen, and Thuy Phuong Nguyen. A mathematical model of exploitative and interference competitions in stage-structured populations. Applied Mathematics and Computation, 501 (??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001936.

Palav:2025:RFO

[PP25] Mansi S. Palav and Vikas H.

Pradhan. Redefined fourth order uniform hyperbolic polynomial B-splines based collocation method for solving advection-diffusion equation. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004533.

Pang:2025:OBF

[PPC^{+}25] Guochen Pang, Xiang Pan, Xiangyong Chen, Jinde Cao, Yang Liu, and Jianlong Qiu. Observer-based finite-time H_{∞} fault-tolerant control for uncertain Markov jump systems against generally bounded transition probabilities via two-step dynamic event-triggered approach. Applied Mathematics and Computation, 499 (??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001341.

Phu:2026:SFN

[PPH26] Nguyen Dinh Phu, Lai Van Phut, and Ngo Van Hoa. Solving fuzzy non-homogeneous linear differential systems with piecewise constant arguments involving the shortmemory variable-order Caputo fractional derivative. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003984.

Prasina:2025:ESC

[PPN^{+}25a] A. Prasina, V. Samuthira Pandi, W. Nancy, K. Thilagam, K. Veena, and A. Muniyappan. Erratum on “Synchronization and chimeras in asymmetrically coupled memristive Tabu learning neuron network” [Appl. Math. Compt. 489 (2025) 129163]. Applied Mathematics and Computation, 492(??):??, May 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324007094.

See [PPN^{+}25b].

Prasina:2025:SCA

[PPN^{+}25b] A. Prasina, V. Samuthira Pandi, W. Nancy, K. Thilagam, K. Veena, and A. Muniyappan. Synchronization and chimeras in asymmetrically coupled memristive Tabu learning neuron network. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324006246. See er-

ratum [PPN^{+}25a].

Padma:2025:CMM

[PPS25] R. Padma, R. Ponalagusamy, and R. Tamil Selvi. Corrigendum to “Mathematical modeling of electro hydrodynamic non-Newtonian fluid flow through tapered arterial stenosis with periodic body acceleration and applied magnetic field” [Applied Mathematics and Computation, 362 (2019) 124453]. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001456.

See [RRR19, RRR20].

Perez:2025:LCS

[PSM25] R. Pérez, W. Sánchez, and H. J. Martínez. On the local convergence of a smoothing structured secant method for nonlinear complementarity. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005563.

Pankavich:2025:CMT

[PSS^{+}25] Stephen Pankavich, Lucas Schauer, Michael J. Schmidt, Nicholas B. Engdahl, Diogo Bolster, and David A. Benson. Convergence of mass transfer particle tracking schemes for the simulation of advection– diffusion–reaction equations. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000852.

Peyrovan:2025:CAC

[PTB25] P. Peyrovan, A. Tari, and H. Brunner. Convergence analysis of collocation solutions for delay Volterra integral equations with weakly singular kernels. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005836.

Qiao:2026:BRG

[QAS26] Junhe Qiao, Shannon Dee Algar, and Thomas Stemler. Belonging and refueling: Group dynamics in foraging flocks. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004734.

Quan:2025:CSE

[QFNW25] Ji Quan, Yixin Feng, Jiacheng Nie, and Xianjia Wang. Catalyst strategy for the evolution of cooperation in indirect reciprocity. Applied Mathematics and Computation, 497 (??):??, July 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001067.

Qian:2025:DLP

[QHZ25] Yanxia Qian, Yunqing Huang, and Yongchao Zhang. Decoupled, linear and positivitypreserving schemes for a modified phase field crystal system incorporating longrange interactions. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324005502.

Quan:2025:IHP

[QLCW25] Ji Quan, Ran Lv, Shengjin Cui, and Xianjia Wang. The impact of heterogeneous preferences on multi-issue repeated social dilemma games with correlated strategy in structured populations. Applied Mathematics and Computation, 505 (??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002358.

Qiao:2025:DPC

[QM25] Hongwei Qiao and Jixiang Meng. Disjoint path covers of star graphs. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005599.

Qiu:2025:DDB

[QMWY25] Xiaojie Qiu, Wenchao Meng, Yingchun Wang, and Qinmin Yang. Data-driven-based fully distributed eventtriggered control for nonlinear multi-agent systems. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000347.

Quan:2025:CDS

[QTW25] Ji Quan, Leyao Tao, and Xianjia Wang. Cooperation dynamics of social dilemmas on an opiniondependent weighted network. Applied Mathematics and Computation, 502 (??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002024.

Qu:2026:DCM

[QWWW26] Qixing Qu, Chengxi Wei, Mingxian Wang, and Yaxin Wang. From disclosure to compliance: Modeling the evolutionary dynamics of tripartite privacy regulation. Applied Mathematics and Computation, 512(??): ??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004989.

Qi:2025:ETS

[QX25a] Xiaojing Qi and Shengyuan Xu. Event-triggered samplingbased singularity-free fixedtime control for nonlinear systems subject to input saturation and unknown control directions. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005319.

Qiao:2025:MAS

[QX25b] Yu Qiao and Xiangtuan Xiong. A mollifier approach to the simultaneous identification of the unknown source and initial distribution in a spacefractional diffusion equation. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006362.

Qiu:2025:RKH

[QXZ25] Renjun Qiu, Ming Xu, and Pengfei Zhu. Reproducing kernel Hilbert space method for high-order linear Fredholm integro-differential equations with variable coefficients. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006222.

Qiu:2025:ITR

[QZ25] Jixiu Qiu and Yonghui Zhou.

Insider trading at a random deadline with correlation between dynamic asset and stochastic liquidity. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005812.

Raikov:2025:KPM

[RB25] I. O. Raikov and Y. M. Beltukov. The kernel polynomial method based on Jacobi polynomials. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006684.

Radojev:2025:MDT

[RBT25] G. Radojev, M. Brdar, and Lj. Teofanov. A modification of Durán-type mesh for singularly perturbed problems. Applied Mathematics and Computation, 495(??): ??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000669.

Ren:2026:IOI

[RHSZ26] Yuanhong Ren, Dingli Hua, Mingxuan Shen, and Guangchen Zhang. Inverse optimal incremental control of nonlinear jump-diffusion systems. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003698.

Rajendran:2025:PTS

[RK25] Sangeetha Rajendran and Palanivel Kaliyaperumal. Prescribed-time synchronization of hyperchaotic fuzzy stochastic PMSM model with an application to secure communications. Applied Mathematics and Computation, 493 (??):??, May 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324007185.

Roy:2025:MVG

[RKS25] Dhanya Roy, Sandi Klavzar, and Aparna Lakshmanan S. Mutual-visibility and general position in double graphs and in Mycielskians. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005927.

Ramezanian:2025:IMF

[RMS25] Reza Ramezanian, Mohammad Hosein Mahdavi, and Shahrooz Shahparvari. Integrated mobile facility production and distribution scheduling planning: a synchronized solution framework. Applied Mathematics and Computation, 494 (??):??, June 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000049.

Rebollo-Neira:2025:RCB

[RNL25] Laura Rebollo-Neira and Jason Laurie. Recursive construction of biorthogonal polynomials for handling polynomial regression. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003042.

Rublev:2025:IAA

[RPD25] G. D. Rublev, A. N. Parshikov, and S. A. Dyachkov. Improving approximation accuracy in Godunov-type smoothed particle hydrodynamics methods. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005897.

Romero:2025:RSM

[RPS^{+}25] Pau Romero, Abel Pedrós, Rafael Sebastian, Miguel Lozano, and Ignacio García-Fernández. A robust shape model for blood vessels analysis. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005393.

Ruan:2025:AMT

[RQWX25] Zongli Ruan, Guobing Qian, Yiwei Wang, and Jia Xiao. Application of multiple twodimensional interpolation methods to bathymetric data on the FVCOM unstructured triangular grid. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002905.

Ruan:2026:SBD

[RR26] Linheng Ruan and Iryna Rybak. Stokes–Brinkman– Darcy models for fluidporous systems: derivation, analysis and validation. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004138.

R:2019:MME

[RRR19] Padma R., Ponalagusamy R., and Tamil Selvi R. Mathematical modeling of electro hydrodynamic non-Newtonian fluid flow through tapered arterial stenosis with periodic body acceleration and applied magnetic field. Applied Mathematics and Computation, 362(??):Article 124453, December 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0096300319304175.

See corrigenda [RRR20, PPS25].

R:2020:CMM

[RRR20] Padma R., Ponalagusamy R., and Tamil Selvi R. Corrigendum to “Mathematical modeling of electro hydrodynamic non-Newtonian fluid flow through tapered arterial stenosis with periodic body acceleration and applied magnetic field” [Applied Mathematics and Computation, 362 (2019) 124453]. Applied Mathematics and Computation, 373(??):??, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300319310239.

See [RRR19, PPS25].

Ramirez:2025:OII

[RS25] Hugo E. Ramírez and Rafael Serrano. Optimal investment with insurable background risk and nonlinear portfolio allocation frictions. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004843.

Rujivan:2025:APV

[Ruj25] Sanae Rujivan. Analytically pricing volatility options and capped/floored volatility swaps with nonlinear payoffs in discrete observation case under the Merton jumpdiffusion model driven by a nonhomogeneous Poisson process. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004909.

Ren:2025:OWT

[RZ25] Guangbin Ren and Xin Zhao. Octonionic wavelet transform and uncertainly principle. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300325001766.

Salehi:2026:DBT

[SA26] Khayyam Salehi and Javad Alidousti. The dynamic behavior of time-delayed Leslie–Gower predator-prey model. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003480.

Suchomelova:2025:CAH

[SBK25] Monika Suchomelová, Michal Benes, and Miroslav Kolár. Computational aspects of hyperbolic curvature flow.

Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000281.

Shen:2025:DDB

[SDH25] Ziwen Shen, Tao Dong, and Tingwen Huang. Datadriven bipartite synchronization control of multi-agent systems with asymmetric input saturation over switching networks. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000074.

Sebek:2025:PTC

[Seb25] Stjepan Sebek. Parking triangles on a continuous ladder. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001584.

Shi:2026:CET

[SFL26] Wei Shi, Ting Fu, and Kai Liu. A class of exponential time-stepping massmomentum-energy-preserving schemes for Korteweg–de Vries equations. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003467.

Song:2026:PGM

[SG26] Jiahui Song and Zaiwu Gong. The power of gaps: Multilayer mechanisms and source tracing of complex dynamics in higher-order networks. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003510.

Su:2025:TDE

[SGH^{+}25] Xin Su, Ruqiang Guo, Shijia Hua, Linjie Liu, and Liang Zhang. Time delay effects on evolutionary dynamics of N-player trust game with diverse investment behaviors. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006386.

Sun:2025:IRW

[SH25] Yang Sun and Ai-Qi Han.

An improved RBF-WENO scheme for hyperbolic conservation laws. Applied Mathematics and Computation, 501(??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001985.

Sun:2025:PPB

[SHY25] Yu Sun, Cheng Hu, and Juan Yu. Pinning passivity and bipartite synchronization of fractional signed networks without gauge transformation. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005289.

Si:2025:ECS

[SIL25] Zehua Si, Takayuki Ito, and Hsuan-Wei Lee. Evolution of cooperation in spatial public goods games with migration and interactive diversity. Applied Mathematics and Computation, 506 (??):??, December 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500270X.

Sahebi:2026:SCS

[SKM26] Hamid Reza Sahebi, Manochehr Kazemi, and Mohamed M. A. Metwali. On solutions of a coupled system of quantum integral equations in Banach spaces. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004047.

Saiparasad:2026:REE

[SKSC26] V. R. Saiparasad, B. Kaviya, D. V. Senthilkumar, and V. K. Chandrasekar. Routes to extreme events in forcedtransmission eco-epidemic model: a dynamical-systems perspective. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003455.

Syu:2026:TTP

[SL26] Jia-Hao Syu and Yuh-Dauh Lyuu. Trapezoid and trapezoidal prism for the maximum relative drawdown: Probability, crash options pricing, and risk. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004345.

Song:2026:OIS

[SLX^{+}26] Jiayi Song, Wenjie Li, Yunzhu Xiao, Ling Chen, Chun Yang, Li Qi, and Wei Wang. Optimal information spreading strategy for containing epidemic spreading on higher-order multiplex networks. Applied Mathematics and Computation, 511 (??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004503.

Suvetha:2025:NFO

[SNP25] R. Suvetha, J. J. Nieto, and P. Prakash. Nonfragile output-feedback control for delayed memristive bidirectional associative memory neural networks against actuator failure. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400482X.

Sommariva:2025:NCH

[Som25] A. Sommariva. Numerical cubature and hyperinterpolation over spherical polygons. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000621.

Sevcik:2025:CMS

[SP25a] Jan Sevcík and Lenka Pribylová. Cycle multistability and synchronization mechanisms in coupled interneurons: In-phase and anti-phase dynamics under current stimuli. Applied Mathematics and Computation, 503(??):??, October 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002267.

Singh:2025:EBP

[SP25b] Sarishti Singh and Geetanjali Panda. Eigenvalue bounds and Perron– Frobenius theory for nonnegative or positive interval matrices. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000566.

Shahab:2024:NNB

[SS24] Muhammad Luthfi Shahab and Hadi Susanto. Neural networks for bifurcation and linear stability analysis of steady states in partial differential equations. Applied Mathematics and Computation, 483(??):??, December 15, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004466.

See corrigendum [SS25].

Shahab:2025:CNN

[SS25] Muhammad Luthfi Shahab and Hadi Susanto. Corrigendum to “Neural networks for bifurcation and linear stability analysis of steady states in partial differential equations” [appl. math. comput. 483 (2024) 128985]. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000463.

See [SS24].

Shahab:2025:FDM

[SSH25] Muhammad Luthfi Shahab, Hadi Susanto, and Haralampos Hatzikirou. A finite difference method with symmetry properties for the highdimensional Bratu equation. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005976.

Shobana:2026:CSM

[SSK^{+}26] N. Shobana, R. Sakthivel, O. M. Kwon, M. Sankar, and M. Asha Safana. Composite sliding mode adaptive tracking control for singularly perturbed semi-Markov jump non-linear systems with fault and disturbances. Applied Mathematics and Computation, 511 (??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004436.

Sennour:2026:SMS

[SST26] M. Sennour, D. Sbibih, and M. Tahrichi. Superconvergent methods for solving two-dimensional Hammerstein integral equations. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500462X.

Shieh:2025:ISP

[ST25] Chung-Tsun Shieh and Tzong-Mo Tsai. Inverse spectral problems for Sturm–Liouville operators with many frozen arguments. Applied Mathematics and Computation, 492 (??):??, May 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006969.

Stawiski:2025:DIS

[Sta25] Marcin Stawiski. Distinguishing infinite star-free graphs. Applied Mathematics and Computation, 495 (??):??, June 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000372.

Shi:2025:DDE

[SW25] Liangyao Shi and Jing Wang. Decentralized dynamic event-triggered passive bipartite synchronization for semi-Markov jump cooperation-competition neural networks under hybrid random cyber-attacks. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002863.

Sun:2025:ETA

[SWLC25] Zong-Yao Sun, Xing-Xing Wang, Jiao-Jiao Li, and Chih-Chiang Chen. Eventtriggered adaptive control for nonlinearly parameterized higher-order systems with sensor fault. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001699.

Song:2025:OBD

[SWW25] Hongchao Song, Zhenlei Wang, and Xin Wang. Observer-based dynamic event-triggered second-level MPC for nonlinear timedelay CPSs under joint hybrid attacks. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001183.

Sun:2025:ISS

[SWXL25] Wenjun Sun, Mingzhu Wang, Jianwei Xia, and Xiaodi Li. Input-to-state stabilization of nonlinear systems with impulsive disturbance via event-triggered impulsive control. Applied Mathematics and Computation, 497(??):??, July 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500092X.

Song:2026:CPL

[SZLG26] Xiaodi Song, Shenggui Zhang, Jing Li, and Shanshan Gao. The consensus performance of leaderfollower multi-agent systems under reversing the directions of arcs. Applied Mathematics and Computation, 509(??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003753.

Shi:2025:ROF

[SZM25] Hongpeng Shi, Zhenli Zhao, and Shuping Ma. Reducedorder full information estimation observer-based dissipative control for nonlinear discrete-time switched singular systems with unknown inputs. Applied Mathematics and Computation, 493 (??):??, May 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007173.

Song:2025:SFF

[SZSW25] Xiaona Song, Ruikang Zhuo, Shuai Song, and Ke Wang. Singularity-free fixed-time adaptive event-triggered faulttolerant control for a QUAV with user-defined tracking performance. Applied Mathematics and Computation, 501(??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001948.

Shen:2025:RSE

[SZY25] Yi Shen, Jiemei Zhao, and Liqi Yu. Reachable set estimation of delayed secondorder memristive neural networks. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004557.

Takesue:2025:ECT

[Tak25] Hirofumi Takesue. Evolution of cooperation in a three-strategy game combining snowdrift and stag hunt games. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000682.

Takesue:2026:HCI

[Tak26] Hirofumi Takesue. Hinderance of cooperation by individual solutions: Evolutionary dynamics of threestrategy games combining the prisoner’s dilemma and stag hunt. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004023.

Tan:2025:SON

[Tan25] Zhijun Tan. Second-order non-uniform and fast twogrid finite element methods for non-linear timefractional mobile/immobile equations with weak regularity. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005046.

Tang:2025:AVE

[TF25] Sichen Tang and Aili Fang. Analysis of viewpoint evolution based on WeiBo data mining. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006738.

Tang:2025:NLL

[TFCL25] Zhuochao Tang, Zhuojia Fu, Meng Chen, and Leevan Ling. A novel localized leastsquares collocation method for coupled bulk-surface problems. Applied Mathematics and Computation, 492(??):??, May 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007112.

Tamimi:2025:CNS

[TGS25] Hengameh Tamimi, Mohammad Bagher Ghaemi, and Reza Saadati. Coupled nonlinear stochastic integral equations in the general form of the predatorprey model. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005848.

Torok:2025:EFI

[THPA25] Csaba Török, Juraj Hudák, Viktor Pristas, and Lubomir Antoni. Explicit forms of interpolating cubic splines and data smoothing. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001389.

Tissaoui:2025:ESE

[TKM25] Yassine Tissaoui, James F. Kelly, and Simone Marras. Efficient spectral element method for the Euler equations on unbounded domains. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005411.

Trevisani:2025:MMN

[TLSVGR25] Davide Trevisani, José Germán López-Salas, Carlos Vázquez, and José Antonio García-Rodríguez. Mathematical models and numerical methods for a capital valuation adjustment (KVA) problem. Applied Mathematics and Computation, 488(??): ??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005666.

Tao:2025:AMC

[TLW^{+}25] Ling Tao, Zhongpan Li, Hong Wen, Simiao Yu, and Zhiqiang Feng. Analysis of multiple contact types within the framework of semi-finite element method. Applied Mathematics and Computation, 502(??):??, October 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002206.

Tominec:2025:SER

[TNL25] Igor Tominec, Murtazo Nazarov, and Elisabeth Larsson. Stability estimates for radial basis function methods applied to linear scalar conservation laws. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004818.

Terao:2025:MVS

[TO25] Takeshi Terao and Katsuhisa Ozaki. Method for verifying solutions of sparse linear systems with general coefficients. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006659.

Tan:2026:IOP

[TQWW26] Qiuyue Tan, Haizhen Qiu, Yiqiao Wang, and Kan Wang. An identifying operation on a 1-planar graph with an application to acyclic coloring. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003406.

Takahara:2025:LSP

[TS25] Akihiro Takahara and Tomoko Sakiyama. Leveraging surrounding past strategies to maintain cooperation in the perverse prisoner’s dilemma. Applied Mathematics and Computation, 493(??):??, May 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324007331.

Tang:2025:NSN

[TX25] Wei Tang and Da Xu. Numerical solution of nonlinear multi-term fractional differential equations based on spline Riesz wavelets. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000864.

Tong:2026:NAE

[TXJW26] Xingcheng Tong, Xue Xiao, Xiaozheng Jin, and Hai Wang. Nonlinear adaptive estimator-based prescribed performance control of perturbed quadcopter systems. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003893.

Tang:2026:SPM

[TZL26] Ling Tang, Yanjun Zhang, and Hanyu Li. On sketchand-project methods for solving tensor equations. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004606.

Ueda:2025:IZD

[Ued25] Masahiko Ueda. On the implementation of zerodeterminant strategies in repeated games. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324006404.

Villarino:2026:DLC

[VL26] Joel P. Villarino and Alvaro Leitao. On deep learning for computing the dynamic initial margin and margin value adjustment. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004059.

Veiga:2026:LRS

[VMM26] Maria Han Veiga, Faezeh Nassajian Mojarrad, and Fatemeh Nassajian Mojarrad. Learning a robust shape parameter for RBF approximation. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004321.

Wang:2025:CET

[Wan25]  Huage Wang.  Calculat-

ing the C-eigenvalues of the three-dimensional piezoelectric tensors directly. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007306.

Wang:2025:GAS

[WDMM25] Xin Wang, Shihong Ding, Li Ma, and Keqi Mei. Global asymptotic stabilization via sampled-data control for a class of time-delayed systems. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006507.

Wettergren:2025:GSD

[Wet25] Thomas A. Wettergren.

Group size dynamics for a group following game with shared rewards. Applied Mathematics and Computation, 501(??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001961.

Wang:2025:HDG

[WF25] Lu Wang and Minfu Feng.

On the hybridizable discontinuous Galerkin method and superconvergence analysis for the diffusive viscous wave equation. Applied Mathematics and Computation, 501(??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001973.

Wang:2025:USS

[WGeL25] Shuaikang Wang, Yongbin Ge, and Sheng en Liu. Unconditionally stable sixthorder structure-preserving scheme for the nonlinear Schrödinger equation with wave operator. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001195.

Wu:2026:IWF

[WH26] Jian Ming Wu and Cong Huang. Improved WENO finite difference method: Treating the multiple discontinuities. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004874.

Wang:2025:EHC

[WHCB25] Mei-Li Wang, Rong-Xia Hao, Jou-Ming Chang, and Sejeong Bang. 2-edge Hamiltonian connectedness: Characterization and results in data center networks. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006581.

Wang:2025:RIT

[WJ25] Q. Wang and Z. M. Jin.

Rainbow independent triangles in complete split graphs. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003157.

Wei:2025:MLI

[WJBL25] Peichao Wei, Wuding Jia, Francesco Belardo, and Muhuo Liu. Maximal Lanzhou index of trees and unicyclic graphs with prescribed diameter. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324005770.

Wang:2025:EFD

[WL25] Yang Wang and Shounan Lu. Environmental-feedback-driven time-varying reputation update solve social dilemmas. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002887.

Wang:2026:GCF

[WL26] Jian Wang and Zhi-Min Li. H_{∞} guaranteed cost filtering for nonlinear system with limited bandwidth and stochastic hybrid attacks. Applied Mathematics and Computation, 512 (??):??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004400.

Wang:2026:SAS

[WLLL26a] Hai-Su Wang, Zhong-Rong Lu, Ji-Ke Liu, and Guang Liu. Semi-analytical solution and nonlinear characterization analysis of fractionalorder nonlinear systems based on the time-domain minimum residual method. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003376.

Wang:2026:SAE

[WLLL26b] Weijie Wang, Hongliang Liu, Haodong Li, and Shoufu Li. Stability analysis of extended canonical Euler splitting method for nonlinear composite stiff neutral functional differential-algebraic equations. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004837.

Wang:2026:EAE

[WLW26] Jiawen Wang, Yangyang Luan, and Xiaoqun Wu. Equilibrium analysis of edgeheterogeneous binary network games. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004485.

Wang:2025:SDB

[WLZ25a] Wei Wang, Jin-Ming Liang, and Hong-Bing Zeng. Sampleddata-based stability and stabilization of Lurie systems. Applied Mathematics and Computation, 501 (??):??, September 15, 2025.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001821.

Wang:2025:PEC

[WLZ25b] Yaoping Wang, Shasha Li, and Zeng Zhao. k-path-edge-connectivity of the complete balanced bipartite graph. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000220.

Wang:2026:ODF

[WLZ26] Guoliang Wang, Yaqiang Lyu, and Yande Zhang. Optimal distributed finite-time consensus control of multiagent systems with Markovian switching topologies and applications to microgrids. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003558.

Wang:2025:TGR

[WMW^{+}25] Chen Wang, Junchi Ma, Chao Wang, Defu Lin, and Weilin Chen. Two graphs: Resolving the periodic reversibility of onedimensional finite cellular automata. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400612X.

Wang:2025:THO

[WMWY25] Shengdi Wang, Tingfu Ma, Lili Wu, and Xiaojia Yang. Two high-order compact finite difference schemes for solving the nonlinear generalized Benjamin–Bona– Mahony–Burgers equation. Applied Mathematics and Computation, 496(??):??, July 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000876.

Wang:2025:CMM

[WMZ25a] Jun Wang, Qiang Ma, and Cheng Zhou. Corrigendum to “Min–min minimization for the fractional l0-regularized problem” [Applied Mathematics and Computation Volume 503, 15 October 2025, 129499]. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/

                                        science/article/pii/S0096300325002929.

See [WMZ25b].

Wang:2025:MMM

[WMZ25b] Jun Wang, Qiang Ma, and Cheng Zhou. Min–min minimization for the fractional l_{0}-regularized problem. Applied Mathematics and Computation, 503 (??):??, October 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002255.

See corrigendum [WMZ25a].

Wendler:2026:NNS

[WPTK26] Anna Wendler, Lena Plötzke, Hannah Tritzschak, and Martin J. Kühn. A nonstandard numerical scheme for a novel SECIR integrodifferential equation-based model allowing nonexponentially distributed stay times. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003625.

Wang:2025:STR

[WQLW25] Jiayu Wang, Lin Qiu, Yingjie Liang, and Fajie Wang. A spatiotemporal radial basis function collocation method based on Hausdorff fractal distance for Hausdorff derivative heat conduction equations in threedimensional anisotropic materials. Applied Mathematics and Computation, 502 (??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002279.

Wu:2026:AFT

[WS26] Xueqi Wu and Wei Sun.

Adaptive fault tolerant optimal control for high-order fully actuated system and its application in spacecraft attitude tracking using reinforcement learning. Applied Mathematics and Computation, 512(??): ??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004990.

Wang:2025:MSB

[WSX^{+}25] H.-Z. Wang, X.-C. Shangguan, D. Xiong, Y.-H. An, and L. Jin. A matrix-separation-based integral inequality for aperiodic sampled-data synchronization of delayed neural networks considering communication delay. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004934.

Wang:2025:CEP

[WSY25] Xianjia Wang, Chuyue Song, and Zhipeng Yang. Cooperation evolution of public goods games under probabilistic punishment rule for reputational altruism. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002322.

Wei:2025:SCN

[WSZM25] Yong Wei, Mouquan Shen, Zheng H. Zhu, and Xianji Meng. Synchronization of complex networks with multiple disturbances via aperiodically intermittent control input. Applied Mathematics and Computation, 502(??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002218.

Wu:2025:GCC

[Wu25]  Hong-Yu Wu.  A gen-

eral class of constraint preconditioners for generalized saddle point linear systems. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400609X.

Wang:2025:LIP

[WWG^{+}25] Yuchen Wang, Huidi Wang, Chao Gao, Kefeng Fan, Hailong Cheng, Zhijie Shen, Zhen Wang, and Matjaž Perc. Learning influence probabilities in diffusion networks without timestamps. Applied Mathematics and Computation, 503(??):??, October 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002280.

Wu:2025:PGN

[WWK25] Qingqin Wu, Weifan Wang, and Jiangxu Kong. 1-planar graphs with no 5-cycles are 5-degenerate. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000311.

Wang:2025:PCF

[WWL25] Yunlong Wang, Weifan Wang, and Runrun Liu. Proper conflict-free 6-coloring of planar graphs without short cycles. Applied Mathematics and Computation, 499(??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001328.

Wang:2025:MES

[WWLZ25] Yuanhua Wang, Ying Wang, Haitao Li, and Wenke Zang. Matrix expressions of symmetric n-player games. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005952.

Wang:2025:NST

[WWQX25] Xin Wang, Xiaoping Wang, Haitao Qi, and Huanying Xu. Numerical simulation of time fractional Allen–Cahn equation based on Hermite neural solver. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006957.

Wei:2025:FRD

[WWS^{+}25] Jiabao Wei, Hai Wang, Kaibo Shi, Shuping He, and Chengcheng Ren. Finiteregion dissipative control for two-dimensional Roesser systems via Markov jumping mechanism. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005678.

Wang:2025:CDD

[WWY^{+}25] Zhishuang Wang, Yicong Wan, Qian Yin, Zhiyong Hong, Qiuxia Xu, and Chengyi Xia. Coupled diffusion dynamics of competitive information and green behaviors on multiplex networks under policy intervention. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000554.

Wang:2025:RFM

[WX25] Huaijin Wang and Chuanju Xu. On recurrence formulae of Müntz polynomials and applications. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006271.

Wang:2025:PEM

[WXD^{+}25] Junwei Wang, Weili Xiong, Feng Ding, Yihong Zhou, and Erfu Yang. Parameter estimation method for separable fractional-order Hammerstein nonlinear systems based on the on-line measurements. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005630.

Wang:2025:CRN

[WXT25] Shengrong Wang, Jie Xie, and Li Tan. Convergence rate of nonlinear delayed neutral McKean– Vlasov SDEs driven by fractional Brownian motions. Applied Mathematics and Computation, 502 (??):??, October 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002048.

Wang:2026:SAR

[WY26a] Gehao Wang and Yuexin Yu. Stability analysis of Runge–Kutta methods for nonlinear delay-integro-differential-algebraic equations. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004011.

Wang:2026:RRR

[WY26b] Jiaoyuan Wang and Yanlong Yang. Research on the resilience of reputation mechanism to the cooperative environment in the face of external shocks. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004394.

Wu:2025:PTL

[WYKS25] Jie Wu, Rongni Yang, Levente Kovacs, and Peng Shi. Predefined-time Lyapunov stability for nonlinear systems via a dynamic event-triggered control strategy. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500284X.

Wu:2025:AEC

[WZJW25] Yifan Wu, Guangming Zhuang, Mengmeng Jiang, and Yanqian Wang. H_{∞} adaptive ETF control for degenerate jump systems under actuator faults and impulsive deception attacks with application to DCMIP device. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006192.

Wang:2026:PTC

[WZLS26] Ce Wang, Wei Zhao, Shaoyu Lv, and Hao Shen. Predefinedtime control of non-strict feedback nonlinear systems subject to input saturation and output constraint: a reinforcement learning method. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500342X.

Wu:2025:BTC

[WZP^{+}25] Donghui Wu, Ying Zhao, Mingchuan Pang, Hong Sang, Shuanghe Yu, and Peng Wang. Bumpless transfer control of switched systems: a dynamic eventtriggering switching control approach. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006167.

Wang:2025:LBS

[WZWC25a] Bo Wang, Fucheng Zou, Junhui Wu, and Jun Cheng. A learning-based sliding mode control for switching systems with dead zone. Applied Mathematics and Computation, 494 (??):??, June 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000104.

Wu:2025:OBF

[WZWC25b] Huaiqin Wu, Xueqing Zhao, Lingyan Wang, and Jinde Cao. Observer-based fixedtime topology identification and synchronization for complex networks via quantized pinning control strategy. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002942.

Wang:2026:GSC

[WZXY26] Huiru Wang, Yongtao Zhou, Wenzhuo Xiong, and Xiaoqiang Yan. Generalized Störmer-Cowell methods for the nonlinear second-order delay integro-differential equations with initial conditions. Applied Mathematics and Computation, 512(??): ??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004965.

Wang:2025:KIC

[WZY25] Yan Wang, Qian Zhang, and Kai Yuan. Kirchhoff indices of Cayley graphs on D_{2n} × Z_{m}. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005939.

Xia:2025:OBS

[XC25] Pei-Zhen Xia and Xiao-Heng Chang. Observer-based security control for nonlinear networked systems: the SCP model of semi-Markov kernel approach. Applied Mathematics and Computation, 498(??):??, August 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001213.

Xin:2025:FID

[XGL25] Qi Xin, Xian-Ming Gu, and Li-Bin Liu. A fast implicit difference scheme with nonuniform discretized grids for the time-fractional Black–Scholes model. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001687.

Xiao:2025:SPG

[XLLC25] Liuchao Xiao, Minghao Li, Zhenzhen Li, and Hongru Chen. The supercloseness property and global superconvergence analysis for the transient conductionconvection equations. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300325000694.

Xie:2026:RLE

[XS26] Kai Xie and Attila Szolnoki.

Reinforcement learning in evolutionary game theory: a brief review of recent developments. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004114.

Xu:2025:PJI

[XSB25] Wei-Ru Xu, Qian-Yu Shu, and Natália Bebiano. A pseudo-Jacobi inverse eigenvalue problem with a rankone modification. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324005794.

Xu:2025:ITO

[XTGZ25] Tongyou Xu, Ying-Ying Tan, Shanshan Gao, and Xuejuan Zhan. Interaction topology optimization by adjustment of edge weights to improve the consensus convergence and prolong the sampling period for a multi-agent system. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001559.

Xu:2025:GWM

[XW25] Si-Ao Xu and Baoyindureng Wu. Graphs whose the maximum size of an odd subgraph equal to ⌊n/2⌋. Applied Mathematics and Computation, 494(??):??, June 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000244.

Xiao:2026:SET

[XWW^{+}26] Shuai Xiao, Zhen Wang, Hongjue Wang, Yueying Liu, and Xia Huang. Switching event-triggered control for mean-square exponential synchronization of complexvalued delayed memristive neural networks under hybrid attacks. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004412.

Xin:2025:EPV

[XZN^{+}25] Yue Xin, Yi Zhang, Idin Noorani, Farshid Mehrdoust, and Jinwu Gao. Estimation of parameters and valuation of options written on multiple assets described by uncertain fractional differential equations. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005708.

Xu:2026:FTC

[XZSC26] Nuo Xu, Yanzheng Zhu, Chun-Yi Su, and Xinkai Chen. Fault-tolerant control for Markov jump PWA systems against actuator lose efficacy faults: an adaptive iterative learning approach. Applied Mathematics and Computation, 512 (??):??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004904.

Yang:2025:CFB

[YBSC25] Yu Yang, Wenshan Bi, Shuai Sui, and C. L. Philip Chen. Command-filter-based neural networks predefined time control for switched nonlinear systems with event-triggering mechanism. Applied Mathematics and Computation, 491 (??):??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006660.

Yang:2025:SSD

[YFZ^{+}25] Feifei Yang, Shihua Fu, Jie Zhong, Xianghui Su, and Hao Zhang. Stabilization of sampled-data generalized asynchronous Boolean control networks under noise interference. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002875.

Yang:2025:ETA

[YGY^{+}25] Yang Yang, Shuocong Geng, Dong Yue, Sergey Gorbachev, and Iakov Korovin. Event-triggered approximately optimized formation control of multiagent systems with unknown disturbances via simplified reinforcement learning. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006106.

Yu:2025:CEM

[YH25] Yue Yu and Liang’an Huo.

A co-evolutionary model of information, behavior, and epidemics in multiplex networks: Incorporating subjective and objective factors. Applied Mathematics and Computation, 499 (??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500133X.

Yan:2025:CFO

[YHGZ25] Jiayuan Yan, Bin Hu, Zhi-Hong Guan, and Ding-Xue Zhang. On controllability of fractional-order impulsive and switching systems with time delay. Applied Mathematics and Computation, 497(??):??, July 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000840.

Yang:2026:OBL

[YHSZ26] Dongsheng Yang, Jing He, Jiayue Sun, and Juan Zhang. Observer-based leader-follower bipartite consensus for linear multiagent systems with communication link faults. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004163.

Yu:2026:CSA

[YK26] Guihai Yu and Yuwei Kang. Can synergy be achieved in global vaccine and medical device trade? Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003303.

Ye:2026:MAB

[YL26] Qingfang Ye and Fengwei Li. The minimum atom-bond sum-connectivity index of graphs with given degree sequence. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003935.

Yu:2025:REF

[YLR^{+}25] Xiaotong Yu, Haili Liang, Xiaoqiang Ren, Zhihai Rong, Xiaofan Wang, and Ming Cao. The role of environmental feedback in promoting cooperation among unequal groups. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001638.

Yuan:2025:CCW

[YM25] Lili Yuan and Jixiang Meng. Classical conditions for (weakly-)antistrong digraphs. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003194.

Yang:2025:RMD

[YNZ25] Luhe Yang, Yongpeng Ning, and Lianzhong Zhang. Reconstruction of micro-dynamics characterizing human decisionmaking behavior in repeated social dilemmas. Applied Mathematics and Computation, 497(??):??, July 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000943.

Yao:2026:CCM

[YQQ^{+}26] Caibin Yao, Yiwen Qi, Ziyu Qu, Wu Wang, and Zonghua Zheng. Consensus control for multi-agent systems with unknown fading channels and limited bandwidth. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004060.

Yu:2025:TRT

[YQYS25] Yuan Yu, Zuojian Qin, Haizhuan Yuan, and Shi Shu. Two-relaxation-time regularized lattice Boltzmann model for convectiondiffusion equation with spatially dependent coefficients. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005964.

Yang:2025:ERT

[YT25] Qing Yang and Yingzhi Tian. Exploring redundant trees in bipartite graphs. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004673.

Yu:2025:QIF

[Yu25]  Jiao Yu.  Quantum in-

tegral favard-type inequality. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001791.

Yu:2025:LBA

[YWJ^{+}25] Yaolong Yu, Zhengtian Wu, Baoping Jiang, Huaicheng Yan, and Yichen Lu. A Lagrange barrier approach for the minimum concave cost supply problem via a logarithmic descent direction algorithm. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005757.

Yang:2025:SSM

[YWR^{+}25] Dongsheng Yang, Hu Wang, Guojian Ren, Yongguang Yu, and Xiao-Li Zhang. Synchronization of short memory fractional coupled neural networks with higher-order interactions via novel intermittent control. Applied Mathematics and Computation, 497(??):??, July 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000906.

Yu:2026:CIP

[YXG26] Xilai Yu, Kefan Xia, and Shiping Gao. Cooperation and the incentive power distribution on multilayer hierarchical networks. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004849.

Yamakawa:2025:CAR

[YY25a] Yuya Yamakawa and Nobuo Yamashita. Convergence analysis of a regularized Newton method with generalized regularization terms for unconstrained convex optimization problems. Applied Mathematics and Computation, 491(??):??, April 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324006805.

Yang:2025:EUM

[YY25b] Qianxi Yang and Yanlong Yang. Effects of an update mechanism based on combinatorial memory and high-reputation learning objects on the evolution of cooperation. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000360.

Yang:2026:DCE

[YY26] Bolin Yang and Guanghui Yang. Dynamics of cooperative evolution with leader-follower hierarchy under information uncertainty in two-layer grid networks. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004126.

Yang:2026:AFN

[YYMY26] Qing Yang, Haisheng Yu, Xiangxiang Meng, and Wenqian Yu. Adaptive fuzzy neural network-based smooth-switching gain dynamic surface control for constrained manipulator with error-based nonlinear disturbance observer. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004680.

Yang:2025:CAE

[YYT25] Yin Yang, Pai Yao, and Emran Tohidi. Convergence analysis of an efficient multistep pseudo-spectral continuous Galerkin approach for solving Volterra integro-differential equations. Applied Mathematics and Computation, 494 (??):??, June 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000116.

You:2025:GNH

[YYW^{+}25] Tao You, Linjiang Yang, Jian Wang, Peng Zhang, Jinchao Chen, and Ying Zhang. The guidance of neutral human populations maintains cooperation in the prisoner’s dilemma game. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005320.

Yuan:2025:MMI

[YYZG25] Bo-Jun Yuan, Zhao-Yu Yang, Lu Zheng, and Shi-Cai Gong. Maximal and maximum induced matchings in connected graphs. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001596.

Yang:2025:PBE

[YZ25] Meng Yang and Junyong Zhai. Predictor-based eventtriggered learning control of networked control systems with false data injection attacks and output delay. Applied Mathematics and Computation, 490(??): ??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400674X.

Ye:2025:WDS

[YZH25] Jing Ye, Jiani Zou, and Miaomiao Han. Weak degeneracy of the square of K_{4}-minor free graphs. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001663.

Yang:2025:FTA

[YZM25] Yayu Yang, Mingzu Zhang, and Jixiang Meng. Fault tolerance assessment for Hamming graphs based on rrestricted R-structure (substructure) fault pattern. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006210.

Zhang:2025:REC

[ZBM25] Hong Zhang, Hong Bian, and Jixiang Meng. Reliability evaluation of conditional recursive networks under h-conditional restriction. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001262.

Zhang:2025:SBS

[ZC25] Zhen Zhang and Wei-Wei Che. SOS-based secure consensus control for stochastic nonlinear MASs against multiple DoS attacks. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003212.

Zhang:2025:OSC

[ZCJL25] Jiannan Zhang, Ping Chen, Zhuo Jin, and Shuanming Li. Optimal strategies for collective defined contribution plans when the stock and labor markets are cointegrated. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006714.

Zhang:2025:OTR

[ZD25] Bin Zhang and Da-Wei Ding. Optimal tracking and regulation performance of networked control systems with additive colored Gaussian noise and finite bandwidth. Applied Mathematics and Computation, 501 (??):??, September 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500195X.

Zhang:2025:GSA

[ZFFM25] Yuanyuan Zhang, Shaochuan Fu, Shucheng Fan, and Fangfang Ma. Game strategy analysis on e-commerce platform supply chain with shared logistics service: a chaos perspective. Applied Mathematics and Computation, 499(??):??, August 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001419.

Zhang:2025:MCS

[ZFW25] Ce Zhang, Likang Feng, and Zhaojing Wu. Maneuvering control of stochastic nonlinear systems with unknown covariance noise. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325001432.

Zaugg:2026:BOV

[ZG26a] Nicola F. Zaugg and Lech A. Grzelak. Basket options with volatility skew: Calibrating a local volatility model by sample rearrangement. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003959.

Zhu:2026:DRU

[ZG26b] Tielei Zhu and Zhihao Ge. Density results using the method of fundamental solutions for a fluidstructure interaction problem. Applied Mathematics and Computation, 512(??): ??, March 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004941.

Zhang:2025:ECA

[Zha25]  Hong Zhang.  Evolu-

tion of cooperation among fairness-seeking agents in spatial public goods game. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006441.

Zhang:2025:LEG

[ZHF25] Jiahao Zhang, Changxiang He, and Rongquan Feng. On the least eigenvalue of genuine strongly 3-walk-regular graphs. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006635.

Zhou:2025:PMC

[ZHL25] Wenjuan Zhou, Rong-Xia Hao, and Yilun Luo. Perfect matching cover indices of generalized rotation snarks. Applied Mathematics and Computation, 487 (??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005629.

Zhang:2025:VRP

[ZHLL25] Weiting Zhang, Guitian He, Maokang Luo, and Wenjie Liang. Valuation of R&D projects of new energy vehicles based on generalized mixed sub-fractional Brownian motion under fuzzy environment. Applied Mathematics and Computation, 507 (??):??, December 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325002851.

Zhang:2026:NAO

[ZHWW26] Tong Zhang, Yiyan Han, Ling Wang, and Xin Wang. Neuro-adaptive optimized control for stochastic systems with state constraints: the non-affine faults case. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004461.

Zhao:2026:SEC

[ZHY26] Shuang Zhao, Xiaomin Hu, and Weihua Yang. Super edge-connectivity of transitive hypergraphs. Applied Mathematics and Computation, 509(??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004242.

Zuo:2026:SMC

[ZHZ^{+}26] Zhiyuan Zuo, Jun Hu, Hongxu Zhang, Hao Liu, and Guangchen Zhang. Sliding mode control for T-S fuzzy delayed systems via event-triggered round-robin protocol. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003595.

Zhang:2025:ZDS

[ZJX25] Zhipeng Zhang, Xiaotong Jiang, and Chengyi Xia. Zero-determinant strategy of finite games with implementation errors and its application into group decisionmaking. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005162.

Zhao:2026:FTS

[ZKZ26] Xin-Yue Zhao, Qingkai Kong, and Jianping Zhou. Finite-time security E2P filter design of fuzzy discretetime singular system. Applied Mathematics and Computation, 511(??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004667.

Zhang:2025:ESV

[ZL25] Wenwen Zhang and Pingrun Li. Existence of solutions for Volterra singular integral equations in the class of differentiable functions. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001754.

Zhang:2025:BFB

[ZLA25] Lili Zhang, Chenglong Liu, and Liwei An. Barrierfunction based adaptive trajectory tracking control for high-order nonlinear systems with collision avoidance. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400465X.

Zhu:2025:DET

[ZLCL25] Guanchao Zhu, Min Luo, Guozeng Cui, and Ze Li. Dynamic event-triggered predefined-time adaptive attitude control for a QUAV with unknown deception attacks. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004648.

Zhang:2026:FFT

[ZLCS26] Yuzhuo Zhang, Mengzhuo Luo, Jun Cheng, and Kaibo Shi. T-S fuzzy fault-tolerant load frequency control for power systems with energy storage system under DoS attacks and actuator failures: a two-step adaptive event-triggered mechanism. Applied Mathematics and Computation, 508 (??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003583.

Zhao:2026:PTR

[ZLH26] Benxin Zhao, Yuan-Xin Li, and Zhongsheng Hou. Prescribed-time reinforcement learning formation control of nonlinear MASs with an unknown dynamic leader. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004370.

Zhou:2025:ACD

[ZLHS25] Mengyu Zhou, Xingwen Liu, Qi Hu, and Feng Shu. Analysis and control of demand response in smart grids: an evolutionary game method. Applied Mathematics and Computation, 488(??):??, March 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005915.

Zhang:2025:OBN

[ZLJY25] Pengcheng Zhang, Yajuan Liu, Shiyu Jiao, and Chen Yang. Observer-based non-fragile control for T-S fuzzy switched systems against cyber attacks: a double-layer PDT switching method. Applied Mathematics and Computation, 495(??):??, June 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325000359.

Zhao:2026:PFT

[ZLL26] Yajing Zhao, Xiaoping Liu, and Cungen Liu. Prescribed finite-time collisionavoidance control for highorder multi-agent systems with external disturbances. Applied Mathematics and Computation, 509(??):??, January 15, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500387X.

Zhou:2026:MPS

[ZLSQ26] Yinglu Zhou, Yinya Li, Andong Sheng, and Guoqing Qi. Multi-pursuer singleevader privacy-preserving differential games. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003388.

Zhang:2025:CCW

[ZLW25] Guozhen Zhang, Xin Liu, and Dajin Wang. Component connectivity of wheel networks. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324005575.

Zhang:2026:MGA

[ZLW^{+}26] Shuo Zhang, Lu Liu, Chunhua Wang, Xiaomeng Zhang, and Rong Ma. Multistability and global attractivity for fractional-order spiking neural networks. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003431.

Zhou:2026:CMS

[ZLY26] Jinqiu Zhou, Danyi Li, and Weigen Yan. Counting matchable spanning trees of join graphs. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004977.

Zhang:2026:RMP

[ZMM26] Daochang Zhang, Leiming Ma, and Dijana Mosić. Representations of the Moore– Penrose inverse of a 2 × 2 block matrix based on Schur complements. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325005016.

Zhang:2025:MES

[ZN25] Jinsen Zhang and Xiaobing Nie. Multiple exponential stability for short memory fractional impulsive Cohen–Grossberg neural networks with time delays. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005277.

Zhang:2026:AAF

[ZQS26] Xinyu Zhang, Jiangang Qi, and Chunyan Sun. Asymptotic analysis of the first eigenvalue for the Sturm– Liouville problems with coupled boundary conditions. Applied Mathematics and Computation, 510(??):??, February 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032500414X.

Zeng:2025:RTG

[ZT25] Yunhua Zeng and Zhijun Tan. An α-robust twogrid finite element method with nonuniform L2-1_{s}igma scheme for the semilinear Caputo–Hadamard timefractional diffusion equations involving initial singularity. Applied Mathematics and Computation, 496 (??):??, July 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000827.

Zhang:2025:DSL

[ZW25] Yonghong Zhang and Ligong Wang. Distance (signless) Laplacian spectra and energies of two classes of cyclic polyomino chains. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005605.

Zhu:2026:ESD

[ZWCT26] Qianlong Zhu, Qi Wu, Wei Chen, and Jun Tao. The evolution of social dilemma based on aspiration-based collective interdependence.

Applied Mathematics and Computation, 508(??):??, January 1, 2026. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003881.

Zou:2025:IIC

[ZWF25] Lantao Zou, Yongjiang Wu, and Lihua Feng. D^{L}-integral and D^{Q}-integral n-Cayley graphs. Applied Mathematics and Computation, 489 (??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006398.

Zhang:2025:ETD

[ZWL25] Zidie Zhang, Daiyong Wu, and Nishan Li. Exploring threshold dynamics in a spatially heterogeneous ecosystem with memorybased diffusion and hunting cooperation on predators. Applied Mathematics and Computation, 491(??): ??, April 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006829.

Zhai:2025:EBE

[ZWQC25] Junchang Zhai, Huanqing Wang, Yuping Qin, and Hongxia Cui. Estimationbased event-triggered fixedtime fuzzy tracking control for high-order nonlinear systems. Applied Mathematics and Computation, 486 (??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005186.

Zhang:2025:SPG

[ZXP^{+}25] Gui Zhang, Xiaojin Xiong, Bin Pi, Minyu Feng, and Matjaz Perc. Spatial public goods games with queueing and reputation. Applied Mathematics and Computation, 505(??):??, November 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002590.

Zhao:2025:ROB

[ZXS^{+}25] Dawei Zhao, Wenkang Xiang, Weizhao Song, Lijuan Xu, Chuan Chen, and Zhen Wang. Reset observerbased containment protocol via event-triggered strategy for multi-agent networks against aperiodic DoS attacks. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001420.

Zhao:2025:FDS

[ZXW^{+}25] Miao Zhao, Jianxiang Xi, Le Wang, Cheng Wang, and Yuanshi Zheng. Fully distributed self-triggered secure consensus for nonlinear multiagent systems with sequential communication link scaling attacks. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006465.

Zeng:2025:DSL

[ZYL25] Sangui Zeng, Xiulan Yang, and Jianren Long. On discrete stochastic p-Laplacian complex-valued Ginzburg– Landau equations driven by superlinear Lévy noise. Applied Mathematics and Computation, 493(??):??, May 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324007288.

Zhang:2026:FBS

[ZYY26] Huan Zhang, Xiaona Yang, and Tianqiu Yu. h-function-based state bounding results of discrete-time delayed systems. Applied Mathematics and Computation, 511 (??):??, February 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325004473.

Zhang:2025:EWB

[ZYZ^{+}25] Ruimin Zhang, Lei Yang, Hui Zhu, Linghui Li, and Yahong Wang. An enhanced wavelet-based numerical scheme with higher convergence order for a class of high-order boundary value problems. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003285.

Zhang:2025:BIS

[ZZ25a] Peng-Li Zhang and Xiao-Dong Zhang. Bounds for the incidence Q-spectral radius of uniform hypergraphs. Applied Mathematics and Computation, 490 (??):??, April 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324006623.

Zhang:2025:LDC

[ZZ25b] Qianwei Zhang and Xinran Zhang. Q-learning driven cooperative evolution with dual-reputation incentive mechanisms. Applied Mathematics and Computation, 507(??):??, December 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325003169.

Zhang:2025:ABT

[ZZ25c] Suhuan Zhang and Fanglai Zhu. Asynchronous bumpless transfer control for stochastic hidden semi-Markovian jump systems against deception attacks. Applied Mathematics and Computation, 506(??):??, December 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032500267X.

Zhao:2025:IDR

[ZZ25d] Jia-Yu Zhao and Tao Zhao.

Improving disturbance rejection and tracking performance of interval type-2 T– S fuzzy systems with timevarying delay. Applied Mathematics and Computation, 489(??):??, March 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006118.

Zheng:2025:ETD

[ZZ25e] Leyi Zheng and Yimin Zhou. Event-triggered distributed consensus control of heterogeneous multi-agent system under communication and actuator faults. Applied Mathematics and Computation, 487(??):??, February 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005381.

Zhou:2025:ADU

[ZZ25f] Dan-Dan Zhou and Ran Zhao. Adaptive distributed unknown input observer for linear systems. Applied Mathematics and Computation, 486(??):??, February 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004880.

Zhang:2026:CEN

[ZZ26a] Qifan Zhang and Shuming Zhou. Conditional (edge) neighbor connectivity of hierarchical star networks. Applied Mathematics and Computation, 509 (??):??, January 15, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003704.

Zheng:2026:BCC

[ZZ26b] Shaobo Zheng and Lei Zhou.

Bipartite containment control of multi-agent systems under channel imperfections: when bit-flips meet packetlosses. Applied Mathematics and Computation, 510 (??):??, February 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004151.

Zhang:2026:HEM

[ZZD26] Yufan Zhang, Xiao Zhang, and Feng Ding. Hierarchical estimation method for fractional-order systems based on the auxiliary model. Applied Mathematics and Computation, 512(??):??, March 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325004746.

Zhu:2025:BAF

[ZZJZ25] Yuxun Zhu, Wenjie Zuo, Jinchen Ji, and Zhengdi Zhang. Bifurcations analysis of a 3D Filippov pestnatural enemy system with stage structure for the prey. Applied Mathematics and Computation, 497 (??):??, July 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325000839.

Zhang:2025:SFB

[ZZL^{+}25] Hui Zhang, Shenglong Zhou, Geoffrey Ye Li, Naihua Xiu, and Yiju Wang. A step function based recursion method for 0/1 deep neural networks. Applied Mathematics and Computation, 488 (??):??, March 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324005903.

Zhao:2026:FFT

[ZZS^{+}26] Junsheng Zhao, Haina Zhao, Ying Song, Zong-Yao Sun, and Dengxiu Yu. Fast finitetime consensus protocol for high-order nonlinear multiagent systems based on event-triggered communication scheme. Applied Mathematics and Computation, 508(??):??, January 1, 2026. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300325003571.

Zhou:2025:QAS

[ZZT25] Hao Zhou, Qianqian Zhang, and Sanyi Tang. Qualitative analysis of the sliding vector field in a Filippov food chain model with integrated pest management strategy. Applied Mathematics and Computation, 490(??):??, April 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324006490.

Zhao:2025:EGB

[ZZTL25] Hui Zhao, Zhenyu Zhang, Igor Tchappi, and Li Li. An evolutionary gamebased Vicsek model with a fixed number of neighbors. Applied Mathematics and Computation, 500 (??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325001304.

Zhang:2025:DPR

[ZZZM25] Yongqiang Zhang, Zehao Zheng, Xiaoming Zhang, and Jinlong Ma. Dynamic punishment-reputation synergy drives cooperation in spatial public goods game. Applied Mathematics and Computation, 506(??):??, December 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300325002711.